### v1

# Low-Altitude ISAC with Rotatable Active and Passive Arrays

Abstract
This paper investigates a low-altitude integrated sensing and communication (ISAC) system that leverages cooperative rotatable active and passive arrays. We consider a downlink scenario where a base station (BS) with an active rotatable array serves multiple communication users and senses low-altitude targets, assisted by a rotatable reconfigurable intelligent surface (RIS). A rotation-aware geometry-based multipath model is developed to capture the impact of three-dimensional (3D) array orientations on both steering vectors and direction-dependent element gains. On this basis, we formulate a new optimization problem that maximizes the downlink sum rate subject to a transmit power budget, RIS unit-modulus constraints, mechanical rotation limits, and a sensing beampattern mean-squared-error (MSE) constraint. To address the resulting highly non-convex problem, we propose a penalty-based alternating-optimization (AO) framework that alternately updates the BS precoder, RIS phase shifts, and BS/RIS array rotation angles. The three blocks are efficiently handled via a convex optimization method based on quadratic-transform (QT) and majorization-minorization (MM), Riemannian conjugate gradient (RCG) on the unit-modulus manifold, and projected gradient descent (PGD) with Barzilai–Borwein step sizes, respectively. Numerical results in low-altitude geometries demonstrate that the proposed jointly rotatable BS–RIS architecture achieves significant sum-rate gains over fixed or partially rotatable baselines while guaranteeing sensing requirements, especially with directional antennas and in interference-limited regimes.
Index Terms:
Low-altitude wireless networks (LAWN), integrated sensing and communication (ISAC), rotatable array (RA), reconfigurable intelligent surface (RIS).
I Introduction
Low-altitude wireless networks (LAWNs) are emerging as a key component of next-generation intelligent infrastructures, interconnecting uncrewed aerial vehicles (UAVs), low-altitude platform stations, aerial robots, and connected ground terminals in the three-dimensional (3D) airspace below a few kilometers [1]-[4]. Typical LAWN applications, such as aerial surveillance, disaster response, smart-city sensing, and low-altitude traffic management, rely not only on reliable broadband connectivity, but also on accurate situational awareness of dynamic low-altitude objects. This naturally calls for integrated sensing and communication (ISAC) designs that can jointly support data transmission and high-resolution sensing within complex, rapidly changing airspace environments [5],[6]. Compared with conventional terrestrial cellular systems, LAWNs must cope with fast-varying spatial geometries and require shaping radio signals to illuminate aerial targets with stringent size–weight–power constraints and capture their reflections, which make efficient ISAC-oriented signal processing particularly challenging.
Multiple-input multiple-output (MIMO) technology and advanced precoding have long been recognized as key tools to exploit spatial degrees of freedom in wireless systems [7],[8]. MIMO systems provide significant gains in beamforming, spatial multiplexing, and diversity by leveraging independent or quasi-independent channel fading, thereby significantly improving spectral efficiency and link reliability. However, in classical designs, antenna arrays are fixed in position and orientation, limiting spatial control to electronic beamforming. For LAWNs, where users and targets are distributed in full 3D and may move at low altitude with complex trajectories, purely electronic beamforming is often not sufficient: system performance becomes strongly dependent on the relative 3D geometry among the base station (BS), scatterers, and low-altitude objects. This motivates the exploitation of geometry-aware architectures that can reshape both the propagation environment and the array response itself for LAWN-specific ISAC.
Reconfigurable intelligent surfaces (RISs) offer a promising approach to tailor the wireless environment via nearly passive phase-shift elements that reflect incident signals toward desired directions [9]-[11]. RIS-aided systems have been extensively investigated for coverage enhancement, interference mitigation, and energy-efficient MIMO operation. More recently, in addition to communication, RIS has been shown to improve wireless sensing, including target detection and localization, by redirecting illumination toward non-line-of-sight (NLoS) targets or acting as a passive reflector from favorable angles, thereby enriching spatial diversity [12]-[14]. In parallel, movable antennas (MA) has recently been proposed to further enhance spatial flexibility by allowing antenna elements or arrays to change their positions within a small region [15]-[18], enabling substantial gains in diversity, multiplexing, and interference management [19], and have been extended to multicast scenarios [20], as well as to two-timescale designs exploiting statistical channel state information (CSI) [21]. These works demonstrate that adding mechanical degrees of freedom to electronic beamforming can significantly improve performance.
However, most existing MA- or RIS-related works treat antenna movement or passive phase control in isolation [22], and their ISAC designs for LAWN-like geometries are much less explored: existing ISAC designs typically assume fixed arrays and do not exploit low-altitude 3D geometry control to shape the transmit beampattern and the multiuser downlink channels simultaneously [23]. Furthermore, the mechanical flexibility in the MA literature is typically modeled as small-scale position changes within a local region. In contrast, realistic low-altitude platforms (e.g., UAV-mounted arrays or rooftop BS panels) can be designed to support rotation of the entire active or passive array via low-cost mechanical actuators [24]. Such rotatable arrays (RA) can jointly adjust their boresight directions, element coordinates, and directional gains, offering a distinct and practically relevant form of geometry control [25]. In summary, RIS-aided ISAC designs that consider rotatable passive arrays and their coupling with active-array rotations at the BS have yet to be systematically studied, especially under explicit constraints on the similarity of sensing beam patterns. As a result, the fundamental question of how to jointly design rotatable active and passive arrays, RIS phase shifts, and multiuser transmit beamforming for low-altitude ISAC remains open.
Motivated by these observations, this paper investigates a low-altitude ISAC architecture with cooperatively rotatable active and passive arrays. We consider a downlink LAWN scenario where a BS equipped with an active rotatable array simultaneously serves multiple single-antenna ground users and senses low-altitude targets in monostatic mode, assisted by a rotatable RIS termed passive array. Both arrays can rotate in 3D, and the resulting orientations jointly affect (i) the element coordinates in the steering vectors and (ii) direction-dependent element/array gains through a practical directivity model. Based on the rotation-aware channel model, we explicitly couple communication and sensing performance and formulate a joint design problem that maximizes the sum rate subject to a sensing constraint, a BS transmit-power budget, and mechanical rotation limits at both the BS and RIS, with design variables given by the BS ISAC precoder, the RIS phase-shift vector, and the 3D rotation angles of the two arrays.
The main contributions of this work are summarized as
  • We propose a LAWN-oriented ISAC architecture in which a rotatable active BS array cooperates with a rotatable passive RIS array to serve multiple communication users and jointly sense low-altitude targets. We develop a rotation-aware 3D multipath channel model that captures both phase shifts from rotated antenna/element positions and direction-dependent antenna/element directional radiation pattern gains. Based on this model, we derive closed-form expressions for the multiuser downlink SINR and the monostatic sensing beampattern, and formulate a new optimization problem that maximizes the communication sum rate under a sensing beampattern mean-squared-error (MSE) constraint, power budget on the BS, unit-modulus constraint on the RIS and mechanical rotation bounds on RA, thereby revealing mechanical rotations as additional spatial degrees of freedom for LAWN ISAC.
  • We convert the constrained ISAC design into a penalized formulation with a hinge penalty on the sensing MSE and develop an alternating optimization (AO) framework that iteratively updates the BS precoder, RIS phase shifts, and array rotations. For the BS precoding block, we combine the quadratic transform (QT) and the Lagrangian dual transform (LDT) with a majorization–minimization (MM) step to obtain a tractable quadratic-constrained quadratic program (QCQP) subproblem that admits a closed-form water-filling-like update. For the RIS block, we exploit the complex-circle manifold structure and design an efficient Riemannian conjugate gradient (RCG) algorithm. For the rotation block, we derive closed-form gradients of both SINR and beampattern MSE with respect to the BS/RIS Euler angles and employ a projected gradient descent (PGD) method with Barzilai–Borwein initialization and Armijo backtracking.
  • We conduct extensive simulations using LAWN-inspired 3D geometries and compare the proposed design with several baselines, including only BS rotation, only RIS rotation, fully fixed arrays, and non-RIS configurations. The results show that the jointly rotatable BS–RIS architecture provides substantial sum-rate gains under the same sensing MSE constraint across varying transmit powers, antenna array sizes, and user densities. The gains are further amplified with realistic directional element patterns, highlighting the importance of aligning the array boresights and main lobes with both user and target directions in a low-altitude 3D space.
The rest of the paper is structured as follows: Section II introduces the system model and problem formulation. Section III presents proposed algorithm. Section IV presents the numerical results, and Section V concludes the paper.
Figure 1: Illustration of a low-altitude ISAC system with rotatable active and passive arrays.
II System Model and Problem Formulation
We consider a downlink RA-enabled ISAC system, as shown in Fig. 1, where a base station (BS) uses an active rotatable uniform planar array (UPA) with $M$ transmit antennas to serve $K$ single-antenna users while sensing multiple targets in a monostatic mode; at the same time, a passive rotatable UPA RIS with $N$ reflecting elements assists both tasks. In addition to active transmit beamforming at the BS and passive phase shifts at the RIS, both arrays can rotate in 3D. These rotations steer the array boresights and reshape the array response to improve, or balance, communication and sensing performance.
II-A Array Geometry and Rotations
We adopt a global right-handed Cartesian coordinate system (CCS) $\mathcal{F}_{0}=(O,x,y,z)$. The BS employs a co-located transmit/receive UPA with size $M=M_{x}\times M_{y}$, and the RIS is a UPA with size $N=N_{x}\times N_{y}$. Let $\overline{\boldsymbol{d}}^{\mathrm{B}}_{m}\in\mathbb{R}^{3}$ and $\overline{\boldsymbol{d}}^{\mathrm{R}}_{n}\in\mathbb{R}^{3}$ denote the unrotated local coordinates of the $m$-th BS transmit antenna and the $n$-th RIS reflecting element, respectively, and let $\boldsymbol{d}_{0}^{\mathrm{B}}$ and $\boldsymbol{d}_{0}^{\mathrm{R}}$ be their array centers in $\mathcal{F}_{0}$. Both BS and RIS employ a motorized mechanism to rotate the arrays as a whole. Based on the system’s ISAC requirements, the proper rotation angles are determined by a central control center, which is connected to the motor controllers via fiber links.
Each array rotates as a rigid body through Euler angles
$\boldsymbol{r}^{\text{B}}\triangleq\big[\zeta_{x}^{\text{B}},\zeta_{y}^{\text{B}},\zeta_{z}^{\text{B}}\big]^{T},$ (1a)
$\boldsymbol{r}^{\text{R}}\triangleq\big[\zeta_{x}^{\text{R}},\zeta_{y}^{\text{R}},\zeta_{z}^{\text{R}}\big]^{T},$ (1b)
respectively, in their local CCS. The corresponding rotation matrix generated by $\boldsymbol{r}^{\text{A}}$, where we define $\text{A}\in\{\text{B,R}\}$ for brevity, represents the rotations about their local axis with the standard intrinsic $z$-$y$-$x$ convention and is given by
$\boldsymbol{R}\left(\zeta_{x}^{\text{A}},\zeta_{y}^{\text{A}},\zeta_{z}^{\text{A}}\right)=\boldsymbol{R}_{x}\left(\zeta_{x}^{\text{A}}\right)\boldsymbol{R}_{y}\left(\zeta_{y}^{\text{A}}\right),\boldsymbol{R}_{z}\left(\zeta_{z}^{\text{A}}\right),$ (2)
where the rotation matrices on each axis are defined by
$\boldsymbol{R}_{x}\left(\zeta_{x}^{\text{A}}\right)=\left[\begin{matrix}1&0&0\\ 0&\cos\zeta_{x}^{\text{A}}&-\sin\zeta_{x}^{\text{A}}\\ 0&\sin\zeta_{x}^{\text{A}}&\cos\zeta_{x}^{\text{A}}\\ \end{matrix}\right],$ (3a)
$\boldsymbol{R}_{y}\left(\zeta_{y}^{\text{A}}\right)=\left[\begin{matrix}\cos\zeta_{y}^{\text{A}}&0&\sin\zeta_{y}^{\text{A}}\\ 0&1&0\\ -\sin\zeta_{y}^{\text{A}}&0&\cos\zeta_{y}^{\text{A}}\\ \end{matrix}\right],$ (3b)
$\boldsymbol{R}_{z}\left(\zeta_{z}^{\text{A}}\right)=\left[\begin{matrix}\cos\zeta_{z}^{\text{A}}&-\sin\zeta_{z}^{\text{A}}&0\\ \sin\zeta_{z}^{\text{A}}&\cos\zeta_{z}^{\text{A}}&0\\ 0&0&1\\ \end{matrix}\right].$ (3c)
The rotated antenna or element coordinates in $\mathcal{F}_{0}$ are then
$\boldsymbol{d}_{m}^{\text{B}}\left(\boldsymbol{r}^{\text{B}}\right)=\boldsymbol{d}_{0}^{\text{B}}+\boldsymbol{R}\left(\boldsymbol{r}^{\text{B}}\right)\overline{\boldsymbol{d}}_{m}^{\text{B}},$ (4a)
$\boldsymbol{d}_{n}^{\text{R}}\left(\boldsymbol{r}^{\text{R}}\right)=\boldsymbol{d}_{0}^{\text{R}}+\boldsymbol{R}\left(\boldsymbol{r}^{\text{R}}\right)\overline{\boldsymbol{d}}_{n}^{\text{R}}.$ (4b)
Note that rotations affect both (i) the position of antennas or elements in the phase term of the steering vector and (ii) the directivity gain of radiation patterns for the arrays, which are jointly captured through direction- and position-dependent gains introduced in the following channel modeling.
II-B Direction and Steering Vectors
For a far-field direction parameterized by an elevation-azimuth pair $(\vartheta,\varphi)$ with elevation from the $xOy$ plane and azimuth from the $x$-axis, define the direction vector
$\boldsymbol{u}(\vartheta,\varphi)\triangleq[\cos\vartheta\cos\varphi,\cos\vartheta\sin\varphi,\sin\vartheta]^{T}.$ (5)
We adopt the geometry channel model to characterize the wireless propagation environment. Let $\mathcal{Q}_{k}=\{1,\dots,Q_{k}\}$, $\mathcal{L}_{k}=\{1,\dots,L_{k}\}$, and $\mathcal{P}=\{1,\dots,P\}$ denote the set of propagation paths between the $k$-th user and the BS, between the $k$-th user and the RIS, and between the RIS and the BS, respectively; specifically, we assume that the first paths in $\mathcal{Q}_{k}$, $\mathcal{L}_{k}$, and $\mathcal{P}$, corresponding to $l_{k}=1$, $q_{k}=1$ and $p=1$, respectively, are the line-of-sight (LoS) paths, while the remaining $L_{k}-1$, $Q_{k}-1$ and P1P-1 are the non-LoS paths. For the targeted sensing area with low-altitude UAVs, we consider merely the single LoS path with a potential reflected echo signal in each desired direction.
Accordingly, we use the following indexed notational shorthands for distinct BS-user (BU), RIS-BS (RB), BS-RIS (BR), and RIS-user (RU) links, respectively
$\boldsymbol{u}_{k,l}^{\text{BU}}\!=\!\left[\cos(\vartheta_{k,l}^{\text{BU}})\cos(\varphi_{k,l}^{\text{BU}}),\cos(\vartheta_{k,l}^{\text{BU}})\sin(\varphi_{k,l}^{\text{BU}}),\sin(\vartheta_{k,l}^{\text{BU}})\right]^{T}\!,$ (6a)
$\boldsymbol{u}_{p}^{\text{RB}}\!=\!\left[\cos(\vartheta_{p}^{\text{RB}})\cos(\varphi_{p}^{\text{RB}}),\cos(\vartheta_{p}^{\text{RB}})\sin(\varphi_{p}^{\text{RB}}),\sin(\vartheta_{p}^{\text{RB}})\right]^{T}\!,$ (6b)
$\boldsymbol{u}_{p}^{\text{BR}}\!=\!\left[\cos(\vartheta_{p}^{\text{BR}})\cos(\varphi_{p}^{\text{BR}}),\cos(\vartheta_{p}^{\text{BR}})\sin(\varphi_{p}^{\text{BR}}),\sin(\vartheta_{p}^{\text{BR}})\right]^{T}\!,$ (6c)
$\boldsymbol{u}_{k,q}^{\text{RU}}\!=\!\left[\cos(\vartheta_{k,q}^{\text{RU}})\cos(\varphi_{k,q}^{\text{RU}}),\cos(\vartheta_{k,q}^{\text{RU}})\sin(\varphi_{k,q}^{\text{RU}}),\sin(\vartheta_{k,q}^{\text{RU}})\right]^{T}\!\!.$ (6d)
The associated steering vectors, jointly parameterized by the coordinates of the rotated elements in (4) and the direction vectors in (6), can be given by
$\!\boldsymbol{t}_{k,l}^{\text{BU}}(\boldsymbol{r}^{\text{B}})\!=\!\big[e^{j\frac{2\pi}{\lambda}(\boldsymbol{u}_{k,l}^{\text{BU}})^{T}\boldsymbol{d}_{1}^{\text{B}}(\boldsymbol{r}^{\text{B}})},\dots,e^{j\frac{2\pi}{\lambda}(\boldsymbol{u}_{k,l}^{\text{BU}})^{T}\boldsymbol{d}_{M}^{\text{B}}(\boldsymbol{r}^{\text{B}})}\big]^{T}\!,\!\!$ (7a)
$\!\boldsymbol{t}_{p}^{\text{BR}}(\boldsymbol{r}^{\text{B}})\!=\!\big[e^{j\frac{2\pi}{\lambda}(\boldsymbol{u}_{p}^{\text{BR}})^{T}\boldsymbol{d}_{1}^{\text{B}}(\boldsymbol{r}^{\text{B}})},\dots,e^{j\frac{2\pi}{\lambda}(\boldsymbol{u}_{p}^{\text{BR}})^{T}\boldsymbol{d}_{M}^{\text{B}}(\boldsymbol{r}^{\text{B}})}\big]^{T}\!,\!\!$ (7b)
$\!\boldsymbol{t}_{p}^{\text{RB}}(\boldsymbol{r}^{\text{R}})\!=\!\big[e^{j\frac{2\pi}{\lambda}(\boldsymbol{u}_{p}^{\text{BR}})^{T}\boldsymbol{d}_{1}^{\text{R}}(\boldsymbol{r}^{\text{R}})},\dots,e^{j\frac{2\pi}{\lambda}(\boldsymbol{u}_{p}^{\text{BR}})^{T}\boldsymbol{d}_{1}^{\text{R}}(\boldsymbol{r}^{\text{R}})}\big]^{T}\!,\!\!$ (7c)
$\!\boldsymbol{t}_{k,q}^{\text{RU}}(\boldsymbol{r}^{\text{B}})\!=\!\big[e^{j\frac{2\pi}{\lambda}(\boldsymbol{u}_{k,q}^{\text{RU}})^{T}\boldsymbol{d}_{1}^{\text{R}}(\boldsymbol{r}^{\text{R}})},\dots,e^{j\frac{2\pi}{\lambda}(\boldsymbol{u}_{k,q}^{\text{RU}})^{T}\boldsymbol{d}_{N}^{\text{R}}(\boldsymbol{r}^{\text{R}})}\big]^{T}\!.\!\!$ (7d)
Similarly, for a target direction $(\vartheta_{a},\varphi_{a})$ within the target sensing area $\mathcal{A}$, the direction steering vectors for BS-target (BT) and RIS-target (RT) links can be respectively given by
$\!\boldsymbol{u}_{a}^{\text{BT}}\!\!=\!\!\left[\cos(\vartheta_{a}^{\text{BT}})\!\cos(\varphi_{a}^{\text{BT}}),\cos(\vartheta_{a}^{\text{BT}})\!\sin(\varphi_{a}^{\text{BT}}),\sin(\vartheta_{a}^{\text{BT}})\right]^{T}\!\!,\!\!$ (8a)
$\!\boldsymbol{u}_{a}^{\text{RT}}\!\!=\!\!\left[\cos(\vartheta_{a}^{\text{RT}})\!\cos(\varphi_{a}^{\text{RT}}),\cos(\vartheta_{a}^{\text{RT}})\!\sin(\varphi_{a}^{\text{RT}}),\sin(\vartheta_{a}^{\text{RT}})\right]^{T}\!\!,\!\!$ (8b)
and
$\boldsymbol{t}_{a}^{\text{BT}}(\boldsymbol{r}^{\text{B}})=\big[e^{j\frac{2\pi}{\lambda}(\boldsymbol{u}_{a}^{\text{BT}})^{T}\boldsymbol{d}_{1}^{\text{B}}(\boldsymbol{r}^{\text{B}})},\dots,e^{j\frac{2\pi}{\lambda}(\boldsymbol{u}_{a}^{\text{BT}})^{T}\boldsymbol{d}_{M}^{\text{B}}(\boldsymbol{r}^{\text{B}})}\big]^{T},$ (9a)
$\boldsymbol{t}_{a}^{\text{RT}}(\boldsymbol{r}^{\text{B}})=\big[e^{j\frac{2\pi}{\lambda}(\boldsymbol{u}_{a}^{\text{RT}})^{T}\boldsymbol{d}_{1}^{\text{R}}(\boldsymbol{r}^{\text{R}})},\dots,e^{j\frac{2\pi}{\lambda}(\boldsymbol{u}_{a}^{\text{RT}})^{T}\boldsymbol{d}_{N}^{\text{R}}(\boldsymbol{r}^{\text{R}})}\big]^{T}.$ (9b)
As shown in (7) and (9), array rotations, which generate position-dependent effects on the steering vector’s phase term, can be exploited to reshape channel responses.
II-C Rotation-Aware Multipath Channels
To capture rotation-dependent directivity in the radiation pattern of BS and RIS arrays, we include non-negative gain terms $G^{\text{B}}\left(\boldsymbol{r}^{\text{B}}\right)$ and $G^{\text{R}}\left(\boldsymbol{r}^{\text{R}}\right)$, which model directional antenna and element patterns under rotation. Let $\bar{\boldsymbol{n}}=[0,0,1]^{T}$ and
$\boldsymbol{n}(\boldsymbol{r}^{\mathrm{A}})=\boldsymbol{R}(\boldsymbol{r}^{\mathrm{A}})\bar{\boldsymbol{n}}$ (10)
be the unrotated and rotated array boresight, respectively, we adopt a practical directivity model
[…]
where $G_{0}^{\text{B}}$ and $G_{0}^{\text{R}}$ denotes the maximum gain achieved when the signal aligns with the boresight, and $p^{\text{B}}$ and $p^{\text{R}}$ are the directivity factors controlling the main-lobe width, following from directivity normalization over the full sphere. Then, with given numbers of resolvable paths $L_{k}$, $P$, and $Q_{k}$, and small-scale complex gains $a_{k,l}$, $b_{p}$, and $c_{k,q}$ that absorb path loss and fading, the BS-to-user channel $\boldsymbol{h}_{k}\left(\boldsymbol{r}^{\text{B}}\right)\in\mathbb{C}^{M}$, BS-to-RIS channel $\boldsymbol{B}\left(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}}\right)\in\mathbb{C}^{M\times N}$, and RIS-to-user channel $\boldsymbol{g}_{k}\left(\boldsymbol{r}^{\text{R}}\right)\in\mathbb{C}^{N}$ are respectively expressed as
$\boldsymbol{h}_{k}\left(\boldsymbol{r}^{\text{B}}\right)=\sum_{l=1}^{L_{k}}{a_{k,l}\sqrt{G_{k,l}^{\text{B}}\left(\boldsymbol{r}^{\text{B}}\right)}}\boldsymbol{t}_{k,l}^{\text{BU}}\left(\boldsymbol{r}^{\text{B}}\right)$ (12a)
$\boldsymbol{B}\left(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}}\right)=\sum_{p=1}^{P}{b_{p}\sqrt{G_{k,p}^{\text{B}}\left(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}}\right)G_{k,p}^{\text{R}}\left(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}}\right)}}$
$\qquad\qquad\qquad\quad\,\,\times\boldsymbol{t}_{p}^{\text{BR}}\left(\boldsymbol{r}^{\text{B}}\right)\left(\boldsymbol{t}_{p}^{\text{RB}}\left(\boldsymbol{r}^{\text{R}}\right)\right)^{H}$ (12b)
$\boldsymbol{g}_{k}\left(\boldsymbol{r}^{\text{R}}\right)=\sum_{q=1}^{Q_{k}}{c_{k,q}\sqrt{G_{k,q}^{\text{R}}\left(\boldsymbol{r}^{\text{R}}\right)}}\boldsymbol{t}_{k,l}^{\text{RU}}\left(\boldsymbol{r}^{\text{R}}\right).$ (12c)
Taking into account direct and cascaded RIS-enabled links, the effective composite channel from BS to the $k$-th user is
$\boldsymbol{f}_{k}(\boldsymbol{r}^{\text{B}}\!,\boldsymbol{r}^{\text{R}}\!,\boldsymbol{\theta})\!=\!\boldsymbol{h}_{k}(\boldsymbol{r}^{\text{B}})\!+\!\boldsymbol{B}(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}})\text{diag}(\boldsymbol{\theta})\boldsymbol{g}_{k}(\boldsymbol{r}^{\text{R}}),\!\!$ (13)
where $\boldsymbol{\theta}=[\theta_{1},\ldots,\theta_{N}]^{T}$ collects the RIS reflecting phase shift coefficients, typically being passive with $|\theta_{n}|=1$. Similarly, for sensing, define the effective steering toward a target direction $(\vartheta_{a},\varphi_{a})$ as
$\boldsymbol{f}_{\text{S},a}(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}},\boldsymbol{\theta})=$ $\sqrt{G_{a}^{\text{B}}\left(\boldsymbol{r}^{\text{B}}\right)}a_{a}\boldsymbol{t}_{a}^{\text{BT}}\left(\boldsymbol{r}^{\text{B}}\right)+\!\boldsymbol{B}(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}})\text{diag}(\boldsymbol{\theta})$
$\times\sqrt{G_{a}^{\text{R}}\left(\boldsymbol{r}^{\text{B}}\right)}c_{a}\boldsymbol{t}_{a}^{\text{RT}}(\boldsymbol{r}^{\text{B}})$ (14)
II-D Transmit Signal, Communication SINR, and Sensing MSE
The BS transmits an ISAC signal
$\boldsymbol{x}=\boldsymbol{W}_{\text{C}}\boldsymbol{s}_{\text{C}}+\boldsymbol{W}_{\text{S}}\boldsymbol{s}_{\text{S}},$ (15)
where $\boldsymbol{W}_{\text{C}}=[\boldsymbol{w}_{1},\dots,\boldsymbol{w}_{K}\boldsymbol{]}\in\mathbb{C}^{M\times K}$ collects $K$ communication beams, $\boldsymbol{W}_{\text{S}}\in\mathbb{C}^{M\times M}$ contains dedicated sensing beams, $\boldsymbol{s}_{\text{C}}=[s_{1},\dots,s_{K}]^{T}$ denotes the communication symbols, and $\boldsymbol{s}_{\text{S}}$ denotes the sensing signals, with $\mathbb{E}\left[\boldsymbol{s}_{\text{S}}\boldsymbol{s}_{\text{S}}^{H}\right]=\boldsymbol{I}_{M}$, $\boldsymbol{s}_{\text{C}}\sim\mathcal{C}\mathcal{N}\left(\boldsymbol{0},\boldsymbol{I}_{K}\right)$, and $\mathbb{E}\left[\boldsymbol{s}_{\text{S}}\boldsymbol{s}_{\text{C}}^{H}\right]=0$. Let $\boldsymbol{W}\triangleq\left[\boldsymbol{W}_{\text{C}},\boldsymbol{W}_{\text{S}}\right]$ denotes the BS ISAC transmit beamforming matrix, $\boldsymbol{s}\triangleq\left[\boldsymbol{s}_{\text{C}}^{T},\boldsymbol{s}_{\text{S}}^{T}\right]^{T}$, the received signal at the communication user $k$ is written as
$y_{k}$ $=\boldsymbol{f}_{k}^{H}(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}},\boldsymbol{\theta})\boldsymbol{Ws}$ (16a)
$=\boldsymbol{f}_{k}^{H}(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}},\boldsymbol{\theta})\boldsymbol{w}_{k}s_{k}+\sum_{i=1,i\neq k}^{K}{\boldsymbol{f}_{k}^{H}(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}},\boldsymbol{\theta})\boldsymbol{w}_{i}}x_{i}$
$\quad+\boldsymbol{f}_{k}^{H}(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}},\boldsymbol{\theta})\boldsymbol{W}_{\text{S}}\boldsymbol{S}_{\text{S}}+n_{k},$ (16b)
where $n_{k}\sim\mathcal{C}\mathcal{N}\left(0,\sigma_{k}^{2}\right)$ is the complex additive white Gaussian noise. Taking further $\boldsymbol{W}\triangleq\left[\boldsymbol{w}_{1},\dots,\boldsymbol{w}_{K+N}\right]$, where $\boldsymbol{w}_{i}$ is the $i$-th column of $\boldsymbol{W}$, the SINR of the $k$-th user is
$\text{SINR}_{k}=\frac{\big|\boldsymbol{f}_{k}^{H}\left(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}},\boldsymbol{\theta}\right)\boldsymbol{w}_{k}\big|^{2}}{\sum_{i=1,i\neq k}^{K+M}{\big|\boldsymbol{f}_{k}^{H}\left(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}},\boldsymbol{\theta}\right)\boldsymbol{w}_{i}\big|^{2}}+\sigma_{k}^{2}},$ (17)
For sensing, the transmit beam pattern toward $(\vartheta_{a},\varphi_{a})$ is
$\mathcal{P}(\vartheta_{a},\varphi_{a})$ $=\mathbb{E}(|\boldsymbol{f}_{\text{S},a}^{H}\left(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}},\boldsymbol{\theta}\right)\boldsymbol{Ws}|^{2})$ (18a)
$=\boldsymbol{f}_{\text{S},a}^{H}(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}},\boldsymbol{\theta})\boldsymbol{WW}^{H}\boldsymbol{f}_{\text{S},a}(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}},\boldsymbol{\theta}).$ (18b)
Using (13) this expands to
$\mathcal{P}(\vartheta_{a},\varphi_{a})$
$=(\boldsymbol{t}_{a}^{\text{BT}}(\boldsymbol{r}^{\text{B}}))^{H}\boldsymbol{WW}^{H}\boldsymbol{t}_{a}^{\text{BT}}(\boldsymbol{r}^{\text{B}})+\!2\sqrt{G_{a}^{\text{B}}(\boldsymbol{r}^{\text{B}})G_{a}^{\text{R}}(\boldsymbol{r}^{\text{R}})}$
$\quad\quad\times\text{Re}\{(\boldsymbol{t}_{a}^{\text{BT}}(\boldsymbol{r}^{\text{B}}))^{H}\boldsymbol{WW}^{H}\boldsymbol{B}(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}})\mathrm{diag}(\boldsymbol{\theta})\boldsymbol{t}_{a}^{\text{RT}}(\boldsymbol{r}^{\text{R}})\}$
$\quad+\sqrt{G_{a}^{\text{B}}(\boldsymbol{r}^{\text{B}})G_{a}^{\text{R}}(\boldsymbol{r}^{\text{R}})}(\boldsymbol{B}\left(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}}\right)\mathrm{diag}(\boldsymbol{\theta})\boldsymbol{t}_{a}^{\text{RT}}(\boldsymbol{r}^{\text{R}}))^{H}$
$\quad\quad\times\boldsymbol{WW}^{H}\boldsymbol{B}(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}})\mathrm{diag}(\boldsymbol{\theta})\boldsymbol{t}_{a}^{\text{RT}}(\boldsymbol{r}^{\text{R}}).$ (19)
Given a discretized grid $\{(\vartheta_{a},\varphi_{a})\}_{a=1}^{A},a\in\{1,\dots,A\}$ with $A$ sampled points in total, which samples the azimuth and elevation angular domain over the target sensing area $\mathcal{A}$, along with the desired beam pattern $\mathcal{P}_{\text{d}}(\vartheta_{a},\varphi_{a})$, the mean squared error (MSE) between $\mathcal{P}$ and $\mathcal{P}_{\text{d}}$ to evaluate the similarity of the sensing beam pattern is defined as []
$\text{MSE}\!=\!\frac{1}{A}\!\sum_{a=1}^{A}\!{\left|\mathcal{P}(\vartheta_{a},\varphi_{a};\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}},\boldsymbol{\theta},\boldsymbol{W})\!-\!\mathcal{P}_{\text{d}}(\vartheta_{a},\varphi_{a})\right|^{2}}\!.\!\!\!$ (20)
A typical ideal desired beam pattern is given by
[…]
where $\mathcal{T}$ is the set of focused grid points in the region $\mathcal{A}$.
II-E Problem Formulation
Accordingly, we jointly optimize the BS and RIS rotations, the RIS phase shifts, and the BS transmit beamforming to maximize the sum-rate performance for communication with the guaranteed sensing requirement on the beam pattern MSE
$\text{(P1)}:\underset{\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}},\boldsymbol{\theta},\boldsymbol{W}}{\max}$ $\sum_{k=1}^{K}{\log_{2}\left(1+\text{SINR}_{k}\right)}$ (22a)
s.t. $\text{MSE}\leq\eta_{\text{th}}$ (22b)
$\|\boldsymbol{W}\|_{F}^{2}=\text{Tr}(\boldsymbol{WW}^{H})\leq P_{\text{B}},$ (22c)
$|\theta_{n}|=1,\forall n\in\mathcal{N},$ (22d)
$\zeta_{i}^{\text{B}}\!\in\left[\zeta_{i,\min}^{\text{B}},\zeta_{i,\max}^{\text{B}}\right],i\in\{x,y,z\},$ (22e)
$\zeta_{i}^{\text{R}}\!\in\left[\zeta_{i,\min}^{\text{R}},\zeta_{i,\max}^{\text{R}}\right],i\in\{x,y,z\},$ (22f)
where $\eta_{\text{th}}$ in (22b) is the maximum allowed sensing MSE, $P_{\mathrm{B}}$ in (22c) is the BS power budget, the unit-modulus constraint (23d) enforces a passive RIS, and the box constraints (22e) and (22f) capture mechanical limits of the array rotations at BS and RIS, respectively.
From the formulated problem, we observe that the rotation variables $(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}})$ provide a geometric control that affects the steering vectors and directive gain, thereby complementing the electromagnetic control offered by active and passive beamforming at the RIS and BS arrays; consequently, jointly optimizing these blocks yields a significant DoF for low-altitude ISAC. However, (P1) is highly non-convex due to unit-modulus constraints, rotation-induced nonlinearities in the steering vector and directive gain, and the coupling among $\boldsymbol{r}^{\text{B}}$, $\boldsymbol{r}^{\text{R}}$, $\boldsymbol{\theta}$, and $\boldsymbol{W}$ in both SINR and MSE expressions.
III Proposed Algorithm
In this section, we convert (P1) to a penalized form and solve it by resorting to AO technique across three blocks: BS transmit beamforming $\boldsymbol{W}$, RIS passive beamforming $\boldsymbol{\theta}$, and array rotations $(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}})$. The three subproblems are handled by a quadratic transform (QT) and a majorization-minimization (MM) step for $\boldsymbol{W}$, Riemannian optimization for $\boldsymbol{\theta}$, and the projection gradient-based method for $(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}})$, respectively.
III-A Penalty Method and AO Framework
The beam pattern similarity constraints (22b), with a high-order non-convex MSE expression, limit the feasible region and hinder the solution of (P1). To handle this difficulty, we incorporate (22b) into the objective function as a hinge penalty term, obtaining a more tractable reformulated problem
$\text{(P1.1)}:\underset{\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}},\boldsymbol{\theta},\boldsymbol{W}}{\max}$ $\sum_{k=1}^{K}{\log_{2}\left(1+\text{SINR}_{k}\left(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}},\boldsymbol{\theta},\boldsymbol{W}\right)\right)}$
$\!\!\!\!-\rho\max\!\left\{\text{MSE}\left(\boldsymbol{r}^{\text{B}}\!,\boldsymbol{r}^{\text{R}}\!,\boldsymbol{\theta}\!,\boldsymbol{W}\right)\!-\!\eta_{\text{th}},0\right\}\!\!\!$ (23)
s.t. $\left(22\text{c}\right),\left(22\text{d}\right),\left(22\text{e}\right),\left(22\text{f}\right)$
where the adjustable penalty parameter $\rho\geq 0$ penalizes violations of (22b) with the hinge penalty function $\max\!\left\{\text{MSE}\!-\!\eta_{\text{th}},0\right\}$. When $\rho\rightarrow\infty$, any solution $\{(\boldsymbol{r}^{\text{B}})^{*},(\boldsymbol{r}^{\text{R}})^{*},\boldsymbol{\theta}^{*},\boldsymbol{W}^{*}\}$ of (P1.1) should meet the condition $\text{MSE}\leq\eta_{\text{th}}$, establishing the equivalence between (P1) and (P1.1). Since setting $\rho$ large at the beginning renders this approach ineffective, we initialize $\rho$ by a small value to find a good starting point, solve problem (P1.1) iteratively with gradually increased $\rho$, until $\text{MSE}\leq\eta_{\text{th}}$ holds and the objective function reaches the convergence criteria. $\rho$ increases only when $\mathrm{MSE}>\eta_{\text{th}}$ persists; once the constraint is satisfied, $\rho$ is kept fixed to avoid unnecessary rate loss.
Even with fixed $\rho$, due to the non-concave objective function and the non-convex constraints with closely coupled variables, (P1.1) remains highly non-convex and challenging to solve directly. Nevertheless, observing that the problem becomes more tractable when optimizing one variable block at a time, we resort to the AO method: at each outer iteration, we cyclically optimize the transmit beamforming $\boldsymbol{W}$, the RIS phase shifts $\boldsymbol{\theta}$, and the array rotations $\{\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}}\}$, while keeping the others fixed, as detailed in the following subsections.
III-B QT- and MM-based Transmit Beamforming
With fixed $(\boldsymbol{\theta},\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}})$, the subproblem optimizing $\boldsymbol{W}$ reads
$\text{(P2)}:\underset{\boldsymbol{W}}{\max}\,\,$ $\sum_{k=1}^{K}{\log_{2}\left(1+\text{SINR}_{k}(\boldsymbol{W})\right)}$
$-\rho\max\left\{\text{MSE}(\boldsymbol{W})-\eta_{\text{th}},0\right\}$ (24)
s.t. $\left(22\text{c}\right),$
where the SINR and MSE expressions reduce to
$\text{SINR}_{k}\left(\boldsymbol{W}\right)=\frac{|\boldsymbol{f}_{k}^{H}\boldsymbol{w}_{k}|^{2}}{\sum_{i=1,i\neq k}^{K+M}{|\boldsymbol{f}_{k}^{H}\boldsymbol{w}_{i}|^{2}}+\sigma_{k}^{2}},$ (25a)
$\text{MSE}\left(\boldsymbol{W}\right)=\frac{1}{A}\sum_{a=1}^{A}{\left|\|\boldsymbol{W}^{H}\boldsymbol{f}_{\text{S},a}\|_{2}^{2}-\mathcal{P}_{\text{d},a}\right|^{2}}.$ (25b)
The power constraint (22c) is convex, and the non-concavity of (24) originates from the fractional SINR and fourth-order MSE expressions, which are sequentially tackled as follows.
III-B1 Sensing MSE term with MM-based second-order upper bound
For brevity, let
$g_{a}\left(\boldsymbol{W}\right)$ $\triangleq\big(\lVert\boldsymbol{W}^{H}\boldsymbol{f}_{\text{S},a}\rVert_{2}^{2}-\mathcal{P}_{\text{d},a}\big)^{2},$ (26a)
$=\big(\text{Tr}(\boldsymbol{W}^{H}\boldsymbol{S}_{a}\boldsymbol{W})-\mathcal{P}_{\text{d},a}\big)^{2},$ (26b)
with the rank-one positive semi-definite matrix defined by
$\boldsymbol{S}_{a}\triangleq\boldsymbol{f}_{\text{S},a}\boldsymbol{f}_{\text{S},a}^{H}\succeq\boldsymbol{0},$ (27)
we have the gradient of (26) being
$\boldsymbol{G}_{a}\left(\boldsymbol{W}\right)$ $\triangleq\nabla g_{a}\left(\boldsymbol{W}\right)$ (28a)
$=4\big(\text{Tr}(\boldsymbol{W}^{H}\boldsymbol{S}_{a}\boldsymbol{W})-\mathcal{P}_{\text{d},a}\big)\boldsymbol{S}_{a}\boldsymbol{W}.$ (28b)
At the current iterate $\boldsymbol{W}^{(t)}$, we apply the second-order Taylor expansion (also known as the Descent Lemma [26]) to construct a convex upper-bound surrogate function for $g_{a}(\boldsymbol{W})$ as
$g_{a}\left(\boldsymbol{W}\right)$ $\leq g_{a}\big(\boldsymbol{W}^{\left(t\right)}\big)+\text{Re}\big\{\text{Tr}\big((\boldsymbol{W}-\boldsymbol{W}^{(t)})^{H}\boldsymbol{G}_{a}(\boldsymbol{W}^{(t)})\big)\big\}$
$\quad+\frac{L_{a}^{\text{lip}}}{2}\big\|\boldsymbol{W}-\boldsymbol{W}^{(t)}\big\|_{F}^{2}$ (29a)
$\triangleq\mathcal{U}_{a}\big(\boldsymbol{W}\mid\boldsymbol{W}^{(t)}\big),$ (29b)
where $L_{a}^{\text{lip}}$ is any Lipschitz constant on the Frobenius ball $\{\boldsymbol{W}:\|\boldsymbol{W}\|_{F}^{2}\leq P_{\mathrm{B}}\}$ for $\nabla g_{a}(\boldsymbol{W})$. Using matrix-norm inequalities, one obtains a choice for a global Lipschitz constant
$L_{a}^{\text{lip}}=12P_{\text{B}}\lVert\boldsymbol{S}_{a}\rVert_{2}^{2}\,+4\mathcal{P}_{\text{d},a}\lVert\boldsymbol{S}_{a}\rVert_{2},$ (30)
with $\|\boldsymbol{S}_{a}\|_{2}=\|\boldsymbol{f}_{\mathrm{S},a}\|_{2}^{2}$.
Proof: For any two points $\boldsymbol{W}_{1}$ and $\boldsymbol{W}_{2}$ on $g_{a}(\boldsymbol{W})$, we have
$\lVert\nabla g_{a}\left(\boldsymbol{W}_{1}\right)-\nabla g_{a}\left(\boldsymbol{W}_{2}\right)\rVert$
$=4\big\|(\text{Tr}(\boldsymbol{W}_{1}^{H}\boldsymbol{S}_{a}\boldsymbol{W}_{1})-\mathcal{P}_{\text{d},a})\boldsymbol{S}_{a}(\boldsymbol{W}_{1}-\boldsymbol{W}_{2})$
$\quad\quad+(\text{Tr}(\boldsymbol{W}_{1}^{H}\boldsymbol{S}_{a}\boldsymbol{W}_{1})-\text{Tr}(\boldsymbol{W}_{2}^{H}\boldsymbol{S}_{a}\boldsymbol{W}_{2}))\boldsymbol{S}_{a}\boldsymbol{W}_{2}\big\|$ (31a)
$\leq 4\lVert\boldsymbol{S}_{a}\rVert_{2}\big(|\text{Tr}(\boldsymbol{W}_{1}^{H}\boldsymbol{S}_{a}\boldsymbol{W}_{1})-\mathcal{P}_{\text{d},a}|\lVert\boldsymbol{W}_{1}-\boldsymbol{W}_{2}\rVert_{F}$
$\quad\quad+|\text{Tr}(\boldsymbol{W}_{1}^{H}\boldsymbol{S}_{a}\boldsymbol{W}_{1})\!-\!\text{Tr}(\boldsymbol{W}_{2}^{H}\boldsymbol{S}_{a}\boldsymbol{W}_{2})|\lVert\boldsymbol{W}_{2}\rVert_{F}\big).\!\!\!$ (31b)
Then, according to the following inequalities for terms in (31b)
$|\text{Tr}(\boldsymbol{W}_{1}^{H}\boldsymbol{S}_{a}\boldsymbol{W}_{1}t)-\text{Tr}(\boldsymbol{W}_{2}^{H}\boldsymbol{S}_{a}\boldsymbol{W}_{2})|$
$=|\text{Tr}((\boldsymbol{W}_{1}-\boldsymbol{W}_{2})^{H}\boldsymbol{S}_{a}\boldsymbol{W}_{1})-\text{Tr}(\boldsymbol{W}_{2}^{H}\boldsymbol{S}_{a}(\boldsymbol{W}_{1}-\boldsymbol{W}_{2}))|$
$\leq\lVert\boldsymbol{W}_{1}-\boldsymbol{W}_{2}\rVert_{F}\lVert\boldsymbol{S}_{a}\rVert_{2}(\lVert\boldsymbol{W}_{1}\rVert_{F}+\lVert\boldsymbol{W}_{2}\rVert_{F}),$ (32)
$|\text{Tr}(\boldsymbol{W}_{1}^{H}\boldsymbol{S}_{a}\boldsymbol{W}_{1})-\mathcal{P}_{\text{d},a}|\leq\lVert\boldsymbol{S}_{a}\rVert_{2}\lVert\boldsymbol{W}_{1}\rVert_{F}^{2}+\mathcal{P}_{\text{d},a},$ (33)
$\lVert\boldsymbol{W}_{1}\rVert_{F},\lVert\boldsymbol{W}_{2}\rVert_{F}\leq\sqrt{P_{B}},$ (34)
we substitute (32)-(34) into (31) and finally obtain
$\lVert\nabla g_{a}\left(\boldsymbol{W}_{1}\right)-\nabla g_{a}\left(\boldsymbol{W}_{2}\right)\rVert$
$\leq 4\lVert\boldsymbol{S}_{a}\rVert_{2}\left(P_{B}\lVert\boldsymbol{S}_{a}\rVert_{2}+\mathcal{P}_{\text{d},a}\right)\lVert\boldsymbol{W}_{1}-\boldsymbol{W}_{2}\rVert_{F}$
$\quad+4\lVert\boldsymbol{S}_{a}\rVert_{2}(2\sqrt{P_{B}}\lVert\boldsymbol{S}_{a}\rVert_{2}\lVert\boldsymbol{W}_{1}-\boldsymbol{W}_{2}\rVert_{F})\sqrt{P_{B}}$
$=\left(12P_{B}\lVert\boldsymbol{S}_{a}\rVert_{2}^{2}+4\lVert\boldsymbol{S}_{a}\rVert_{2}\mathcal{P}_{\text{d},a}\right)\lVert\boldsymbol{W}_{1}-\boldsymbol{W}_{2}\rVert_{F},$ (35)
which is exactly the definition of the Lipschitz constant.    $\blacksquare$
Consequently, replacing each $g_{a}(\boldsymbol{W})$ by its upper bound yields a concave penalization term $\rho\max\{\mathcal{U}_{a}(\boldsymbol{W}\mid\boldsymbol{W}^{(t)})-\!\eta_{\text{th}},0\}$. At each iteration, (29) provides a tight lower bound on $\boldsymbol{W}^{(t)}$ to maximize the objective function (24), guaranteeing non-decreasing penalized objective values in updates to $\boldsymbol{W}$ with a standard MM step.
III-B2 Communication SINR term with Lagrangian dual transform (LDT) and QT for fractions
Introducing auxiliary real-valued variables $\{\mu_{k}>0,\forall k\}$, we remove the fractional term from the logarithmic function using the LDT [27]
$\log_{2}\Big(1+\frac{|\boldsymbol{f}_{k}^{H}\boldsymbol{w}_{k}|^{2}}{\sum_{i=1,i\neq k}^{K+M}{|\boldsymbol{f}_{k}^{H}\boldsymbol{w}_{i}|^{2}}+\sigma_{k}^{2}}\Big)$
$=\underset{\mu_{k}\geq 0}{\max}\,\log_{2}(1\!+\!\mu_{k})\!-\!\frac{\mu_{k}}{\ln 2}\!+\!\frac{1+\mu_{k}}{\ln 2}\frac{|\boldsymbol{f}_{k}^{H}\boldsymbol{w}_{k}|^{2}}{\sum_{i=1}^{K+M}{|\boldsymbol{f}_{k}^{H}\boldsymbol{w}_{i}|^{2}}\!+\!\sigma_{k}^{2}}.$ (36)
Next, introducing complex-valued variables $\{y_{k},\forall k\}$, we apply QT to transform the fractional SINR term into
$\frac{|\boldsymbol{f}_{k}^{H}\boldsymbol{w}_{k}|^{2}}{\sum_{i=1}^{K+M}{|\boldsymbol{f}_{k}^{H}\boldsymbol{w}_{i}|^{2}}+\sigma_{k}^{2}}$
$=\underset{y_{k}\in\mathbb{C}}{\max}\,\,2\text{Re}\{y_{k}^{*}\boldsymbol{f}_{k}^{H}\boldsymbol{w}_{k}\}\!-\!\left|y_{k}\right|^{2}\Big(\!\sum_{i=1}^{K+M}\!{|\boldsymbol{f}_{k}^{H}\boldsymbol{w}_{i}|^{2}}\!+\!\sigma_{k}^{2}\Big).\!\!$ (37)
These auxiliary variables can be updated in closed form
$\mu_{k}^{*}=\frac{|\boldsymbol{f}_{k}^{H}\boldsymbol{w}_{k}|^{2}}{\sum_{i=1,i\neq k}^{K+M}{|\boldsymbol{f}_{k}^{H}\boldsymbol{w}_{i}|^{2}}+\sigma_{k}^{2}},$ (38)
$y_{k}^{*}=\frac{\boldsymbol{f}_{k}^{H}\boldsymbol{w}_{k}}{\sum_{i=1}^{K+M}{|\boldsymbol{f}_{k}^{H}\boldsymbol{w}_{i}|^{2}}+\sigma_{k}^{2}}.$ (39)
Consequently, the communication sum rate is equivalently transformed into a non-fractional quadratic function of $\boldsymbol{W}$
$\mathcal{Y}\left(\boldsymbol{W}\mid\mu_{k},y_{k}\right)$
$\!=\!\frac{1\!\!+\!\mu_{k}}{\ln 2}\Big(2\text{Re}\{y_{k}^{*}\boldsymbol{f}_{k}^{H}\boldsymbol{w}_{k}\}\!-\!\left|y_{k}\right|^{2}\!\Big(\!\sum_{i=1}^{K+M}{\!\!}|\boldsymbol{f}_{k}^{H}\boldsymbol{w}_{i}|^{2}\!\!+\!\sigma_{k}^{2}\Big)\!\Big).\!\!\!$ (40)
III-B3 Closed form solution for the QCQP
With $\{\mu_{k},y_{k}\}$ fixed and $g_{a}$ replaced by $\mathcal{U}_{a}$, (P2) reduces to the convex quadratic constrained quadratic program (QCQP) stated as below
$\text{(P2.1):}\underset{\boldsymbol{W}}{\max}\,$ $\sum_{k=1}^{K}{\mathcal{Y}\left(\boldsymbol{W}\mid\mu_{k},y_{k}\right)}-\rho\max\{\mathcal{U}_{a}(\boldsymbol{W}\mid\boldsymbol{W}^{(t)}),0\}$ (41)
s.t. $\left(22\text{c}\right).$
In general, (P2.1) can be solved using convex optimization tools. However, we derive an efficient solution based on Lagrange duality and the Karush-Kuhn-Tucker (KKT) conditions. Specifically, if the MSE threshold is violated at the current $\boldsymbol{W}^{(t)}$, the objective function (41) can be compactly rewritten as a summation with a linear and a quadratic term
$\underset{\boldsymbol{W}}{\max}\,$ $-\text{tr}(\boldsymbol{W}^{H}\boldsymbol{QW})\,\,+2\text{Re}\{\text{Tr}(\boldsymbol{P}^{H}\boldsymbol{W})\}\quad\text{s.t.}\,\,\text{(22c)},$ (42)
where $\boldsymbol{Q}$ and $\boldsymbol{P}$ are defined by
$\boldsymbol{Q}=\boldsymbol{Q}_{\text{C}}+\rho\boldsymbol{Q}_{\text{S}},\!\!$ (43a)
$\boldsymbol{P}=\boldsymbol{P}_{C}+\rho\boldsymbol{P}_{\text{S}},$ (43b)
with equivalent expressions from (29) and (40) as
$\boldsymbol{Q}_{C}=\frac{1}{\ln 2}\sum_{k=1}^{K}{\left(1+\mu_{k}\right)\left|y_{k}\right|^{2}\boldsymbol{f}_{k}\boldsymbol{f}_{k}^{H}},$ (44a)
$\boldsymbol{Q}_{\text{S}}=\frac{1}{2A}\sum_{a=1}^{A}{L_{a}^{\text{lip}}}\boldsymbol{I}_{M+K},$ (44b)
$\boldsymbol{P}_{C}=\frac{1}{\ln 2}\left[\left(1+\mu_{1}\right)y_{1}\boldsymbol{f}_{1},\dots,\left(1+\mu_{K}\right)y_{K}\boldsymbol{f}_{K}\right],$ (44c)
$\boldsymbol{P}_{\text{S}}=\frac{1}{2A}\sum_{a=1}^{A}{\boldsymbol{G}_{a}\left(\boldsymbol{W}^{\left(t\right)}\right)+L_{a}^{\text{lip}}\boldsymbol{W}^{\left(t\right)}}.$ (44d)
The corresponding Lagrangian with a multiplier $\nu>0$ is
$\mathcal{L}(\boldsymbol{W},\nu)=$ $-\text{tr}(\boldsymbol{W}^{H}\boldsymbol{QW})+2\text{Re}\{\text{tr}(\boldsymbol{P}^{H}\boldsymbol{W})t\}$
$-\nu(\lVert\boldsymbol{W}\rVert_{F}^{2}-P_{\text{B}}),$ (45)
with KKT conditions for stationarity, complementary slackness, primal feasibility, and dual feasibility, respectively, being
$-\boldsymbol{QW}^{*}+\boldsymbol{P}-\nu^{*}\boldsymbol{I}=0,$ (46a)
$\nu^{*}\left(\lVert\boldsymbol{W}^{*}\rVert_{F}^{2}-P_{\text{B}}\right)=0,$ (46b)
$\lVert\boldsymbol{W}^{*}\rVert_{F}^{2}\leq P_{\text{B}},$ (46c)
$\nu^{*}\geq 0,$ (46d)
which yields the water-filling-like solution
$\boldsymbol{W}\left(\nu\right)=\left(\boldsymbol{Q}+\nu\boldsymbol{I}\right)^{-1}\boldsymbol{P},\nu\geq 0,$ (47)
where $\nu$ is found by bisection search to satisfy the power constraint (22c) with equality. Since $\nu$ enters as a scalar shift on the identity, we precompute the eigenvalue decomposition (EVD) $\boldsymbol{Q}=\boldsymbol{U\varLambda U}^{H}$ with $\boldsymbol{\varLambda}=\text{diag}(\lambda_{1},\dots,\lambda_{M})$ and $\lambda_{m}$ being the eigenvalue; thus, for $\nu>-\min\{\lambda_{m}\}_{m=1}^{M}$, we have
$\boldsymbol{W}\left(\nu\right)$ $=\boldsymbol{U}\left(\boldsymbol{\varLambda}+\nu\boldsymbol{I}\right)^{-1}\boldsymbol{U}^{H}\boldsymbol{P}$ (48a)
$=\boldsymbol{U}\mathrm{diag}\left(\left[\frac{1}{\lambda_{1}+\nu},\dots,\frac{1}{\lambda_{M}+\nu}\right]\right)\boldsymbol{U}^{H}\boldsymbol{P},$ (48b)
so each bisection step costs only a diagonal scaling, without refactorization, and evaluates
$\lVert\boldsymbol{W}\left(\nu\right)\rVert_{F}^{2}=\sum_{i=1}^{M}{\frac{\lVert[\boldsymbol{U}^{H}\boldsymbol{P}]_{i,:}\rVert_{2}^{2}}{(\lambda_{i}+\nu)^{2}}},$ (49)
where $\cdot]_{i,:}$ denote the $i$-th row of a matrix. Furthermore, if the MSE threshold is satisfied in current $\boldsymbol{W}^{(t)}$, we let $\boldsymbol{Q}=\boldsymbol{Q}_{\text{C}}$ and $\boldsymbol{P}=\boldsymbol{P}_{C}$ and perform the same procedures in (43)-(49).
III-B4 Overall algorithm and complexity analysis for subproblem 1
The overall low-complexity algorithm for solving (P2) is summarized in Algorithm 1. Each inner iteration performs one EVD of an $M{\times}M$ Hermitian matrix at cost $\mathcal{O}(M^{3})$ and a scalar bisection whose per-step cost is $\mathcal{O}(M(K+M))$ to calculate $\|\boldsymbol{W}(\nu)\|_{F}^{2}$; the number of bisection steps grows only logarithmically with precision. The MM construction provides a tight lower bound of the penalized objective; therefore, the penalized objective in (P2) is non-decreasing over $\boldsymbol{W}$-updates. Under the power ball, the sequence $\{\boldsymbol{W}^{(t)}\}$ admits limit points that satisfy the first-order stationarity conditions of the surrogate built from MM and QT.
Algorithm 1 Low-complexity transmit beamforming based on QT, $M$, and bisection search for (P2)
1:  Input: $\{\boldsymbol{f}_{k}\}_{k=1}^{K}$, $\{\boldsymbol{f}_{\mathrm{S},a}\}_{a=1}^{A}$, $\{\mathcal{P}_{\mathrm{d},a}\}$, $P_{\mathrm{B}}$, $\rho$, $\eta_{\mathrm{th}}$.
2:  Initialization: $\boldsymbol{W}^{(0)}$, using zero forcing in the first outer iterate, or using previous outer iterate results; Set $t=0$.
3:  repeat
4:   Compute $\|\boldsymbol{S}_{a}\|_{2}^{2}$ via (27), $L_{a}^{\text{lip}}$ via (30), $\boldsymbol{G}_{a}(\boldsymbol{W}^{(t)})$ via (28), and $\mathcal{U}(\boldsymbol{W}\mid\boldsymbol{W}^{(t)})$ via (29).
5:   Build $\boldsymbol{Q}_{\mathrm{C}},\boldsymbol{P}_{\mathrm{C}}$ from (44); if $\mathcal{U}(\boldsymbol{W}^{(t)}\mid\boldsymbol{W}^{(t)})>\eta_{\mathrm{th}}$, also build $\boldsymbol{Q}_{\mathrm{S}},\boldsymbol{P}_{\mathrm{S}}$, else set them to zero. Set $\boldsymbol{Q}$ and $\boldsymbol{P}$ via (43).
6:   Compute EVD $\boldsymbol{Q}=\boldsymbol{U}\boldsymbol{\Lambda}\boldsymbol{U}^{H}$. Find $\nu^{\star}$ by bisection using (45); set $\boldsymbol{W}^{(t{+}1)}=\boldsymbol{W}(\nu^{\star})$ via (48).
7:   Update auxiliaries $\mu_{k}^{(t{+}1)}$ by (38) and $y_{k}^{(t{+}1)}$ by(39).
8:   $t\leftarrow t{+}1$.
9:  until the relative increment on the objective value of (42) or $\|\boldsymbol{W}^{(t)}{-}\boldsymbol{W}^{(t-1)}\|_{F}/\|\boldsymbol{W}^{(t-1)}\|_{F}$ below a predetermined threshold $\varepsilon$, or $t$ hits the maximum number of iterations $t_{\max}$.
10:  Output: $\boldsymbol{W}^{\star}=\boldsymbol{W}^{(t)}$.
III-C Subproblem 2: Passive Beamforming
Given fixed $\boldsymbol{W}$ and $(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}})$, we optimize $\boldsymbol{\theta}$ via
$\text{(P3)}:\underset{\boldsymbol{\theta}}{\max}\,\,$ $\sum_{k=1}^{K}{\log_{2}\left(1+\text{SINR}_{k}(\boldsymbol{\theta})\right)}$
$-\rho\max\left\{\text{MSE}(\boldsymbol{\theta})-\eta_{\text{th}},0\right\}$ (50)
s.t. $\left(22\text{d}\right).$
The unique challenge in RIS passive beamforming is the unit-modulus constraint (22d). Nevertheless, it inherently defines a Riemannian manifold, known as the complex circle manifold
$\mathcal{R}_{\boldsymbol{\theta}}=\left\{\boldsymbol{\theta}\in\mathbb{C}^{N}:|\theta_{n}|=1,\forall n\in\mathcal{N}\right\},$ (51)
which enables the use of manifold optimization techniques to efficiently exploit the smoothness of objective functions and the structure of manifold feasible regions.
III-C1 Problem reformulation
First, we rewrite expressions in (50) with respect to $\boldsymbol{\theta}$, and recast an unconstrained manifold optimization problem. Specifically, for the $a$-th angular sample, $\mathcal{P}_{a}(\vartheta,\varphi;\boldsymbol{\theta})$ in (18) expands as a quadratic form
$\mathcal{P}_{a}(\vartheta,\varphi;\boldsymbol{\theta})$ $=\boldsymbol{\theta}^{H}\boldsymbol{\Xi}_{\text{S},a}\boldsymbol{\theta}+2\Re\{\boldsymbol{\xi}_{\text{S},a}^{H}\boldsymbol{\theta}\}+\xi_{\text{S},a},$ (52)
where $\boldsymbol{\Xi}_{\text{S},a}$, $\boldsymbol{\xi}_{\text{S},a}$, and $\xi_{\text{S},a}$ are computed based on (14) by
$\!\boldsymbol{\Xi}_{\mathrm{S},a}=G^{\mathrm{R}}(\boldsymbol{r}^{\mathrm{R}})\mathrm{diag}(\boldsymbol{t}_{a}^{\mathrm{RT}})\boldsymbol{B}^{H}\boldsymbol{W}\boldsymbol{W}^{H}\boldsymbol{B}\mathrm{diag}(\boldsymbol{t}_{a}^{\mathrm{RT}})^{H},\!\!$ (53a)
$\!\boldsymbol{\xi}_{\mathrm{S},a}^{H}=\!\sqrt{G^{\mathrm{B}}(\boldsymbol{r}^{\mathrm{B}})G^{\mathrm{R}}(\boldsymbol{r}^{\mathrm{R}})}(\boldsymbol{t}_{a}^{\mathrm{BT}})^{H}\boldsymbol{W}\boldsymbol{W}^{H}\boldsymbol{B}\mathrm{diag}(\boldsymbol{t}_{a}^{\mathrm{RT}}),\!\!\!$ (53b)
$\!\xi_{\mathrm{S},a}=G^{\mathrm{B}}(\boldsymbol{r}^{\mathrm{B}})(\boldsymbol{t}_{a}^{\mathrm{BT}})^{H}\boldsymbol{W}\boldsymbol{W}^{H}\boldsymbol{t}_{a}^{\mathrm{BT}}.\!\!$ (53c)
Similarly, the signal or interference term of any beam $\boldsymbol{w}_{i}$ for the $k$-th communication user in the SINR expression with effective downlink channel $\boldsymbol{f}_{k}(\boldsymbol{\theta})$ can be written as
$\big|\boldsymbol{f}_{k}^{H}(\boldsymbol{\theta})\boldsymbol{w}_{i}\big|^{2}$ $=\boldsymbol{\theta}^{H}\boldsymbol{\Xi}_{\mathrm{C},k,i}\boldsymbol{\theta}+2\Re\{\boldsymbol{\xi}_{\mathrm{C},k,i}^{H}\boldsymbol{\theta}\}+\xi_{\mathrm{C},k,i},$ (54)
with $\boldsymbol{\Xi}_{\text{C},k,i}$, $\boldsymbol{\xi}_{\text{C},k,i}$, and $\xi_{\text{C},k,i}$ computed based on (13)
$\boldsymbol{\Xi}_{\mathrm{C},k,i}$ $=\mathrm{diag}(\boldsymbol{g}_{k}^{*})\,\boldsymbol{B}^{H}\boldsymbol{w}_{i}\boldsymbol{w}_{i}^{H}\boldsymbol{B}\,\mathrm{diag}(\boldsymbol{g}_{k}),$ (55a)
$\boldsymbol{\xi}_{\mathrm{C},k,i}^{H}$ $=\boldsymbol{h}_{k}^{H}\boldsymbol{w}_{i}\boldsymbol{w}_{i}^{H}\boldsymbol{B}\,\mathrm{diag}(\boldsymbol{g}_{k}),$ (55b)
$\xi_{\mathrm{C},k,i}$ $=\boldsymbol{h}_{k}^{H}\boldsymbol{w}_{i}\boldsymbol{w}_{i}^{H}\boldsymbol{h}_{k}.$ (55c)
Define real-valued scalars $S_{a}(\boldsymbol{\theta})$, $C_{1,k}(\boldsymbol{\theta})$, and $C_{2,k}(\boldsymbol{\theta})$ as
$\!\!\!S_{a}(\boldsymbol{\theta})$ $=\boldsymbol{\theta}^{H}\boldsymbol{\Xi}_{\mathrm{S},a}\boldsymbol{\theta}+2\Re\{\boldsymbol{\xi}_{\mathrm{S},a}^{H}\boldsymbol{\theta}\}+\xi_{\mathrm{S},a}-\mathcal{P}_{\mathrm{d},a},$ (56a)
$\!\!\!C_{1,k}(\boldsymbol{\theta})$ $=\boldsymbol{\theta}^{H}\boldsymbol{\Xi}_{\mathrm{C},k,k}\boldsymbol{\theta}+2\Re\{\boldsymbol{\xi}_{\mathrm{C},k,k}^{H}\boldsymbol{\theta}\}+\xi_{\mathrm{C},k,k},$ (56b)
$\!\!\!C_{2,k}(\boldsymbol{\theta})$ $\!\!=\!\!\sum_{i=1,i\neq k}^{K+M}\!\!\!\Big(\!\boldsymbol{\theta}^{H}\boldsymbol{\Xi}_{\mathrm{C},k,i}\boldsymbol{\theta}\!+\!2\Re\{\boldsymbol{\xi}_{\mathrm{C},k,i}^{H}\boldsymbol{\theta}\!\}\!+\!\xi_{\mathrm{C},k,i}\!\Big)\!\!+\!\sigma_{k}^{2},\!\!\!$ (56c)
we can rewrite the objective function in (58) as
$F\left(\boldsymbol{\theta}\right)=$ $\sum_{k=1}^{K}{\log_{2}}\Big(1+\frac{C_{1,k}\left(\boldsymbol{\theta}\right)}{C_{2,k}\left(\boldsymbol{\theta}\right)}\Big)$
$-\rho\max\Big\{\frac{1}{A}\sum_{a=1}^{A}{|S_{a}\left(\boldsymbol{\theta}\right)|^{2}}-\eta_{\text{th}},0\Big\},$ (57)
which further arrives at a standard Riemannian unconstrained manifold optimization problem stated as
$\text{(P3.1):}\,\,\min_{\boldsymbol{\theta}\in\mathcal{R}_{\boldsymbol{\theta}}}$ $\,\,-F\left(\boldsymbol{\theta}\right).$
III-C2 RCG method
To solve (P3.1), we adopt the Riemannian conjugate gradient (RCG) method, which generalizes the classical conjugate gradient algorithm over curved spaces, as a manifold is a topological space that locally resembles the Euclidean space [28],[29]. At any point, tangent vectors give all feasible updating directions, and their set is the tangent space
$T_{\boldsymbol{\theta}}=\big\{\boldsymbol{\varsigma}\in\mathbb{C}^{N}:\;\Re\{\boldsymbol{\varsigma}\odot\boldsymbol{\theta}^{*}\}=\boldsymbol{0}\big\}.$ (58)
Based on this definition, the RCG method proceeds by i) computing the Riemannian gradient, ii) determining the conjugate search direction, and iii) applying a retraction operation to ensure that updated variables remain on the manifold.
The Riemannian gradient is a vector field on the manifold obtained by projecting the Euclidean gradient onto the tangent space. Using Wirtinger derivatives calculus for the real-valued $F(\boldsymbol{\theta})$ with all $\boldsymbol{\xi}_{\text{S},a}$ and $\boldsymbol{\xi}_{\text{C},k,i}$ being Hermitian, we obtain
$\nabla_{\boldsymbol{\theta}}S_{a}(\boldsymbol{\theta})$ $=2\boldsymbol{\Xi}_{\mathrm{S},a}\boldsymbol{\theta}+2\boldsymbol{\xi}_{\mathrm{S},a},$ (59)
$\nabla_{\boldsymbol{\theta}}C_{1,k}(\boldsymbol{\theta})$ $=2\boldsymbol{\Xi}_{\mathrm{C},k,k}\boldsymbol{\theta}+2\boldsymbol{\xi}_{\mathrm{C},k,k},$ (60)
$\nabla_{\boldsymbol{\theta}}C_{2,k}(\boldsymbol{\theta})$ $=\sum_{i=1,\,i\neq k}^{K+M}\big(2\boldsymbol{\Xi}_{\mathrm{C},k,i}\boldsymbol{\theta}+2\boldsymbol{\xi}_{\mathrm{C},k,i}\big).$ (61)
Using the chain rule, the complete Euclidean gradient expression $\nabla_{\boldsymbol{\theta}}F(\boldsymbol{\theta})$ can be computed by (62) at the top of the next page, which is suitable for the case that the MSE threshold does not satisfy in the current $\boldsymbol{\theta}^{(\jmath)}$; otherwise, omit the MSE-related terms in (62).
$\nabla_{\boldsymbol{\theta}}F\left(\boldsymbol{\theta}\right)=\frac{1}{\ln 2}\sum_{k=1}^{K}{\frac{C_{2,k}\left(\boldsymbol{\theta}\right)\,\nabla_{\boldsymbol{\theta}}C_{1,k}\left(\boldsymbol{\theta}\right)-C_{1,k}\left(\boldsymbol{\theta}\right)\,\nabla_{\boldsymbol{\theta}}C_{2,k}\left(\boldsymbol{\theta}\right)}{C_{2,k}\left(\boldsymbol{\theta}\right)\left(C_{1,k}\left(\boldsymbol{\theta}\right)+C_{2,k}\left(\boldsymbol{\theta}\right)\right)}}-\frac{\rho}{A}\sum_{a=1}^{A}{2}S_{a}\left(\boldsymbol{\theta}\right)\,\nabla_{\boldsymbol{\theta}}S_{a}\left(\boldsymbol{\theta}\right)$ (62a)
$\qquad\,\,\,\quad=\frac{2}{\ln 2}\sum_{k=1}^{K}{\frac{C_{2,k}\left(\mathbf{\Xi}_{\text{C},k,k}\boldsymbol{\theta}+\boldsymbol{\xi}_{\text{C},k,k}\right)-C_{1,k}\sum_{i=1,i\neq k}^{K}{\mathbf{\Xi}_{\text{C},k,i}\boldsymbol{\theta}+\boldsymbol{\xi}_{\text{C},k,i}}}{C_{2,k}\left(C_{1,k}+C_{2,k}\right)}}-\frac{4\rho}{A}\sum_{a=1}^{A}{S_{a}}\left(\boldsymbol{\theta}\right)\left(\mathbf{\Xi}_{\text{S},a}\boldsymbol{\theta}+\boldsymbol{\xi}_{\text{S},a}\right)$ (62b)
$\overline${\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ }
Then, the Riemannian gradient is obtained via the orthogonal projection $\mathsf{Proj}$ for the Euclidean vector $\nabla_{\boldsymbol{\theta}}F(\boldsymbol{\theta})$ onto the tangent space $T_{\boldsymbol{\theta}}$ as follows
$\mathsf{Rgrad}F\left(\boldsymbol{\theta}\right)$ $=\mathsf{Proj}_{\boldsymbol{\theta}}\left(\nabla_{\boldsymbol{\theta}}F\left(\boldsymbol{\theta}\right)\right)$ (63a)
$=\nabla_{\boldsymbol{\theta}}F\left(\boldsymbol{\theta}\right)-\text{Re}\{\nabla_{\boldsymbol{\theta}}F\left(\boldsymbol{\theta}\right)\odot\boldsymbol{\theta}^{*}\}\odot\boldsymbol{\theta},$ (63b)
where $\odot$ denotes the Hadamard product. Next, in iteration $\jmath$, the descent direction $\boldsymbol{\tau}^{(\jmath)}$ for $\boldsymbol{\theta}^{(\jmath)}$ is determined using the Polak-Ribiere formula. For brevity, the function $F(\boldsymbol{\theta}^{(\jmath)})$ is abbreviated as $F^{(\jmath)}$ in the following. Let $\nabla_{\mathcal{R}_{\boldsymbol{\theta}}}F^{(\jmath)}$ denote the Riemannian gradient in iteration $\jmath$. The conjugate gradient update coefficient $\chi^{(\jmath)}$, which adjusts the new search direction accounting for the curvature of the gradient, is
$\chi^{\left(\jmath\right)}=\frac{\left<\nabla_{\mathcal{R}_{\boldsymbol{\theta}}}F^{\left(\jmath\right)},\nabla_{\mathcal{R}_{\boldsymbol{\theta}}}F^{\left(\jmath\right)}-\nabla_{\mathcal{R}_{\boldsymbol{\theta}}}F^{\left(\jmath-1\right)}\right>}{\left<\nabla_{\mathcal{R}_{\boldsymbol{\theta}}}F^{\left(\jmath-1\right)},\nabla_{\mathcal{R}_{\boldsymbol{\theta}}}F^{\left(\jmath-1\right)}\right>},$ (64)
where $\left\langle\cdot,\cdot\right\rangle$ denotes the Riemannian metric, being Euclidean inner product here. The descent direction is then updated as
$\boldsymbol{\tau}^{(\jmath)}=-\nabla_{\mathcal{R}_{\boldsymbol{\theta}}}F^{\left(\jmath\right)}+\beta_{\boldsymbol{\theta}}^{(\jmath)}\mathsf{Proj}_{\boldsymbol{\theta}^{(\jmath)}}(\boldsymbol{\tau}^{(\jmath-1)}),$ (65)
here the projection represents a simple vector transport operation. After updating the variable by moving along the descent direction, we map the new point back onto the manifold using the retraction operation, ensuing that the updated variables $\boldsymbol{\theta}^{(\jmath+1)}$ satisfy their manifold constraints, as follows
$\boldsymbol{\theta}^{\left(\jmath+1\right)}$ $=\mathsf{Retr}_{\boldsymbol{\theta}^{\left(\jmath\right)}}(\varpi^{\left(\jmath\right)}\boldsymbol{\tau}^{\left(\jmath\right)})$ (66a)
$=\left[\frac{(\boldsymbol{\theta}^{\left(\jmath\right)}+\varpi^{\left(\jmath\right)}\boldsymbol{\tau}^{\left(\jmath\right)})_{n}}{|(\boldsymbol{\theta}^{\left(\jmath\right)}+\varpi^{\left(\jmath\right)}\boldsymbol{\tau}^{\left(\jmath\right)})_{n}|}\right],$ (66b)
where $\varpi^{(\jmath)}$ is the step size determined by an Armijo backtracking line search method, and the element-wise normalization ensures that $\boldsymbol{\theta}^{(\jmath+1)}$ lies on the complex circle manifold.
Algorithm 2 RCG for Passive RIS Beamforming (P3)
1:Input: $\{\boldsymbol{\Xi}_{\text{S},a},\boldsymbol{\xi}_{\text{S},a},\xi_{\text{S},a}\}$ and $\{\boldsymbol{\Xi}_{\text{C},k,i},\boldsymbol{\xi}_{\text{C},k,i},\xi_{\text{C},k,i}\}$.
2:Initialization: $\boldsymbol{\theta}^{(0)}$ , $\boldsymbol{\tau}^{(0)}=-\mathsf{Rgrad}\,F(\boldsymbol{\theta}^{(0)})$.
3:repeat
4:  Compute $\nabla_{\boldsymbol{\theta}}F(\boldsymbol{\theta}^{(\jmath)})$ via (62) and $\mathsf{Rgrad}F(\boldsymbol{\theta})$ via (63).
5:  Update $\chi^{(\jmath)}$ via (64) and search direction $\boldsymbol{\tau}^{(\jmath)}$ via (65).
6:  Armijo backtracking line search to get $\varpi^{(\jmath)}$.
7:  Perform retraction via (66) and obtain $\boldsymbol{\theta}^{(\jmath+1)}$.
8:  $\jmath\leftarrow\jmath+1$
9:until Convergence: $\|\nabla_{\boldsymbol{\theta}}F(\boldsymbol{\theta}^{(\jmath)})\|/\sqrt{N}\leq\varepsilon$ or $\jmath\geq\jmath_{\max}$.
10:Output: $\boldsymbol{\theta}^{\star}=\boldsymbol{\theta}^{(\jmath)}$.
III-C3 Overall algorithm and complexity analysis for subproblem 2
Algorithm 2 summarizes the manifold optimization procedure, guaranteeing convergence to a stationary point under standard regularity conditions [39]. The precomputation to build all $N\times N$ matrices $\{\boldsymbol{\varXi}_{\text{S},a}\}_{a=1}^{A}$ and $\{\boldsymbol{\varXi}_{\text{C},k,i}\}_{k\in[K],i\in[K+M]}$ scales as $\mathcal{O}(MN^{2}+MN(K+M)+(A+K(K+M))N^{2})$. For each RCG iteration, including objective and Riemannian gradient evaluations, line search with a constant number of trials, costs $\mathcal{O}((A+K(K+M))N^{2})$, since for every grid $a$ and pair $(k,i)$ one performs products of the form $\boldsymbol{\varXi}_{\text{C},k,i}\boldsymbol{\theta}$ or $\boldsymbol{\varXi}_{\text{S},a}\boldsymbol{\theta}$ and a few additional $\mathcal{O}(N^{2})$ accumulations. The total computational complexity over $T_{\mathrm{RCG}}$ iterations is thus $\mathcal{O}(T_{\text{RCG}}(A+K(K+M))N^{2})$ plus the aforementioned precomputation. These costs can be further reduced by avoiding explicit $N\times N$ formations and reusing the rank–1 decomposition for $\boldsymbol{B}^{H}\boldsymbol{W}\boldsymbol{W}^{H}\boldsymbol{B}$ in (53a).
III-D Subproblem 3: Array Rotation Angles
Given $\boldsymbol{W}$ and $\boldsymbol{\theta}$ from the previous subproblems, we jointly optimize the array rotations $(\boldsymbol{r}^{\mathrm{B}},\boldsymbol{r}^{\mathrm{R}})$ of the BS and RIS via
$\text{(P4)}:\underset{\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}}}{\max}$ $\sum_{k=1}^{K}{\log_{2}\left(1+\text{SINR}_{k}\left(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}}\right)\right)}$
$-\rho\max\{\text{MSE}\left(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}}\right)-\eta_{\text{th}},0\}\!\!\!$ (67)
s.t. $\left(22\text{e}\right),\left(22\text{f}\right)$
Due to the highly non-concave relationships involving the cosine and exponential functions in the directivity gain and steering vectors, respectively, in the objective function, and the relatively simple box constraints on the array rotation angles, we employ the projection gradient method (PGD) [30], an effective strategy for this nonlinear maximization problem with boundary value constraints.
III-D1 Gradient computation
The key step of PGD is to derive the gradient expression for the complex objective function (67). In the following, we consider the case that the rotations $(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}})^{(\ell)}$ in the current iteration violate the MSE threshold constraint. Thus, the MSE term is involved in the cost function; otherwise, the cost and gradient functions relate only to the communication sum rate. We adopt $\boldsymbol{r}^{\text{A}}$ for any array rotation variable with $\text{A}\in\{\text{B,R}\}$ for brevity as defined in (2). Further, define a six-dimensional vector $\boldsymbol{r}$ for all rotations as below
$\boldsymbol{r}\triangleq[(\boldsymbol{r}^{\text{B}})^{T},(\boldsymbol{r}^{\text{R}})^{T}]^{T}=\left[r_{x}^{\text{B}},r_{y}^{\text{B}},r_{z}^{\text{B}},r_{x}^{\text{R}},r_{y}^{\text{R}},r_{z}^{\text{R}}\right]^{T}\in\mathbb{R}^{6}.$ (68)
The gradient of (67) w.r.t. the generic rotation vector $\boldsymbol{r}$ is
$\frac{\partial F\left(\boldsymbol{r}\right)}{\partial\boldsymbol{r}}=$ $-\frac{2\rho}{A}\sum_{a=1}^{A}{\left(\mathcal{P}_{a}\left(\boldsymbol{r}\right)-\mathcal{P}_{\text{d},a}\right)\frac{\partial\mathcal{P}_{a}\left(\boldsymbol{r}\right)}{\partial\boldsymbol{r}}}$
$+\frac{1}{\ln 2}\sum_{k=1}^{K}{\frac{1}{1+\text{SINR}_{k}\left(\boldsymbol{r}\right)}\frac{\partial\text{SINR}_{k}\left(\boldsymbol{r}\right)}{\partial\boldsymbol{r}}},$ (69)
where $\frac{\partial\text{SINR}_{k}(\boldsymbol{r})}{\partial\boldsymbol{r}}$ and $\frac{\partial\mathcal{P}_{a}(\boldsymbol{r})}{\partial\boldsymbol{r}}$ are the two main computational branches.
Define the signal and interference terms in the SINR expression by $C_{1,k}\left(\boldsymbol{r}\right)=|\boldsymbol{f}_{k}^{H}\left(\boldsymbol{r}\right)\boldsymbol{w}_{k}|^{2}$ and $C_{2,k}\left(\boldsymbol{r}\right)=\sum_{i\neq k}{|}\boldsymbol{f}_{k}^{H}\left(\boldsymbol{r}\right)\boldsymbol{w}_{i}|^{2}+\sigma_{k}^{2}$, respectively. Then we have
$\frac{\partial\,\text{SINR}_{k}\left(\boldsymbol{r}\right)}{\partial\boldsymbol{r}}=\frac{\tfrac{\partial C_{1,k}\left(\boldsymbol{r}\right)}{\partial\boldsymbol{r}}C_{2,k}\left(\boldsymbol{r}\right)-C_{1,k}\left(\boldsymbol{r}\right)\tfrac{\partial C_{2,k}\left(\boldsymbol{r}\right)}{\partial\boldsymbol{r}}}{\left(C_{2,k}\left(\boldsymbol{r}\right)\right)^{2}},$ (70)
with
$\frac{\partial C_{1,k}(\boldsymbol{r})}{\partial\boldsymbol{r}}=2\text{Re}\Big\{\Big(\frac{\partial\boldsymbol{f}_{k}(\boldsymbol{r})}{\partial\boldsymbol{r}}\Big)^{H}\boldsymbol{w}_{k}\boldsymbol{w}_{k}^{H}\boldsymbol{f}_{k}(\boldsymbol{r})\Big\},$ (71a)
$\frac{\partial C_{2,k}(\boldsymbol{r})}{\partial\boldsymbol{r}}=\sum_{i\neq k}{2}\text{Re}\Big\{\Big(\frac{\partial\boldsymbol{f}_{k}(\boldsymbol{r})}{\partial\boldsymbol{r}}\Big)^{H}\boldsymbol{w}_{i}\boldsymbol{w}_{i}^{H}\boldsymbol{f}_{k}(\boldsymbol{r})\Big\}.$ (71b)
According to the effective user channel (13), we derive
$\frac{\partial\boldsymbol{f}_{k}(\boldsymbol{r})}{\partial\boldsymbol{r}^{\text{B}}}=\frac{\partial\boldsymbol{h}_{k}(\boldsymbol{r}^{\text{B}})}{\partial\boldsymbol{r}^{\text{B}}}+\frac{\partial\boldsymbol{B}(\boldsymbol{r})}{\partial\boldsymbol{r}^{\text{B}}}\mathrm{diag}(\boldsymbol{\theta})\boldsymbol{g}_{k}(\boldsymbol{r}^{\text{R}}),$ (72a)
$\frac{\partial\boldsymbol{f}_{k}(\boldsymbol{r})}{\partial\boldsymbol{r}^{\text{R}}}=\frac{\partial\boldsymbol{B}(\boldsymbol{r})}{\partial\boldsymbol{r}^{\text{R}}}\,\mathrm{diag}(\boldsymbol{\theta})\boldsymbol{g}_{k}(\boldsymbol{r}^{\text{R}})+\boldsymbol{B}(\boldsymbol{r})\mathrm{diag}(\boldsymbol{\theta})\frac{\partial\boldsymbol{g}_{k}(\boldsymbol{r}^{\text{R}})}{\partial\boldsymbol{r}^{\text{R}}},$ (72b)
where $\frac{\partial\boldsymbol{h}_{k}(\boldsymbol{r}^{\text{B}})}{\partial\boldsymbol{r}^{\text{B}}}$, $\frac{\partial\boldsymbol{g}_{k}(\boldsymbol{r}^{\text{R}})}{\partial\boldsymbol{r}^{\text{R}}}$, $\frac{\partial\boldsymbol{B}(\boldsymbol{r})}{\partial\boldsymbol{r}^{\text{R}}}$, and $\frac{\partial\boldsymbol{B}(\boldsymbol{r})}{\partial\boldsymbol{r}^{\text{B}}}$ can be computed in (88) at the top of the next page.
Then we compute the derivatives of the steering vectors. We denote $\mathrm{X}\in\{\mathrm{BT},\mathrm{RT},\mathrm{BR},\mathrm{RB},\mathrm{BU},\mathrm{RU}\},$ and omit the index for all the steering vectors with the same element-wise structure. For a generic steering vector,
$\boldsymbol{t}^{\text{X}}(\boldsymbol{r}^{\text{A}})\triangleq\big[e^{j\frac{2\pi}{\lambda}(\boldsymbol{u}^{\text{X}})^{T}\boldsymbol{d}_{1}^{\text{A}}\left(\boldsymbol{r}^{\text{A}}\right)},...,e^{j\frac{2\pi}{\lambda}\left(\boldsymbol{u}^{\text{X}}\right)^{T}\boldsymbol{d}_{J}^{\text{A}}(\boldsymbol{r}^{\text{A}})}\big]^{T},$ (73)
with element positions $\boldsymbol{d}_{v}^{\text{A}}\left(\boldsymbol{r}^{\text{A}}\right)=\boldsymbol{d}_{0}^{\text{A}}+\boldsymbol{R}\left(\boldsymbol{r}^{\text{A}}\right)\overline{\boldsymbol{d}}_{v}^{\text{A}}$ as depicted in (4) and $v\in\{m,n\}$ denoting the index of BS or RIS element, the following relationships hold based on chain rules
$\frac{\partial\boldsymbol{t}^{\text{X}}(\boldsymbol{r}^{\text{A}})}{\partial\boldsymbol{r}^{\text{A}}}=\mathrm{diag}\left(\boldsymbol{\omega}^{\text{X}}\left(\boldsymbol{r}^{\text{A}}\right)\right)\,\boldsymbol{t}^{\text{X}}\left(\boldsymbol{r}^{\text{A}}\right),$ (74a)
$\omega_{v}^{\text{X}}(\boldsymbol{r}^{\text{A}})=j\frac{2\pi}{\lambda}\left(\boldsymbol{u}^{\text{X}}\right)^{T}\frac{\partial\boldsymbol{d}_{v}^{\text{A}}\left(\boldsymbol{r}^{\text{A}}\right)}{\partial\boldsymbol{r}^{\text{A}}},$ (74b)
$\frac{\partial\boldsymbol{d}_{v}^{\text{A}}(\boldsymbol{r}^{\text{A}})}{\partial\boldsymbol{r}^{\text{A}}}\!=\!\Big[\frac{\partial\boldsymbol{R}(\boldsymbol{r}^{\text{A}})}{\partial r_{x}^{\text{A}}}\overline{\boldsymbol{d}}_{v}^{\text{A}},\frac{\partial\boldsymbol{R}(\boldsymbol{r}^{\text{A}})}{\partial r_{y}^{\text{A}}}\overline{\boldsymbol{d}}_{v}^{\text{A}},\frac{\partial\boldsymbol{R}(\boldsymbol{r}^{\text{A}})}{\partial r_{z}^{\text{A}}}\overline{\boldsymbol{d}}_{v}^{\text{A}}\Big].\!\!\!$ (74c)
The derivatives of the rotation matrix $\boldsymbol{R}\left(\boldsymbol{r}^{\text{A}}\right)=\boldsymbol{R}_{x}\left(r_{x}^{\text{A}}\right)\boldsymbol{R}_{y}\left(r_{y}^{\text{A}}\right)\boldsymbol{R}_{z}\left(r_{z}^{\text{A}}\right)$ w.r.t each rotation variable are
$\frac{\partial\boldsymbol{R}(\boldsymbol{r}^{\text{A}})}{\partial r_{x}^{\text{A}}}=\frac{\partial\boldsymbol{R}_{x}(r_{x}^{\text{A}})}{\partial r_{x}^{\text{A}}}\boldsymbol{R}_{y}(r_{y}^{\text{A}})\boldsymbol{R}_{z}(r_{z}^{\text{A}}),$ (75a)
$\frac{\partial\boldsymbol{R}(\boldsymbol{r}^{\text{A}})}{\partial r_{y}^{\text{A}}}=\boldsymbol{R}_{x}(r_{x}^{\text{A}})\frac{\partial\boldsymbol{R}_{y}(r_{y}^{\text{A}})}{\partial r_{y}^{\text{A}}}\boldsymbol{R}_{z}(r_{z}^{\text{A}}),$ (75b)
$\frac{\partial\boldsymbol{R}(\boldsymbol{r}^{\text{A}})}{\partial r_{z}^{\text{A}}}=\boldsymbol{R}_{x}(r_{x}^{\text{A}})\boldsymbol{R}_{y}(r_{y}^{\text{A}})\frac{\partial\boldsymbol{R}_{z}(r_{z}^{\text{A}})}{\partial r_{z}^{\text{A}}},$ (75c)
and according to the definition in (3), we have
$\frac{\partial\boldsymbol{R}_{x}\left(r_{x}^{\text{A}}\right)}{\partial r_{x}^{\text{A}}}=\left[\begin{matrix}0&0&0\\ 0&-\sin r_{x}^{\text{A}}&-\cos r_{x}^{\text{A}}\\ 0&\cos r_{x}^{\text{A}}&-\sin r_{x}^{\text{A}}\\ \end{matrix}\right],$ (76a)
$\frac{\partial\boldsymbol{R}_{y}\left(r_{y}^{\text{A}}\right)}{\partial r_{y}^{\text{A}}}=\left[\begin{matrix}-\sin r_{y}^{\text{A}}&0&\cos r_{y}^{\text{A}}\\ 0&0&0\\ -\cos r_{y}^{\text{A}}&0&-\sin r_{y}^{\text{A}}\\ \end{matrix}\right],$ (76b)
$\frac{\partial\boldsymbol{R}_{z}\left(r_{z}^{\text{A}}\right)}{\partial r_{z}^{\text{A}}}=\left[\begin{matrix}-\sin r_{z}^{\text{A}}&-\cos r_{z}^{\text{A}}&0\\ \cos r_{z}^{\text{A}}&-\sin r_{z}^{\text{A}}&0\\ 0&0&0\\ \end{matrix}\right].$ (76c)
For the directional gains and their derivatives, let $\bar{\boldsymbol{n}}=[0,0,1]^{T}$ and $\boldsymbol{n}(\boldsymbol{r}^{\text{A}})=\boldsymbol{R}(\boldsymbol{r}^{\text{A}})\bar{\boldsymbol{n}}$ be the array boresight. According to the practical directivity model (11), we have
$\tfrac{\partial\sqrt{G}}{\partial\boldsymbol{r}}=\frac{1}{2\sqrt{G}}\frac{\partial G}{\partial\boldsymbol{r}}.$ (77a)
$\frac{\partial G^{\text{A}}}{\partial\boldsymbol{r}^{\text{A}}}=G_{0}^{\text{A}}\,p^{\text{A}}\big(-\boldsymbol{n}^{T}\boldsymbol{u}\big)^{p^{\text{A}}-1}\Big(-\frac{\partial\boldsymbol{n}^{T}}{\partial\boldsymbol{r}^{\text{A}}}\boldsymbol{u}\Big),$ (77b)
$\frac{\partial\boldsymbol{n}^{T}}{\partial\boldsymbol{r}^{\text{A}}}=\bar{\boldsymbol{n}}^{T}\frac{\partial\boldsymbol{R}(\boldsymbol{r}^{\text{A}})}{\partial\boldsymbol{r}^{\text{A}}}.$ (77c)
At the non-differentiable boundary, subgradients are used.
$\,\,\frac{\partial\boldsymbol{h}_{k}(\boldsymbol{r}^{\text{B}})}{\partial\boldsymbol{r}^{\text{B}}}=\sum_{l=1}^{L_{k}}{a_{k,l}\bigg[\frac{\partial\sqrt{G_{k,l}^{\text{B}}(\boldsymbol{r}^{\text{B}})}}{\partial\boldsymbol{r}^{\text{B}}}\boldsymbol{t}_{k,l}^{\text{BU}}(\boldsymbol{r}^{\text{B}})+\sqrt{G_{k,l}^{\text{B}}(\boldsymbol{r}^{\text{B}})}\frac{\partial\boldsymbol{t}_{k,l}^{\text{BU}}(\boldsymbol{r}^{\text{B}})}{\partial\boldsymbol{r}^{\text{B}}}\bigg]}$ (78a)
$\,\,\frac{\partial\boldsymbol{g}_{k}(\boldsymbol{r}^{\text{R}})}{\partial\boldsymbol{r}^{\text{R}}}=\sum_{q=1}^{Q_{k}}{c_{k,q}\bigg[\frac{\partial\sqrt{G_{k,q}^{\text{R}}(\boldsymbol{r}^{\text{R}})}}{\boldsymbol{r}^{\text{R}}}\boldsymbol{t}_{k,l}^{\text{RU}}(\boldsymbol{r}^{\text{R}})+\sqrt{G_{k,q}^{\text{R}}(\boldsymbol{r}^{\text{R}})}\frac{\boldsymbol{t}_{k,l}^{\text{RU}}(\boldsymbol{r}^{\text{R}})}{\partial\boldsymbol{r}^{\text{R}}}\bigg]}.$ (78b)
$\,\,\frac{\partial\boldsymbol{B}(\boldsymbol{r})}{\partial\boldsymbol{r}^{\text{B}}}=\sum_{p=1}^{P}{b_{p}}\bigg[\frac{\partial\sqrt{G_{k,p}^{\text{B}}(\boldsymbol{r}^{\text{B}})}}{\partial\boldsymbol{r}^{\text{B}}}\sqrt{G_{k,p}^{\text{R}}(\boldsymbol{r}^{\text{R}})}\boldsymbol{t}_{p}^{\text{BR}}(\boldsymbol{r}^{\text{B}})(\boldsymbol{t}_{p}^{\text{RB}}(\boldsymbol{r}^{\text{R}}))^{H}+\sqrt{G_{k,p}^{\text{B}}(\boldsymbol{r}^{\text{B}})G_{k,p}^{\text{R}}(\boldsymbol{r}^{\text{R}})}\frac{\partial\boldsymbol{t}_{p}^{\text{BR}}(\boldsymbol{r}^{\text{B}})}{\partial\boldsymbol{r}^{\text{B}}}(\boldsymbol{t}_{p}^{\text{RB}}(\boldsymbol{r}^{\text{R}}))^{H}\bigg],$ (78c)
$\,\,\frac{\partial\boldsymbol{B}(\boldsymbol{r})}{\partial\boldsymbol{r}^{\text{R}}}=\sum_{p=1}^{P}{b_{p}}\bigg[\sqrt{G_{k,p}^{\text{B}}(\boldsymbol{r}^{\text{B}})}\frac{\partial\sqrt{G_{k,p}^{\text{R}}(\boldsymbol{r}^{\text{R}})}}{\partial\boldsymbol{r}^{\text{R}}}\boldsymbol{t}_{p}^{\text{BR}}(\boldsymbol{r}^{\text{B}}(\boldsymbol{r}^{\text{B}}))(\boldsymbol{t}_{p}^{\text{RB}}(\boldsymbol{r}^{\text{R}}))^{H}+\sqrt{G_{k,p}^{\text{B}}(\boldsymbol{r}^{\text{B}})G_{k,p}^{\text{R}}(\boldsymbol{r}^{\text{R}})}\boldsymbol{t}_{p}^{\text{BR}}(\boldsymbol{r}^{\text{B}})\Big(\frac{\partial\boldsymbol{t}_{p}^{\text{RB}}(\boldsymbol{r}^{\text{R}})}{\partial\boldsymbol{r}^{\text{R}}}\Big)^{H}\bigg],$ (78d)
$\overline${\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ }
According to the sensing beam pattern (18) and the sensing channel model (14), for the beam-pattern derivative,
$\frac{\partial\mathcal{P}_{a}}{\partial\boldsymbol{r}^{\text{A}}}=2\Re\Big\{\Big(\frac{\partial\boldsymbol{f}_{\text{S},a}(\boldsymbol{r})}{\partial\boldsymbol{r}^{\mathrm{A}}}\Big)^{H}\boldsymbol{W}\boldsymbol{W}^{H}\boldsymbol{f}_{\text{S},a}(\boldsymbol{r})\Big\},$ (79)
where $\frac{\partial\boldsymbol{f}_{\text{S},a}(\boldsymbol{r})}{\partial\boldsymbol{r}^{\mathrm{A}}}$ is computed in (80).
$\frac{\partial\boldsymbol{f}_{\text{S},a}\left(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}}\right)}{\partial\boldsymbol{r}^{\text{B}}}=a_{a}\frac{\partial\left(\sqrt{G_{a}^{\text{B}}\left(\boldsymbol{r}^{\text{B}}\right)}\boldsymbol{t}^{\text{BT}}\left(\boldsymbol{r}^{\text{B}}\right)\right)}{\partial\boldsymbol{r}^{\text{B}}}+c_{a}\frac{\partial\boldsymbol{B}\left(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}}\right)}{\partial\boldsymbol{r}^{\text{B}}}\mathrm{diag}\left(\boldsymbol{\theta}\right)\sqrt{G_{a}^{\text{R}}\left(\boldsymbol{r}^{\text{R}}\right)}\,\boldsymbol{t}^{\text{RT}}\left(\boldsymbol{r}^{\text{R}}\right),$ (80a)
$\frac{\partial\boldsymbol{f}_{\text{S},a}\left(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}}\right)}{\partial\boldsymbol{r}^{\text{R}}}=a_{a}\frac{\partial\boldsymbol{B}\left(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}}\right)}{\partial\boldsymbol{r}^{\text{R}}}\mathrm{diag}\left(\boldsymbol{\theta}\right)\sqrt{G_{a}^{\text{R}}}\boldsymbol{t}^{\text{RT}}+c_{a}\boldsymbol{B}\left(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}}\right)\mathrm{diag}\left(\boldsymbol{\theta}\right)\frac{\partial\left(\sqrt{G_{a}^{\text{R}}\left(\boldsymbol{r}^{\text{R}}\right)}\,\boldsymbol{t}^{\text{RT}}\left(\boldsymbol{r}^{\text{R}}\right)\right)}{\partial\boldsymbol{r}^{\text{R}}}$ (80b)
$\overline${\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ }
It is also convenient to gather partials into Jacobians $\boldsymbol{J}_{\text{S},a}^{(\text{A})}\triangleq\left[\frac{\partial\boldsymbol{f}_{\text{S},a}}{\partial r_{x}^{\text{A}}},\frac{\partial\boldsymbol{f}_{\text{S},a}}{\partial r_{y}^{\text{A}}},\frac{\partial\boldsymbol{f}_{\text{S},a}}{\partial r_{z}^{\text{A}}}\right]\in\mathbb{C}^{N_{t}\times 3}$ similar to (74), so that (79) compacts to
$\frac{\partial\mathcal{P}_{a}}{\partial\boldsymbol{r}^{\mathrm{A}}}=2\,\Re\!\Big\{\boldsymbol{J}_{\text{S},a}^{(\text{A})\,H}\boldsymbol{W}\boldsymbol{W}^{H}\boldsymbol{f}_{\text{S},a}\Big\}\in\mathbb{R}^{3}.$ (81)
III-D2 Projected-gradient framework for the rotation block.
Let $\mathcal{D}\triangleq\mathcal{D}_{\text{B}}\times\mathcal{D}_{\text{R}}$ be the box of angle limits in (P4) with
$\mathcal{D}_{\text{A}}=\{\boldsymbol{r}^{\text{A}}:\boldsymbol{r}_{\text{lb}}^{\text{A}}\leq\boldsymbol{r}^{\text{A}}\leq\boldsymbol{r}_{\text{ub}}^{\text{A}}\},$ (82)
and denote $\boldsymbol{r}_{\text{lb}}\!\triangleq\![(\boldsymbol{r}_{\text{lb}}^{\text{B}})^{T}\!,(\boldsymbol{r}_{\text{lb}}^{\text{R}})^{T}]^{T}$, and $\boldsymbol{r}_{\text{ub}}\!\triangleq\![(\boldsymbol{r}_{\text{ub}}^{\text{B}})^{T}\!,(\boldsymbol{r}_{\text{ub}}^{\text{R}})^{T}]^{T}$. Let $\Pi_{\mathcal{D}}(\cdot)$ be the Euclidean projection onto $\mathcal{D}$. We perform projected gradient ascent on $F$ with the projected step
$\boldsymbol{\beta}\left(\alpha\right)=\Pi_{\mathcal{D}}\left(\boldsymbol{r}+\alpha\nabla_{\boldsymbol{r}}F\left(\boldsymbol{r}\right)\right)-\boldsymbol{r},$ (83)
for any step size $\alpha>0$. Define the projected-gradient mapping $\boldsymbol{H}_{\alpha}\left(\boldsymbol{r}\right)\triangleq\frac{1}{\alpha}\boldsymbol{\beta}\left(\alpha\right)$, for any $\alpha>0$, $\boldsymbol{r}^{\star}$ is a first-order stationary point of (P4) if and only if $\boldsymbol{H}_{\alpha}\left(\boldsymbol{r}^{\star}\right)=\boldsymbol{0}$, according to the KKT condition. Hence $\|\boldsymbol{H}_{\alpha}(\boldsymbol{r})\|_{2}$ is a valid criticality measure for stopping. Because $\mathcal{D}$ is a box, the projection (83) reduces to the componentwise clipping
$\boldsymbol{\beta}^{(\ell)}(\alpha)=\text{clip}(\boldsymbol{r}^{\left(\ell\right)}+\alpha\nabla_{\boldsymbol{r}}F^{(\ell)}(\boldsymbol{r}))_{[r_{\text{lb}},r_{\text{ub}}]}-\boldsymbol{r}^{(\ell)},$ (84)
where the clip $\text{clip}\left(x\right)_{\left[y,z\right]}=\min\left\{\max\left\{x,y\right\},z\right\}$ acts elementwise. Given the previous $\left(\boldsymbol{r},F\left(\boldsymbol{r}\right),\nabla F\left(\boldsymbol{r}\right)\right)^{(\ell-1)}$, the stepsize uses Barzilai–Borwein initialization
$\alpha_{\text{BB}}=\frac{\|\boldsymbol{r}^{\left(\ell\right)}-\boldsymbol{r}^{\left(\ell-1\right)}\|^{2}}{\left(\boldsymbol{r}^{\left(\ell\right)}-\boldsymbol{r}^{\left(\ell-1\right)}\right)^{T}\left(\nabla F\left(\boldsymbol{r}^{\left(\ell\right)}\right)-\nabla F\left(\boldsymbol{r}^{\left(\ell-1\right)}\right)\right)},$ (85)
and clip to $\alpha_{\min},\alpha_{\max}]$ to ensure robustness. An Armijo backtracking is then performed until the ascent rule holds
$F\left(\boldsymbol{r}+\boldsymbol{\beta}\left(\alpha\right)\right)\geq F\left(\boldsymbol{r}\right)+c_{1}\nabla_{\boldsymbol{r}}F\left(\boldsymbol{r}\right)^{T}\boldsymbol{\beta}\left(\alpha\right),$ (86)
with $c_{1}\!\in\!(0,1)$ and the backtracking using $\alpha\!\leftarrow\!c_{2}\alpha$, $c_{2}\in(0,1)$. The iteration terminates when it meets the stopping criteria that the projected gradient norm is small, i.e., $\|\boldsymbol{H}_{\alpha}(\boldsymbol{r}^{(\ell)})\|_{2}\ \leq\ \varepsilon_{\text{pg}}$, or the relative step is small, i.e., $\frac{\|\boldsymbol{r}^{(\ell)}-\boldsymbol{r}^{(\ell-1)}\|_{2}}{\max\{1,\|\boldsymbol{r}^{(\ell-1)}\|_{2}\}}\ \leq\ \varepsilon_{\text{rel}}$. For box constraints, $\boldsymbol{H}_{\alpha}(\boldsymbol{r}^{(\ell)})$ admits the following componentwise form for $\imath$-th entry independent of $\alpha$
$\!\!\!\big[\boldsymbol{H}_{\alpha}(\boldsymbol{r}^{(\ell)})\big]_{\imath}\!=\!\!\begin{cases}[\nabla_{\boldsymbol{r}}F^{\left(\ell\right)}(\boldsymbol{r})]_{\imath},\!&\!\!\!r_{\text{lb},i}\!\!<r_{i}\!<r_{\text{ub},i}\!,\\ \max\{0,[\nabla_{\boldsymbol{r}}F^{(\ell)}(\boldsymbol{r})]_{\imath}\},\!&\!\!\!r_{\imath}=r_{\text{lb},\imath},\\ \min\{0,[\nabla_{\boldsymbol{r}}F^{(\ell)}(\boldsymbol{r})]_{\imath}\},\!&\!\!\!r_{\imath}=r_{\text{ub},\imath}.\end{cases}\!\!\!\!$ (87)
The gradient cost is dominated by Jacobians (80) of rotation-related steering vectors and derivatives (78) with matrix-chain products. Because the directional-gain model is piecewise smooth, we use subgradients at the visibility boundary and a smooth gating for stable line search.
Algorithm 3 PGD for Array Rotation Subproblem (P4)
1:  Input: fixed $(\boldsymbol{W},\boldsymbol{\theta})$, box $\mathcal{D}$, tolerances $\varepsilon_{\mathrm{pg}},\varepsilon_{\mathrm{rel}}>0$, and maximum iterations $I_{\text{PGD}}^{\text{max}}$.
2:  Initialization: initial point $\boldsymbol{r}^{(0)}\in\mathcal{D}$
3:  for $\ell=0,1,\ldots,I_{\text{PGD}}^{\text{max}}-1$ do
4:   Evaluate the cost $F(\boldsymbol{r}^{(\ell)})$ and the gradient $\nabla_{\boldsymbol{r}}F(\boldsymbol{r}^{(\ell)})$ using (67)-(81).
5:   if $\ell>0$ then
6:    Barzilai-Borwein stepsize initialization $\alpha^{(\ell)}\!\leftarrow\!\alpha_{\text{BB}}$ via (85).
7:   else
8:    $\alpha^{(\ell)}\leftarrow\alpha_{0}$.
9:   end if
10:   Find stepsize $\alpha^{(\ell)}$ via Armijo backtracking to satisfy (86).
11:   Determine projected step $\boldsymbol{\beta}^{(\ell)}(\alpha)$ via (84)
12:   Update optimization variables by $\boldsymbol{r}^{(\ell+1)}=\boldsymbol{r}^{(\ell)}+\boldsymbol{\beta}^{(\ell)}(\alpha)$.
13:   Compute projected gradient mapping $\boldsymbol{H}_{\alpha}(\boldsymbol{r}^{(\ell)})$ via (87).
14:   if $\|\boldsymbol{H}_{\alpha}(\boldsymbol{r}^{(\ell+1)})\|_{2}\leq\varepsilon_{\mathrm{pg}}$ or $\dfrac{\|\boldsymbol{r}^{(\ell+1)}-\boldsymbol{r}^{(\ell)}\|_{2}}{\max\{1,\|\boldsymbol{r}^{(\ell)}\|_{2}\}}\leq\varepsilon_{\mathrm{rel}}$ then
15:    break;
16:   end if
17:  end for
18:  Output: $\boldsymbol{r}^{\mathrm{B}}\leftarrow\boldsymbol{r}^{(\ell+1)}_{1:3}$, $\boldsymbol{r}^{\mathrm{R}}\leftarrow\boldsymbol{r}^{(\ell+1)}_{4:6}$.
III-D3 Overall algorithm under penalty-based AO framework for original problem 1
Algorithm 4 summarizes the overall projected gradient method. For the per-iteration computational complexity, directly forming $\boldsymbol{B}$ and the Jacobians as dense $M\times N$ matrices yields $\mathcal{O}(PMN)$ computations per rotation axis, and evaluating the sensing gradients over $A$ grids would add $\mathcal{O}(AMN)$. However, using the rank-$P$ factorization for $\boldsymbol{B}$ and its derivatives, all costly matrix–vector products, involving $\boldsymbol{B}$ and $\partial\boldsymbol{B}/\partial r_{\imath}^{\mathrm{A}}$, are computed without materializing $M\times N$ blocks and costing $\mathcal{O}(NP)$ plus $\mathcal{O}(MP)$. With these reuses, one PGD iteration with full gradient of $F$ has the leading cost: i) precomputation on steering vectors and their Jacobians with $\mathcal{O}\big(P(M+N)\big)$, ii) sensing and communication channel–Jacobian products $\mathcal{O}\big((A+K)P(N+M)\big)$, iii) use $\boldsymbol{W}(\boldsymbol{W}^{H}\boldsymbol{f})$ to apply $\boldsymbol{W}\boldsymbol{W}^{H}$ in (71) and (81) with $\mathcal{O}\!\big((M+K)M(A+K)\big)$. Therefore, the per-iteration complexity scales as $\mathcal{O}\big(P(M+N)(A+K+1)+(M+K)M(A+K)\big)$ with a constant coefficient affected by backtracking searches.
III-E Overall algorithm
Finally, we adopt a penalty-assisted AO-based algorithm framework over $(\boldsymbol{W},\boldsymbol{\theta},\boldsymbol{r}^{\mathrm{B}},\boldsymbol{r}^{\mathrm{R}})$, resulting in an overall algorithm with nested loops summarized in Algorithm 4, where we obtain a suboptimal solution for the original problem (P1). At the $\delta$-th outer iteration, with the other blocks fixed, we first solve (P2) by Algorithm 1 to update $\boldsymbol{W}^{(\delta+1)}$; next solve (P3) by Algorithm 2 to update $\boldsymbol{\theta}^{(\delta+1)}$; then solve (P4) by Algorithm 3 (projected-gradient over the box $\mathcal{D}$) to obtain $(\boldsymbol{r}^{\mathrm{B}},\boldsymbol{r}^{\mathrm{R}})^{(\delta+1)}$. After the three inner loops, the penalty weight is refreshed as $\rho^{(\delta+1)}\leftarrow\gamma_{\rho}\rho^{(\delta)}$ if the MSE constraint is violated (i.e., $\mathrm{MSE}>\eta_{\text{th}}$); otherwise $\rho^{(\delta+1)}=\rho^{(\delta)}$. Each inner solver uses its own line search/tolerances to guarantee ascent of the penalized objective and is warm-started from the previous iterate. The outer loop terminates when the fractional increase of the penalized objective in (P1.1) falls below $\varepsilon_{\mathrm{AO}}$ and the MSE constraint is satisfied ($\mathrm{MSE}\leq\eta_{\text{th}}$).
Algorithm 4 Penalty-assisted AO-based Algorithm for (P1)
1:  Input: $\boldsymbol{u}_{k,l}^{\text{BU}},\boldsymbol{u}_{p}^{\text{RB}},\boldsymbol{u}_{p}^{\text{BR}},\boldsymbol{u}_{k,q}^{\text{RU}},\boldsymbol{u}_{a}^{\text{BT}},\boldsymbol{u}_{a}^{\text{RT}}$, $\gamma_{\rho}\!>\!0$, and $\mathcal{P}_{\text{d}}\left(\vartheta_{a},\varphi_{a}\right)$ .
2:  Initialization: $\boldsymbol{W}^{\left(0\right)},\boldsymbol{\theta}^{\left(0\right)},\left(\boldsymbol{r}^{\text{A}},\boldsymbol{r}^{\text{B}}\right)^{\left(0\right)}$, $\rho^{(0)}$ and $\delta=0$
3:  repeat
4:   Solve subproblem (P2) via Algorithm 1 to obtain $\boldsymbol{W}^{\left(\delta+1\right)}$.
5:   Solve subproblem (P3) via Algorithm 2 to obtain $\boldsymbol{\theta}^{\left(\delta+1\right)}$.
6:   Solve subproblem (P4) via Algorithm 3 to obtain $(\boldsymbol{r}^{\text{A}}\!,\!\boldsymbol{r}^{\text{B}})^{(\delta\!+\!1)}$.
7:   if Current MSE value violates the threshold $\eta_{\text{th}}$ then
8:    $\rho^{(\delta+1)}=\gamma\rho^{(\delta)}$.
9:   else
10:    $\rho^{(\delta+1)}=\rho^{(\delta)}$.
11:   end if
12:   $\delta\leftarrow\delta+1$.
13:  until the fractional increase of the objective value of problem (P1.1) is smaller than a threshold $\varepsilon_{\text{AO}}$ and the MSE is below the threshold $\eta_{\text{th}}$.
14:  Output: $\boldsymbol{W}\leftarrow\boldsymbol{W}^{(\delta)},\boldsymbol{\theta}\leftarrow\boldsymbol{\theta}^{(\delta)},\boldsymbol{r}^{\text{A}}\leftarrow\boldsymbol{r}^{\text{A},(\delta)}$, $\boldsymbol{r}^{\text{B}}\leftarrow\boldsymbol{r}^{\text{B},(\delta)}$.
IV Simulation Results
This section presents numerical results for the proposed rotatable BS–RIS ISAC design. Unless otherwise stated, the BS is equipped with a $M=M_{x}M_{y}=2\times 2$ antennas by default with inter-element spacing $d_{x}=d_{y}=\lambda/2$ for a carrier wavelength $\lambda=0.1$ m. The RIS is modeled as a $6\times 6$ UPA with spacing $e_{x}=e_{y}=\lambda/2$, located at $(10,0,0)$ m relative to the BS. The unrotated BS and RIS boresights are aligned with the positive $z$-axis, and their orientations are parameterized by rotation vectors $\boldsymbol{r}^{\mathrm{B}}$ and $\boldsymbol{r}^{\mathrm{R}}$. The small-scale fading of all links follows a geometry-based multipath model: each BS–user and RIS–user channel comprises $L_{k}=Q_{k}=2$ paths with independently generated azimuth and elevation angles, uniformly distributed over $\pi,\pi]$ and $\pi/3,\pi/3]$, respectively, and complex Gaussian path gains with normalized variance. The cascaded BS–RIS channel contains $P=2$ paths with angles and gains drawn independently. Thermal noise at each user is modeled as circularly symmetric complex Gaussian with variance $\sigma_{k}^{2}=10^{-6}$. For sensing, the BS forms a monostatic beampattern over a two-dimensional angular grid with $A_{\mathrm{az}}=11$ azimuth and $A_{\mathrm{el}}=6$ elevation samples, uniformly spanning $\phi\in[-\pi,\pi]$ and $\vartheta\in[-\pi/4,\pi/4]$. The desired sensing pattern corresponds to two spotlight-like angular sectors, and the target threshold of sensing MSE is $\eta_{\mathrm{th}}=0.2$. The total BS transmit power is $P_{\mathrm{B}}=30$ dBm.
We adopt both half-space isotropic and directional element patterns. In the directional case, the BS and RIS element gains follow the piecewise model (11), while $G_{0,\mathrm{B}}=G_{0,\mathrm{R}}=2$ are constant in the isotopic case. For each random channel realization, the proposed alternating optimization algorithm jointly updates the BS precoder $\boldsymbol{W}$, the RIS phase vector $\boldsymbol{\theta}$, and the rotations $(\boldsymbol{r}^{\mathrm{B}},\boldsymbol{r}^{\mathrm{R}})$ so as to maximize the downlink sum rate $\sum_{k=1}^{K}\log_{2}(1+\mathrm{SINR}_{k})$ under the sensing MSE constraint. To assess the gain of spatial geometry control and RIS assistance, we compare the fully rotatable BS–RIS architecture (“Rot-BS+Rot-RIS”) with four benchmarks: (i) only the BS is rotatable while the RIS orientation is fixed (“Rot-BS+Fix-RIS”); (ii) only the RIS is rotatable (“Fix-BS+Rot-RIS”); (iii) both BS and RIS are fixed (“Fix-BS+Fix-RIS”); and two non-RIS baseline (iv) (“Rot-BS+No-RIS”) and (v) (“Fix-BS+No-RIS”), where the RIS–user link is disabled by setting the RIS–user path gains to zero. All reported curves are obtained by averaging the communication and sensing metrics over a large number of independent channel realizations.
Figure 2: Sum rate versus maximum transmit power under $p^{\text{R}}=p^{\text{B}}=0$ with $N=6\times 6$, $M=2\times 2$ and $K=2$.
Figure 3: Sum rate versus maximum transmit power under $p^{\text{R}}=p^{\text{B}}=2$ with $N=6\times 6$, $M=2\times 2$ and $K=2$.
Figs. 2 and 3 illustrate the sum rate versus the BS transmit power under the sensing beampattern MSE constraint for a $2\times 2$ transmit array, two communication users, and two sensing directions. In the half-space isotropic case (Fig. 2), all schemes benefit from increasing power, showing roughly linear growth at low–to–moderate SNR and gradual saturation when multiuser interference and the sensing constraint become dominant. Across the entire power range, the proposed joint BS–RIS rotation (“Rot-BS+Rot-RIS”) consistently achieves the highest sum rate, followed by the partially rotatable schemes. At the same time, the non-RIS baselines perform worse due to their reduced effective aperture. However, the RIS array rotations seem not to be effective in the half-space isotropic case, because rotations in this case cannot reap the directivity gains; one can say that isotropic elements avoid severe signal degradation in side-lobe directions. With a directional half-space pattern (Fig. 3, element directivity factor $2$), all curves shift upward, confirming the SNR and throughput gains brought by element directivity. At the same time, the rate gap between Rot-BS+Rot-RIS and the benchmark schemes becomes more pronounced, as the proposed design better aligns the main lobes with user directions and steers reduced-gain regions toward the sensing directions, thereby alleviating the performance penalty of the sensing constraint. This indicates that geometric adaptation, especially when both active and passive arrays are rotatable, becomes increasingly valuable in the interference-limited regime and practical BS antenna or RIS element radiation patterns.
Figure 4: Sum rate versus the number of BS antennas under $p^{\text{R}}=p^{\text{B}}=0$ with $N=6\times 6$, $M_{x}=1$ $P=30$ dBm, and $K=3$.
Figure 5: Sum rate versus the number of BS antennas under $p^{\text{R}}=p^{\text{B}}=2$ with $N=6\times 6$, $M_{x}=1$ $P=30$ dBm, and $K=3$.
Figs. 4 and 5 examine the impact of the number of transmit antennas for a $1\times M_{y}$ BS array with three users. In the isotropic case (Fig. 4), the sum rate of all schemes increases with $M_{y}$, since a larger array aperture provides stronger array gain and finer spatial resolution to separate the three users while shaping the sensing beampattern. Schemes with only BS rotation or only RIS rotation still benefit from adding antennas, but their curves remain noticeably below the fully rotatable configuration, indicating that partial geometric adaptation cannot fully unlock the potential of a large array in the presence of both multiuser interference and sensing requirements. The non-RIS baselines lie at the bottom in all $M_{y}$, highlighting the importance of the extra reflecting aperture with more spatial resources. In the directional case (Fig. 5), all curves are further elevated and the advantage of Rot-BS+Rot-RIS becomes more marked. The directivity amplifies the relative advantage of the joint BS–RIS rotation: by aligning the main lobe of the transmit array and the dominant reflecting directions of the RIS with the user angles while steering reduced-gain regions toward the sensing directions, the proposed design can more efficiently decouple communication beams from sensing beams, leading to a larger rate gain over the partially rotatable and fully fixed benchmarks. In summary, combining array scaling with geometric adaptation is effective for one-dimensional arrays in ISAC systems, especially with directional arrays with practical radiation patterns.
Figure 6: Sum rate versus the number of communication users under $p^{\text{R}}=p^{\text{B}}=2$ with $N=6\times 6$, $M=2\times 4$, and $P=22$ dBm.
Figure 7: Sum rate versus the number of communication users under $p^{\text{R}}=p^{\text{B}}=2$ with $N=6\times 6$, $M=2\times 2$, and $P=30$ dBm.
Figs. 6 investigate user scaling for a $2\times 4$ BS array with fixed transmit power $P_{\mathrm{B}}=22$ dBm. The sum rate increases with the number of users due to multiuser multiplexing gains, but the growth gradually slows as the system becomes more interference-limited and the available spatial degrees of freedom must be shared between user separation and sensing beampattern shaping. The Rot-BS+Rot-RIS scheme consistently delivers the highest sum rate, demonstrating that joint geometry control is especially beneficial in the multiuser regime, whereas partial rotation or fixed architectures suffer from stronger interference and tighter sensing constraints.
Fig. 7 illustrates the impact of the maximum BS rotation range $\zeta$ on the achievable sum rate for different RIS array rotation ranges. As $\zeta$ increases, the sum rate of all schemes improves and gradually saturates, since a larger rotation range provides more geometric degrees of freedom to steer the BS beams toward the users and away from interference directions. Across the whole range of $\zeta$, both rotatable-RIS configurations significantly outperform the fixed-IRS and no-IRS baselines, confirming that jointly adapting the BS orientation and the IRS reflection direction is highly beneficial. Moreover, the curves for “rot-IRS $90^{\circ}$” and “rot-IRS $45^{\circ}$” almost coincide once $\zeta$ exceeds $30^{\circ}$, indicating that a moderate IRS rotation capability already captures most of the gain brought by passive array rotation, while the BS rotation range remains the dominant factor in leveraging geometric flexibility.
V Conclusion
This paper studied a low-altitude ISAC system with cooperative rotatable active and passive arrays. We considered a downlink transmission where a BS equipped with an active RA serves multiple ground users and senses low-altitude targets in monostatic mode, assisted by a rotatable RIS, i.e., passive RA. To capture the impact of 3D array rotations, we developed a rotation-aware geometry-based multipath channel model in which BS–user, BS–RIS, and RIS–user links are parameterized by path-dependent 3D angles and complex gains, and array rotations jointly affect steering vectors and direction-dependent element gains. On this basis, we formulated a unified design that maximizes the multiuser sum rate under a transmit-power budget, RIS unit-modulus constraints, mechanical rotation limits, and a sensing beampattern MSE constraint. To tackle the resulting highly coupled and non-convex problem, we proposed a penalty-based AO framework that alternately updates the BS ISAC precoder, RIS phase-shift vector, and BS/RIS rotation angles. The BS precoding block is solved via a QT– and MM-aided method, leading to a tractable QCQP; the RIS block is handled on the complex-circle manifold using an RGD algorithm; and the rotation block is optimized via PGD with Barzilai–Borwein step sizes and Armijo backtracking. Extensive simulations under LAWN-inspired 3D geometries showed that the proposed jointly rotatable BS–RIS architecture achieves significant sum-rate gains over fixed or partially rotatable baselines while satisfying stringent sensing MSE requirements, with the advantages further amplified with directional antennas.

### v2

@@
 Abstract
-This paper investigates a low-altitude integrated sensing and communication (ISAC) system that leverages cooperative rotatable active and passive arrays.
-We consider a downlink scenario where a base station (BS) with an active rotatable array serves multiple communication users and senses low-altitude targets, assisted by a rotatable reconfigurable intelligent surface (RIS).
-A rotation-aware geometry-based multipath model is developed to capture the impact of three-dimensional (3D) array orientations on both steering vectors and direction-dependent element gains.
-On this basis, we formulate a new optimization problem that maximizes the downlink sum rate subject to a transmit power budget, RIS unit-modulus constraints, mechanical rotation limits, and a sensing beampattern mean-squared-error (MSE) constraint.
-To address the resulting highly non-convex problem, we propose a penalty-based alternating-optimization (AO) framework that alternately updates the BS precoder, RIS phase shifts, and BS/RIS array rotation angles.
-The three blocks are efficiently handled via a convex optimization method based on quadratic-transform (QT) and majorization-minorization (MM), Riemannian conjugate gradient (RCG) on the unit-modulus manifold, and projected gradient descent (PGD) with Barzilai–Borwein step sizes, respectively.
-Numerical results in low-altitude geometries demonstrate that the proposed jointly rotatable BS–RIS architecture achieves significant sum-rate gains over fixed or partially rotatable baselines while guaranteeing sensing requirements, especially with directional antennas and in interference-limited regimes.
+This paper investigates a low-altitude integrated sensing and communication (ISAC) system that leverages cooperatively rotatable active and passive arrays.
+We consider a downlink scenario in which a base station (BS) equipped with a three-dimensional (3D) rotatable active array serves multiple users while sensing low-altitude targets in monostatic mode, assisted by a rotatable reconfigurable intelligent surface (RIS).
+To capture the impact of array orientations, we develop a rotation-aware geometry-based multipath model in which 3D rotations jointly affect the phases of the steering vector via rotated element coordinates, and the direction-dependent element and antenna radiation gains via a practical directivity model.
+On this basis, we formulate a joint design problem that maximizes a weighted-sum utility of the downlink sum rate and the sensing beampattern similarity, where the latter is quantified by a normalized mean-squared error (NMSE).
+We introduce a scaling parameter and update it in closed form once per outer alternating-optimization (AO) iteration.
+The resulting non-convex problem is tackled via an efficient AO framework that alternately updates the BS precoder, RIS phase shifts, and array rotation angles, using a quadratic transform and majorization-minimization-based closed-form updates, Riemannian conjugate gradient on the unit-modulus manifold, and projected gradient ascent under rotation box constraints, respectively.
+Numerical results demonstrate that the proposed jointly rotatable arrays at the BS and RIS achieve a favorable rate-sensing trade-off and performance gains compared with fixed or partially rotatable baselines.
 Index Terms:
@@
 I Introduction
-Low-altitude wireless networks (LAWNs) are emerging as a key component of next-generation intelligent infrastructures, interconnecting uncrewed aerial vehicles (UAVs), low-altitude platform stations, aerial robots, and connected ground terminals in the three-dimensional (3D) airspace below a few kilometers [1]-[4].
-Typical LAWN applications, such as aerial surveillance, disaster response, smart-city sensing, and low-altitude traffic management, rely not only on reliable broadband connectivity, but also on accurate situational awareness of dynamic low-altitude objects.
-This naturally calls for integrated sensing and communication (ISAC) designs that can jointly support data transmission and high-resolution sensing within complex, rapidly changing airspace environments [5],[6].
-Compared with conventional terrestrial cellular systems, LAWNs must cope with fast-varying spatial geometries and require shaping radio signals to illuminate aerial targets with stringent size–weight–power constraints and capture their reflections, which make efficient ISAC-oriented signal processing particularly challenging.
-Multiple-input multiple-output (MIMO) technology and advanced precoding have long been recognized as key tools to exploit spatial degrees of freedom in wireless systems [7],[8].
+Low-altitude wireless networks (LAWNs) are emerging as a key component of next-generation intelligent infrastructures, interconnecting uncrewed aerial vehicles (UAVs), low-altitude platform stations, aerial robots, and connected ground terminals in the three-dimensional (3D) airspace below a few kilometers [1, 2, 3, 4].
+Typical LAWN applications, such as aerial surveillance, disaster response, smart-city sensing, and low-altitude traffic management, rely not only on reliable broadband connectivity but also on accurate situational awareness of dynamic low-altitude objects.
+This naturally calls for integrated sensing and communication (ISAC) designs that can jointly support data transmission and high-resolution sensing within complex, rapidly changing airspace environments [5, 6].
+Compared with conventional terrestrial cellular systems, LAWNs must cope with rapidly varying spatial geometries and require shaping radio signals to illuminate aerial targets while meeting stringent size–weight–power constraints and capturing their reflections, which make efficient ISAC-oriented signal processing particularly challenging.
+Multiple-input multiple-output (MIMO) technology and advanced precoding have long been recognized as key tools to exploit spatial degrees of freedom (DoF) in wireless systems [7, 8].
 MIMO systems provide significant gains in beamforming, spatial multiplexing, and diversity by leveraging independent or quasi-independent channel fading, thereby significantly improving spectral efficiency and link reliability.
 However, in classical designs, antenna arrays are fixed in position and orientation, limiting spatial control to electronic beamforming.
-For LAWNs, where users and targets are distributed in full 3D and may move at low altitude with complex trajectories, purely electronic beamforming is often not sufficient: system performance becomes strongly dependent on the relative 3D geometry among the base station (BS), scatterers, and low-altitude objects.
+For LAWNs, where users and targets are distributed in full 3D and may move at low altitude with complex trajectories, purely electronic beamforming is often not sufficient, as system performance becomes strongly dependent on the relative 3D geometry among the base station (BS), scatterers, and low-altitude objects.
 This motivates the exploitation of geometry-aware architectures that can reshape both the propagation environment and the array response itself for LAWN-specific ISAC.
-Reconfigurable intelligent surfaces (RISs) offer a promising approach to tailor the wireless environment via nearly passive phase-shift elements that reflect incident signals toward desired directions [9]-[11].
+Reconfigurable intelligent surfaces (RISs) offer a promising approach to tailor the wireless environment via nearly passive phase-shift elements that reflect incident signals toward desired directions [9, 10, 11, 12, 13].
 RIS-aided systems have been extensively investigated for coverage enhancement, interference mitigation, and energy-efficient MIMO operation.
-More recently, in addition to communication, RIS has been shown to improve wireless sensing, including target detection and localization, by redirecting illumination toward non-line-of-sight (NLoS) targets or acting as a passive reflector from favorable angles, thereby enriching spatial diversity [12]-[14].
-In parallel, movable antennas (MA) has recently been proposed to further enhance spatial flexibility by allowing antenna elements or arrays to change their positions within a small region [15]-[18], enabling substantial gains in diversity, multiplexing, and interference management [19], and have been extended to multicast scenarios [20], as well as to two-timescale designs exploiting statistical channel state information (CSI) [21].
-These works demonstrate that adding mechanical degrees of freedom to electronic beamforming can significantly improve performance.
-However, most existing MA- or RIS-related works treat antenna movement or passive phase control in isolation [22], and their ISAC designs for LAWN-like geometries are much less explored: existing ISAC designs typically assume fixed arrays and do not exploit low-altitude 3D geometry control to shape the transmit beampattern and the multiuser downlink channels simultaneously [23].
+More recently, in addition to communication, RIS has been shown to improve wireless sensing, including target detection and localization, by redirecting illumination toward non-line-of-sight (NLoS) targets or acting as a passive reflector from favorable angles, thereby enriching spatial diversity [14, 15, 16].
+In parallel, movable antennas (MA) has recently been proposed to further enhance spatial flexibility by allowing antenna elements or arrays to change their positions within a small region [17, 18, 19, 20], enabling substantial gains in diversity, multiplexing, and interference management [21], and have been extended to multicast scenarios [22], as well as to two-timescale designs exploiting statistical channel state information (CSI) [23].
+These and other related works on MA-enabled systems demonstrate that adding mechanical DoF to electronic beamforming can significantly improve performance [24, 25, 26, 27].
+However, most existing MA- or RIS-related works treat antenna movement or passive phase tuning in isolation [28], and their ISAC designs for LAWN-like geometries are much less explored.
+Existing ISAC designs typically assume fixed arrays and do not exploit low-altitude 3D geometry control to shape the transmit beampattern and the multiuser downlink channels simultaneously [29].
 Furthermore, the mechanical flexibility in the MA literature is typically modeled as small-scale position changes within a local region.
-In contrast, realistic low-altitude platforms (e.g., UAV-mounted arrays or rooftop BS panels) can be designed to support rotation of the entire active or passive array via low-cost mechanical actuators [24].
-Such rotatable arrays (RA) can jointly adjust their boresight directions, element coordinates, and directional gains, offering a distinct and practically relevant form of geometry control [25].
+In contrast, realistic low-altitude platforms, e.g., UAV-mounted arrays or rooftop BS panels, can be designed to support rotation of the entire active or passive array via low-cost mechanical actuators [30, 31, 32].
+Such rotatable arrays (RA) can jointly adjust their boresight directions, element coordinates, and directional gains, offering a distinct and practically relevant form of geometry control [33, 34].
 In summary, RIS-aided ISAC designs that consider rotatable passive arrays and their coupling with active-array rotations at the BS have yet to be systematically studied, especially under explicit constraints on the similarity of sensing beampatterns.
-As a result, the fundamental question of how to jointly design rotatable active and passive arrays, RIS phase shifts, and multiuser transmit beamforming for low-altitude ISAC remains open.
+As a result, the fundamental question of how to jointly design rotatable active and passive arrays, RIS phase shifts, and multiuser transmit beamforming for low-altitude ISAC remains open [35].
 Motivated by these observations, this paper investigates a low-altitude ISAC architecture with cooperatively rotatable active and passive arrays.
-We consider a downlink LAWN scenario where a BS equipped with an active rotatable array simultaneously serves multiple single-antenna ground users and senses low-altitude targets in monostatic mode, assisted by a rotatable RIS termed passive array.
-Both arrays can rotate in 3D, and the resulting orientations jointly affect (i) the element coordinates in the steering vectors and (ii) direction-dependent element/array gains through a practical directivity model.
-Based on the rotation-aware channel model, we explicitly couple communication and sensing performance and formulate a joint design problem that maximizes the sum rate subject to a sensing constraint, a BS transmit-power budget, and mechanical rotation limits at both the BS and RIS, with design variables given by the BS ISAC precoder, the RIS phase-shift vector, and the 3D rotation angles of the two arrays.
-The main contributions of this work are summarized as
-  • We propose a LAWN-oriented ISAC architecture in which a rotatable active BS array cooperates with a rotatable passive RIS array to serve multiple communication users and jointly sense low-altitude targets.
-We develop a rotation-aware 3D multipath channel model that captures both phase shifts from rotated antenna/element positions and direction-dependent antenna/element directional radiation pattern gains.
-Based on this model, we derive closed-form expressions for the multiuser downlink SINR and the monostatic sensing beampattern, and formulate a new optimization problem that maximizes the communication sum rate under a sensing beampattern mean-squared-error (MSE) constraint, power budget on the BS, unit-modulus constraint on the RIS and mechanical rotation bounds on RA, thereby revealing mechanical rotations as additional spatial degrees of freedom for LAWN ISAC.
-  • We convert the constrained ISAC design into a penalized formulation with a hinge penalty on the sensing MSE and develop an alternating optimization (AO) framework that iteratively updates the BS precoder, RIS phase shifts, and array rotations.
+We consider a downlink LAWN scenario in which a BS equipped with an active rotatable array simultaneously serves multiple single-antenna ground users and senses low-altitude targets in monostatic mode, assisted by a rotatable RIS, termed a passive array.
+Both arrays can rotate in 3D, and the resulting orientations jointly affect the element coordinates and equivalent phases in the steering vectors and direction-dependent array gains through a practical radiation model.
+Based on the rotation-aware channel model, we formulate a joint design problem on the BS precoder, the RIS phase shift vector, and the rotation angles of arrays, which maximizes a weighted-sum ISAC utility of the downlink sum rate and the sensing beampattern similarity, subject to the BS transmit-power budget, RIS unit-modulus constraints, and mechanical rotation limits at both the BS and the RIS.
+The main contributions of this work are summarized as follows
+  • We propose a LAWN-oriented ISAC architecture where a rotatable active BS array cooperates with a rotatable passive RIS array.
+We develop a rotation-aware 3D multipath channel model that captures both phase shifts of steering vectors induced from rotated antenna and element positions and directional path gains induced from direction-dependent antenna and element radiation patterns.
+We formulate an optimization problem that maximizes the weighted-sum utility on communication sum rate and sensing beampattern similarity, which is characterized by normalized mean-squared error (NMSE) between the generated and desired beampattern, by jointly designing BS transmit beamforming, RIS phase shifts, and active and passive array rotations, subjected to power budget on the BS, unit-modulus constraint on the RIS, and mechanical rotation bounds on both RAs.
+  • We develop an alternating optimization (AO) framework that iteratively updates the BS precoder, RIS phase shifts, and array rotations.
 For the BS precoding block, we combine the quadratic transform (QT) and the Lagrangian dual transform (LDT) with a majorization–minimization (MM) step to obtain a tractable quadratic-constrained quadratic program (QCQP) subproblem that admits a closed-form water-filling-like update.
 For the RIS block, we exploit the complex-circle manifold structure and design an efficient Riemannian conjugate gradient (RCG) algorithm.
-For the rotation block, we derive closed-form gradients of both SINR and beampattern MSE with respect to the BS/RIS Euler angles and employ a projected gradient descent (PGD) method with Barzilai–Borwein initialization and Armijo backtracking.
+For the rotation block, we derive closed-form gradients of both SINR and beampattern NMSE with respect to the BS and RIS Euler angles and employ a projected gradient descent (PGD) method with Barzilai–Borwein initialization and Armijo backtracking.
   • We conduct extensive simulations using LAWN-inspired 3D geometries and compare the proposed design with several baselines, including only BS rotation, only RIS rotation, fully fixed arrays, and non-RIS configurations.
-The results show that the jointly rotatable BS–RIS architecture provides substantial sum-rate gains under the same sensing MSE constraint across varying transmit powers, antenna array sizes, and user densities.
-The gains are further amplified with realistic directional element patterns, highlighting the importance of aligning the array boresights and main lobes with both user and target directions in a low-altitude 3D space.
+The results show that the jointly rotatable BS and RIS architecture provides substantial gains in the weighted ISAC utility of the communication rate and beampattern NMSE across varying transmit powers, antenna array sizes, and user densities.
+The gains are further amplified by practical directional radiation patterns, highlighting the importance of aligning the array boresights and main lobes with both the user and target directions in low-altitude 3D space.
+By sweeping the trade-off weight, we further characterize the Pareto frontier between the sum rate and the NMSE, thereby quantifying how joint array rotations expand the ISAC trade-off boundary.
 The rest of the paper is structured as follows: Section II introduces the system model and problem formulation.
-Section III presents proposed algorithm.
+Section III presents the proposed algorithm.
 Section IV presents the numerical results, and Section V concludes the paper.
-Figure 1: Illustration of a low-altitude ISAC system with rotatable active and passive arrays.
+Notations: Scalars are denoted by italic letters, and vectors and matrices are denoted by bold-face lower and upper-case letters, respectively.
+$\mathbb{R}^{x}$ and $\mathbb{R}^{x\times y}$ denote the space of real vectors or matrices, and $\mathbb{C}^{x}$ and $\mathbb{C}^{x\times y}$ denote the space of complex vectors or matrices, respectively.
+$\left|x\right|$ denotes the modulus of a complex-valued scalar $x$.
+For a vector $\boldsymbol{x}$, $\left\|\boldsymbol{x}\right\|$ denotes its Euclidean norm and $\boldsymbol{x}^{*}$ denotes its conjugate.
+The distribution of a circularly symmetric complex Gaussian random vector with mean vector $\boldsymbol{x}$ and covariance matrix $\boldsymbol{\varSigma}$ is denoted by $\mathcal{CN}\left(\boldsymbol{x},\boldsymbol{\varSigma}\right)$.
+The Euclidean gradient of a scalar function $f(\boldsymbol{x})$ with a vector variable $\boldsymbol{x}$ is denoted by $\nabla f(\boldsymbol{x})$.
+For a square matrix $\boldsymbol{S}$, $\mathrm{Tr}\left(\boldsymbol{S}\right)$ denote its trace.
+For any general matrix, $\boldsymbol{M}^{H}$ and $\left[\boldsymbol{M}\right]_{ij}$ denote its conjugate transpose and $(i,j)$th element,respectively.
+$\boldsymbol{I}$ denotes an identity matrix, respectively.
+$\mathbb{E}[\cdot]$ stands for the expectation operation.
+Figure 1: A low-altitude ISAC system with rotatable active and passive arrays.
 II System Model and Problem Formulation
 We consider a downlink RA-enabled ISAC system, as shown in Fig.
-1, where a base station (BS) uses an active rotatable uniform planar array (UPA) with $M$ transmit antennas to serve $K$ single-antenna users while sensing multiple targets in a monostatic mode; at the same time, a passive rotatable UPA RIS with $N$ reflecting elements assists both tasks.
+1, where a BS uses an active rotatable uniform planar array (UPA) with $M$ transmit antennas to serve $K$ single-antenna users while sensing multiple targets in monostatic mode; at the same time, a passive rotatable RIS arranged as a UPA with $N$ reflecting elements assists both tasks.
 In addition to active transmit beamforming at the BS and passive phase shifts at the RIS, both arrays can rotate in 3D.
-These rotations steer the array boresights and reshape the array response to improve, or balance, communication and sensing performance.
+These rotations steer the array boresights and reshape the array response to improve or balance the performance between communication and sensing.
 II-A Array Geometry and Rotations
-We adopt a global right-handed Cartesian coordinate system (CCS) $\mathcal{F}_{0}=(O,x,y,z)$.
-The BS employs a co-located transmit/receive UPA with size $M=M_{x}\times M_{y}$, and the RIS is a UPA with size $N=N_{x}\times N_{y}$.
-Let $\overline{\boldsymbol{d}}^{\mathrm{B}}_{m}\in\mathbb{R}^{3}$ and $\overline{\boldsymbol{d}}^{\mathrm{R}}_{n}\in\mathbb{R}^{3}$ denote the unrotated local coordinates of the $m$-th BS transmit antenna and the $n$-th RIS reflecting element, respectively, and let $\boldsymbol{d}_{0}^{\mathrm{B}}$ and $\boldsymbol{d}_{0}^{\mathrm{R}}$ be their array centers in $\mathcal{F}_{0}$.
+We adopt a global right-handed Cartesian coordinate system (CCS) $\mathcal{G}=(O,\mathrm{x},\mathrm{y},\mathrm{z})$.
+The BS employs a co-located transmit/receive UPA with size $M=M_{\text{col}}\times M_{\text{row}}$ of $M_{\text{col}}$ columns and $M_{\text{row}}$ rows, and the RIS is also a UPA with size $N=N_{\text{col}}\times N_{\text{row}}$ of $N_{\text{col}}$ columns and $N_{\text{row}}$ rows.
+Let $\overline{\boldsymbol{d}}^{\text{B}}_{m}\in\mathbb{R}^{3}$ and $\overline{\boldsymbol{d}}^{\text{R}}_{n}\in\mathbb{R}^{3}$ denote the unrotated local coordinates of the $m$-th BS transmit antenna and the $n$-th RIS reflecting element, respectively, and let $\boldsymbol{d}_{0}^{\text{B}}$ and $\boldsymbol{d}_{0}^{\text{R}}$ be their array centers in $\mathcal{G}$.
 Both BS and RIS employ a motorized mechanism to rotate the arrays as a whole.
-Based on the system’s ISAC requirements, the proper rotation angles are determined by a central control center, which is connected to the motor controllers via fiber links.
-Each array rotates as a rigid body through Euler angles
-$\boldsymbol{r}^{\text{B}}\triangleq\big[\zeta_{x}^{\text{B}},\zeta_{y}^{\text{B}},\zeta_{z}^{\text{B}}\big]^{T},$ (1a)
-$\boldsymbol{r}^{\text{R}}\triangleq\big[\zeta_{x}^{\text{R}},\zeta_{y}^{\text{R}},\zeta_{z}^{\text{R}}\big]^{T},$ (1b)
+Based on the system’s ISAC requirements, the proper rotation angles are determined by a central control center, which is connected to the motor controllers via fiber links [^1] In practice, mechanical rotations can be updated on a slow timescale (e.g., sector-level adaptation or quasi-static geometry changes), whereas the BS digital precoder and RIS phases can be refreshed on the fast timescale of each channel-coherence block.
+Since the mechanical control involves only six scalar angles (3D Euler angles for the BS and RIS), the control signaling overhead is negligible compared with per-element electronic tuning.
+When the angles change, the array responses are updated deterministically; thus, standard pilot-based estimation and tracking remain applicable, and two-timescale designs using long-term statistical CSI for rotations are also feasible..
+Both the active array at BS and the passive RIS array rotate as a rigid body through Euler angles denoted by
+$\boldsymbol{r}^{\text{B}}\triangleq\big[r_{\mathrm{x}}^{\text{B}},r_{\mathrm{y}}^{\text{B}},r_{\mathrm{z}}^{\text{B}}\big]^{T},\quad\boldsymbol{r}^{\text{R}}\triangleq\big[r_{\mathrm{x}}^{\text{R}},r_{\mathrm{y}}^{\text{R}},r_{\mathrm{z}}^{\text{R}}\big]^{T},$ (1)
 respectively, in their local CCS.
-The corresponding rotation matrix generated by $\boldsymbol{r}^{\text{A}}$, where we define $\text{A}\in\{\text{B,R}\}$ for brevity, represents the rotations about their local axis with the standard intrinsic $z$-$y$-$x$ convention and is given by
-$\boldsymbol{R}\left(\zeta_{x}^{\text{A}},\zeta_{y}^{\text{A}},\zeta_{z}^{\text{A}}\right)=\boldsymbol{R}_{x}\left(\zeta_{x}^{\text{A}}\right)\boldsymbol{R}_{y}\left(\zeta_{y}^{\text{A}}\right),\boldsymbol{R}_{z}\left(\zeta_{z}^{\text{A}}\right),$ (2)
+In the following, we denote $\text{A}\in\{\text{B, R}\}$ as the notations for either BS or RIS array for brevity.
+Then, the corresponding rotation matrix generated by $\boldsymbol{r}^{\text{A}}$ represents the rotations about their local axis with the standard intrinsic $\mathrm{x}$-$\mathrm{y}$-$\mathrm{z}$ convention and is given by
+$\boldsymbol{R}\left(r_{\mathrm{x}}^{\text{A}},r_{\mathrm{y}}^{\text{A}},r_{\mathrm{z}}^{\text{A}}\right)=\boldsymbol{R}_{\mathrm{x}}\left(r_{\mathrm{x}}^{\text{A}}\right)\boldsymbol{R}_{\mathrm{y}}\left(r_{\mathrm{y}}^{\text{A}}\right)\boldsymbol{R}_{\mathrm{z}}\left(r_{\mathrm{z}}^{\text{A}}\right),$ (2)
 where the rotation matrices on each axis are defined by
-$\boldsymbol{R}_{x}\left(\zeta_{x}^{\text{A}}\right)=\left[\begin{matrix}1&0&0\\ 0&\cos\zeta_{x}^{\text{A}}&-\sin\zeta_{x}^{\text{A}}\\ 0&\sin\zeta_{x}^{\text{A}}&\cos\zeta_{x}^{\text{A}}\\ \end{matrix}\right],$ (3a)
-$\boldsymbol{R}_{y}\left(\zeta_{y}^{\text{A}}\right)=\left[\begin{matrix}\cos\zeta_{y}^{\text{A}}&0&\sin\zeta_{y}^{\text{A}}\\ 0&1&0\\ -\sin\zeta_{y}^{\text{A}}&0&\cos\zeta_{y}^{\text{A}}\\ \end{matrix}\right],$ (3b)
-$\boldsymbol{R}_{z}\left(\zeta_{z}^{\text{A}}\right)=\left[\begin{matrix}\cos\zeta_{z}^{\text{A}}&-\sin\zeta_{z}^{\text{A}}&0\\ \sin\zeta_{z}^{\text{A}}&\cos\zeta_{z}^{\text{A}}&0\\ 0&0&1\\ \end{matrix}\right].$ (3c)
-The rotated antenna or element coordinates in $\mathcal{F}_{0}$ are then
+$\boldsymbol{R}_{\mathrm{x}}\left(r_{\mathrm{x}}^{\text{A}}\right)=\left[\begin{matrix}1&0&0\\ 0&\cos r_{\mathrm{x}}^{\text{A}}&-\sin r_{\mathrm{x}}^{\text{A}}\\ 0&\sin r_{\mathrm{x}}^{\text{A}}&\cos r_{\mathrm{x}}^{\text{A}}\\ \end{matrix}\right],$ (3a)
+$\boldsymbol{R}_{\mathrm{y}}\left(r_{\mathrm{y}}^{\text{A}}\right)=\left[\begin{matrix}\cos r_{\mathrm{y}}^{\text{A}}&0&\sin r_{\mathrm{y}}^{\text{A}}\\ 0&1&0\\ -\sin r_{\mathrm{y}}^{\text{A}}&0&\cos r_{\mathrm{y}}^{\text{A}}\\ \end{matrix}\right],$ (3b)
+$\boldsymbol{R}_{\mathrm{z}}\left(r_{\mathrm{z}}^{\text{A}}\right)=\left[\begin{matrix}\cos r_{\mathrm{z}}^{\text{A}}&-\sin r_{\mathrm{z}}^{\text{A}}&0\\ \sin r_{\mathrm{z}}^{\text{A}}&\cos r_{\mathrm{z}}^{\text{A}}&0\\ 0&0&1\\ \end{matrix}\right].$ (3c)
+The rotated antenna or element coordinates in $\mathcal{G}$ are then
 $\boldsymbol{d}_{m}^{\text{B}}\left(\boldsymbol{r}^{\text{B}}\right)=\boldsymbol{d}_{0}^{\text{B}}+\boldsymbol{R}\left(\boldsymbol{r}^{\text{B}}\right)\overline{\boldsymbol{d}}_{m}^{\text{B}},$ (4a)
@@
 II-B Direction and Steering Vectors
-For a far-field direction parameterized by an elevation-azimuth pair $(\vartheta,\varphi)$ with elevation from the $xOy$ plane and azimuth from the $x$-axis, define the direction vector
+For a far-field direction parameterized by an elevation-azimuth pair $(\vartheta,\varphi)$ with elevation from the $xOy$ plane and azimuth from the $x$-axis, we define the departure direction vector as the unit propagation direction pointing from the array center to the far-field point
 $\boldsymbol{u}(\vartheta,\varphi)\triangleq[\cos\vartheta\cos\varphi,\cos\vartheta\sin\varphi,\sin\vartheta]^{T}.$ (5)
@@
 For the targeted sensing area with low-altitude UAVs, we consider merely the single LoS path with a potential reflected echo signal in each desired direction.
-Accordingly, we use the following indexed notational shorthands for distinct BS-user (BU), RIS-BS (RB), BS-RIS (BR), and RIS-user (RU) links, respectively
+Accordingly, we use the following indexed notational shorthands for distinct BS-user (BU), RIS-BS (RB), BS-RIS (BR), and RIS-user (RU) links, respectively, and define the direction vectors as follows
 $\boldsymbol{u}_{k,l}^{\text{BU}}\!=\!\left[\cos(\vartheta_{k,l}^{\text{BU}})\cos(\varphi_{k,l}^{\text{BU}}),\cos(\vartheta_{k,l}^{\text{BU}})\sin(\varphi_{k,l}^{\text{BU}}),\sin(\vartheta_{k,l}^{\text{BU}})\right]^{T}\!,$ (6a)
@@
 $\!\boldsymbol{t}_{p}^{\text{BR}}(\boldsymbol{r}^{\text{B}})\!=\!\big[e^{j\frac{2\pi}{\lambda}(\boldsymbol{u}_{p}^{\text{BR}})^{T}\boldsymbol{d}_{1}^{\text{B}}(\boldsymbol{r}^{\text{B}})},\dots,e^{j\frac{2\pi}{\lambda}(\boldsymbol{u}_{p}^{\text{BR}})^{T}\boldsymbol{d}_{M}^{\text{B}}(\boldsymbol{r}^{\text{B}})}\big]^{T}\!,\!\!$ (7b)
-$\!\boldsymbol{t}_{p}^{\text{RB}}(\boldsymbol{r}^{\text{R}})\!=\!\big[e^{j\frac{2\pi}{\lambda}(\boldsymbol{u}_{p}^{\text{BR}})^{T}\boldsymbol{d}_{1}^{\text{R}}(\boldsymbol{r}^{\text{R}})},\dots,e^{j\frac{2\pi}{\lambda}(\boldsymbol{u}_{p}^{\text{BR}})^{T}\boldsymbol{d}_{1}^{\text{R}}(\boldsymbol{r}^{\text{R}})}\big]^{T}\!,\!\!$ (7c)
-$\!\boldsymbol{t}_{k,q}^{\text{RU}}(\boldsymbol{r}^{\text{B}})\!=\!\big[e^{j\frac{2\pi}{\lambda}(\boldsymbol{u}_{k,q}^{\text{RU}})^{T}\boldsymbol{d}_{1}^{\text{R}}(\boldsymbol{r}^{\text{R}})},\dots,e^{j\frac{2\pi}{\lambda}(\boldsymbol{u}_{k,q}^{\text{RU}})^{T}\boldsymbol{d}_{N}^{\text{R}}(\boldsymbol{r}^{\text{R}})}\big]^{T}\!.\!\!$ (7d)
+$\!\boldsymbol{t}_{p}^{\text{RB}}(\boldsymbol{r}^{\text{R}})\!=\!\big[e^{j\frac{2\pi}{\lambda}(\boldsymbol{u}_{p}^{\text{RB}})^{T}\boldsymbol{d}_{1}^{\text{R}}(\boldsymbol{r}^{\text{R}})},\dots,e^{j\frac{2\pi}{\lambda}(\boldsymbol{u}_{p}^{\text{RB}})^{T}\boldsymbol{d}_{N}^{\text{R}}(\boldsymbol{r}^{\text{R}})}\big]^{T}\!,\!\!$ (7c)
+$\!\boldsymbol{t}_{k,q}^{\text{RU}}(\boldsymbol{r}^{\text{R}})\!=\!\big[e^{j\frac{2\pi}{\lambda}(\boldsymbol{u}_{k,q}^{\text{RU}})^{T}\boldsymbol{d}_{1}^{\text{R}}(\boldsymbol{r}^{\text{R}})},\dots,e^{j\frac{2\pi}{\lambda}(\boldsymbol{u}_{k,q}^{\text{RU}})^{T}\boldsymbol{d}_{N}^{\text{R}}(\boldsymbol{r}^{\text{R}})}\big]^{T}\!.\!\!$ (7d)
 Similarly, for a target direction $(\vartheta_{a},\varphi_{a})$ within the target sensing area $\mathcal{A}$, the direction steering vectors for BS-target (BT) and RIS-target (RT) links can be respectively given by
@@
 $\boldsymbol{t}_{a}^{\text{BT}}(\boldsymbol{r}^{\text{B}})=\big[e^{j\frac{2\pi}{\lambda}(\boldsymbol{u}_{a}^{\text{BT}})^{T}\boldsymbol{d}_{1}^{\text{B}}(\boldsymbol{r}^{\text{B}})},\dots,e^{j\frac{2\pi}{\lambda}(\boldsymbol{u}_{a}^{\text{BT}})^{T}\boldsymbol{d}_{M}^{\text{B}}(\boldsymbol{r}^{\text{B}})}\big]^{T},$ (9a)
-$\boldsymbol{t}_{a}^{\text{RT}}(\boldsymbol{r}^{\text{B}})=\big[e^{j\frac{2\pi}{\lambda}(\boldsymbol{u}_{a}^{\text{RT}})^{T}\boldsymbol{d}_{1}^{\text{R}}(\boldsymbol{r}^{\text{R}})},\dots,e^{j\frac{2\pi}{\lambda}(\boldsymbol{u}_{a}^{\text{RT}})^{T}\boldsymbol{d}_{N}^{\text{R}}(\boldsymbol{r}^{\text{R}})}\big]^{T}.$ (9b)
+$\boldsymbol{t}_{a}^{\text{RT}}(\boldsymbol{r}^{\text{R}})=\big[e^{j\frac{2\pi}{\lambda}(\boldsymbol{u}_{a}^{\text{RT}})^{T}\boldsymbol{d}_{1}^{\text{R}}(\boldsymbol{r}^{\text{R}})},\dots,e^{j\frac{2\pi}{\lambda}(\boldsymbol{u}_{a}^{\text{RT}})^{T}\boldsymbol{d}_{N}^{\text{R}}(\boldsymbol{r}^{\text{R}})}\big]^{T}.$ (9b)
 As shown in (7) and (9), array rotations, which generate position-dependent effects on the steering vector’s phase term, can be exploited to reshape channel responses.
@@
 Let $\bar{\boldsymbol{n}}=[0,0,1]^{T}$ and
-$\boldsymbol{n}(\boldsymbol{r}^{\mathrm{A}})=\boldsymbol{R}(\boldsymbol{r}^{\mathrm{A}})\bar{\boldsymbol{n}}$ (10)
+$\boldsymbol{n}(\boldsymbol{r}^{\text{A}})=\boldsymbol{R}(\boldsymbol{r}^{\text{A}})\bar{\boldsymbol{n}},$ (10)
 be the unrotated and rotated array boresight, respectively, we adopt a practical directivity model
@@
 […]
-where $G_{0}^{\text{B}}$ and $G_{0}^{\text{R}}$ denotes the maximum gain achieved when the signal aligns with the boresight, and $p^{\text{B}}$ and $p^{\text{R}}$ are the directivity factors controlling the main-lobe width, following from directivity normalization over the full sphere.
-Then, with given numbers of resolvable paths $L_{k}$, $P$, and $Q_{k}$, and small-scale complex gains $a_{k,l}$, $b_{p}$, and $c_{k,q}$ that absorb path loss and fading, the BS-to-user channel $\boldsymbol{h}_{k}\left(\boldsymbol{r}^{\text{B}}\right)\in\mathbb{C}^{M}$, BS-to-RIS channel $\boldsymbol{B}\left(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}}\right)\in\mathbb{C}^{M\times N}$, and RIS-to-user channel $\boldsymbol{g}_{k}\left(\boldsymbol{r}^{\text{R}}\right)\in\mathbb{C}^{N}$ are respectively expressed as
-$\boldsymbol{h}_{k}\left(\boldsymbol{r}^{\text{B}}\right)=\sum_{l=1}^{L_{k}}{a_{k,l}\sqrt{G_{k,l}^{\text{B}}\left(\boldsymbol{r}^{\text{B}}\right)}}\boldsymbol{t}_{k,l}^{\text{BU}}\left(\boldsymbol{r}^{\text{B}}\right)$ (12a)
-$\boldsymbol{B}\left(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}}\right)=\sum_{p=1}^{P}{b_{p}\sqrt{G_{k,p}^{\text{B}}\left(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}}\right)G_{k,p}^{\text{R}}\left(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}}\right)}}$
-$\qquad\qquad\qquad\quad\,\,\times\boldsymbol{t}_{p}^{\text{BR}}\left(\boldsymbol{r}^{\text{B}}\right)\left(\boldsymbol{t}_{p}^{\text{RB}}\left(\boldsymbol{r}^{\text{R}}\right)\right)^{H}$ (12b)
-$\boldsymbol{g}_{k}\left(\boldsymbol{r}^{\text{R}}\right)=\sum_{q=1}^{Q_{k}}{c_{k,q}\sqrt{G_{k,q}^{\text{R}}\left(\boldsymbol{r}^{\text{R}}\right)}}\boldsymbol{t}_{k,l}^{\text{RU}}\left(\boldsymbol{r}^{\text{R}}\right).$ (12c)
+where $G_{0}^{\text{B}}$ and $G_{0}^{\text{R}}$ denotes the maximum gain achieved when the signal aligns with the boresight, and $b^{\text{B}}$ and $b^{\text{R}}$ are the directivity factors controlling the main-lobe width, following from directivity normalization over the full sphere.
+Then, with given numbers of resolvable paths $L_{k}$, $P$, and $Q_{k}$, and small-scale complex gains $\delta_{k,l}$, $\gamma_{p}$, and $\psi_{k,q}$ that absorb path loss and fading, the BS-to-user channel $\boldsymbol{h}_{k}\left(\boldsymbol{r}^{\text{B}}\right)\in\mathbb{C}^{M}$, BS-to-RIS channel $\boldsymbol{B}\left(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}}\right)\in\mathbb{C}^{M\times N}$, and RIS-to-user channel $\boldsymbol{g}_{k}\left(\boldsymbol{r}^{\text{R}}\right)\in\mathbb{C}^{N}$ are respectively expressed as
+$\boldsymbol{h}_{k}\left(\boldsymbol{r}^{\text{B}}\right)=\sum_{l=1}^{L_{k}}{\delta_{k,l}\sqrt{G_{k,l}^{\text{B}}\left(\boldsymbol{r}^{\text{B}}\right)}}\boldsymbol{t}_{k,l}^{\text{BU}}\left(\boldsymbol{r}^{\text{B}}\right),$ (12a)
+$\boldsymbol{B}\left(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}}\right)=\sum_{p=1}^{P}{\gamma_{p}\sqrt{G_{p}^{\text{B}}\left(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}}\right)G_{p}^{\text{R}}\left(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}}\right)}}$
+$\qquad\qquad\qquad\quad\,\,\times\boldsymbol{t}_{p}^{\text{BR}}\left(\boldsymbol{r}^{\text{B}}\right)\left(\boldsymbol{t}_{p}^{\text{RB}}\left(\boldsymbol{r}^{\text{R}}\right)\right)^{H},$ (12b)
+$\boldsymbol{g}_{k}\left(\boldsymbol{r}^{\text{R}}\right)=\sum_{q=1}^{Q_{k}}{\psi_{k,q}\sqrt{G_{k,q}^{\text{R}}\left(\boldsymbol{r}^{\text{R}}\right)}}\boldsymbol{t}_{k,q}^{\text{RU}}\left(\boldsymbol{r}^{\text{R}}\right).$ (12c)
 Taking into account direct and cascaded RIS-enabled links, the effective composite channel from BS to the $k$-th user is
-$\boldsymbol{f}_{k}(\boldsymbol{r}^{\text{B}}\!,\boldsymbol{r}^{\text{R}}\!,\boldsymbol{\theta})\!=\!\boldsymbol{h}_{k}(\boldsymbol{r}^{\text{B}})\!+\!\boldsymbol{B}(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}})\text{diag}(\boldsymbol{\theta})\boldsymbol{g}_{k}(\boldsymbol{r}^{\text{R}}),\!\!$ (13)
+$\boldsymbol{f}_{k}(\boldsymbol{r}^{\text{B}}\!,\boldsymbol{r}^{\text{R}}\!,\boldsymbol{\theta})\!=\!\boldsymbol{h}_{k}(\boldsymbol{r}^{\text{B}})\!+\!\boldsymbol{B}(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}})\mathrm{diag}(\boldsymbol{\theta})\boldsymbol{g}_{k}(\boldsymbol{r}^{\text{R}}),\!\!$ (13)
 where $\boldsymbol{\theta}=[\theta_{1},\ldots,\theta_{N}]^{T}$ collects the RIS reflecting phase shift coefficients, typically being passive with $|\theta_{n}|=1$.
 Similarly, for sensing, define the effective steering toward a target direction $(\vartheta_{a},\varphi_{a})$ as
-$\boldsymbol{f}_{\text{S},a}(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}},\boldsymbol{\theta})=$ $\sqrt{G_{a}^{\text{B}}\left(\boldsymbol{r}^{\text{B}}\right)}a_{a}\boldsymbol{t}_{a}^{\text{BT}}\left(\boldsymbol{r}^{\text{B}}\right)+\!\boldsymbol{B}(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}})\text{diag}(\boldsymbol{\theta})$
-$\times\sqrt{G_{a}^{\text{R}}\left(\boldsymbol{r}^{\text{B}}\right)}c_{a}\boldsymbol{t}_{a}^{\text{RT}}(\boldsymbol{r}^{\text{B}})$ (14)
-II-D Transmit Signal, Communication SINR, and Sensing MSE
+$\boldsymbol{f}_{\text{S},a}(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}},\boldsymbol{\theta})=$ $\sqrt{G_{a}^{\text{B}}\left(\boldsymbol{r}^{\text{B}}\right)}\boldsymbol{t}_{a}^{\text{BT}}\left(\boldsymbol{r}^{\text{B}}\right)+\!\boldsymbol{B}(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}})\mathrm{diag}(\boldsymbol{\theta})$
+$\times\sqrt{G_{a}^{\text{R}}\left(\boldsymbol{r}^{\text{R}}\right)}\boldsymbol{t}_{a}^{\text{RT}}(\boldsymbol{r}^{\text{R}}).$ (14)
+II-D Transmit Signal, Communication SINR, and Sensing NMSE
 The BS transmits an ISAC signal
 $\boldsymbol{x}=\boldsymbol{W}_{\text{C}}\boldsymbol{s}_{\text{C}}+\boldsymbol{W}_{\text{S}}\boldsymbol{s}_{\text{S}},$ (15)
-where $\boldsymbol{W}_{\text{C}}=[\boldsymbol{w}_{1},\dots,\boldsymbol{w}_{K}\boldsymbol{]}\in\mathbb{C}^{M\times K}$ collects $K$ communication beams, $\boldsymbol{W}_{\text{S}}\in\mathbb{C}^{M\times M}$ contains dedicated sensing beams, $\boldsymbol{s}_{\text{C}}=[s_{1},\dots,s_{K}]^{T}$ denotes the communication symbols, and $\boldsymbol{s}_{\text{S}}$ denotes the sensing signals, with $\mathbb{E}\left[\boldsymbol{s}_{\text{S}}\boldsymbol{s}_{\text{S}}^{H}\right]=\boldsymbol{I}_{M}$, $\boldsymbol{s}_{\text{C}}\sim\mathcal{C}\mathcal{N}\left(\boldsymbol{0},\boldsymbol{I}_{K}\right)$, and $\mathbb{E}\left[\boldsymbol{s}_{\text{S}}\boldsymbol{s}_{\text{C}}^{H}\right]=0$.
+where $\boldsymbol{W}_{\text{C}}=[\boldsymbol{w}_{1},\dots,\boldsymbol{w}_{K}\boldsymbol{]}\in\mathbb{C}^{M\times K}$ collects $K$ communication beams, $\boldsymbol{W}_{\text{S}}\in\mathbb{C}^{M\times M}$ contains dedicated sensing beams, $\boldsymbol{s}_{\text{C}}=[s_{\text{C},1},\dots,s_{\text{C},K}]^{T}$ denotes the communication symbols, and $\boldsymbol{s}_{\text{S}}$ denotes the sensing signals, with $\mathbb{E}\left[\boldsymbol{s}_{\text{S}}\boldsymbol{s}_{\text{S}}^{H}\right]=\boldsymbol{I}_{M}$, $\boldsymbol{s}_{\text{C}}\sim\mathcal{C}\mathcal{N}\left(\boldsymbol{0},\boldsymbol{I}_{K}\right)$, and $\mathbb{E}\left[\boldsymbol{s}_{\text{S}}\boldsymbol{s}_{\text{C}}^{H}\right]=0$.
 Let $\boldsymbol{W}\triangleq\left[\boldsymbol{W}_{\text{C}},\boldsymbol{W}_{\text{S}}\right]$ denotes the BS ISAC transmit beamforming matrix, $\boldsymbol{s}\triangleq\left[\boldsymbol{s}_{\text{C}}^{T},\boldsymbol{s}_{\text{S}}^{T}\right]^{T}$, the received signal at the communication user $k$ is written as
-$y_{k}$ $=\boldsymbol{f}_{k}^{H}(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}},\boldsymbol{\theta})\boldsymbol{Ws}$ (16a)
-$=\boldsymbol{f}_{k}^{H}(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}},\boldsymbol{\theta})\boldsymbol{w}_{k}s_{k}+\sum_{i=1,i\neq k}^{K}{\boldsymbol{f}_{k}^{H}(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}},\boldsymbol{\theta})\boldsymbol{w}_{i}}x_{i}$
+$c_{k}$ $=\boldsymbol{f}_{k}^{H}(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}},\boldsymbol{\theta})\boldsymbol{Ws}$ (16a)
+$=\boldsymbol{f}_{k}^{H}(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}},\boldsymbol{\theta})\boldsymbol{w}_{k}s_{\text{C},k}+\sum_{i=1,i\neq k}^{K}{\boldsymbol{f}_{k}^{H}(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}},\boldsymbol{\theta})\boldsymbol{w}_{i}}s_{\text{C},i}$
 $\quad+\boldsymbol{f}_{k}^{H}(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}},\boldsymbol{\theta})\boldsymbol{W}_{\text{S}}\boldsymbol{s}_{\text{S}}+n_{k},$ (16b)
 where $n_{k}\sim\mathcal{C}\mathcal{N}\left(0,\sigma_{k}^{2}\right)$ is the complex additive white Gaussian noise.
-Taking further $\boldsymbol{W}\triangleq\left[\boldsymbol{w}_{1},\dots,\boldsymbol{w}_{K+N}\right]$, where $\boldsymbol{w}_{i}$ is the $i$-th column of $\boldsymbol{W}$, the SINR of the $k$-th user is
+Taking further $\boldsymbol{W}\triangleq\left[\boldsymbol{w}_{1},\dots,\boldsymbol{w}_{K+M}\right]$, where $\boldsymbol{w}_{i}$ is the $i$-th column of $\boldsymbol{W}$, the SINR of the $k$-th user is
 $\text{SINR}_{k}=\frac{\big|\boldsymbol{f}_{k}^{H}\left(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}},\boldsymbol{\theta}\right)\boldsymbol{w}_{k}\big|^{2}}{\sum_{i=1,i\neq k}^{K+M}{\big|\boldsymbol{f}_{k}^{H}\left(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}},\boldsymbol{\theta}\right)\boldsymbol{w}_{i}\big|^{2}}+\sigma_{k}^{2}},$ (17)
@@
 $=\boldsymbol{f}_{\text{S},a}^{H}(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}},\boldsymbol{\theta})\boldsymbol{WW}^{H}\boldsymbol{f}_{\text{S},a}(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}},\boldsymbol{\theta}).$ (18b)
-Using (13) this expands to
+Using (13), (18) expands to
 $\mathcal{P}(\vartheta_{a},\varphi_{a})$
@@
 $\quad\quad\times\boldsymbol{WW}^{H}\boldsymbol{B}(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}})\mathrm{diag}(\boldsymbol{\theta})\boldsymbol{t}_{a}^{\text{RT}}(\boldsymbol{r}^{\text{R}}).$ (19)
-Given a discretized grid $\{(\vartheta_{a},\varphi_{a})\}_{a=1}^{A},a\in\{1,\dots,A\}$ with $A$ sampled points in total, which samples the azimuth and elevation angular domain over the target sensing area $\mathcal{A}$, along with the desired beam pattern $\mathcal{P}_{\text{d}}(\vartheta_{a},\varphi_{a})$, the mean squared error (MSE) between $\mathcal{P}$ and $\mathcal{P}_{\text{d}}$ to evaluate the similarity of the sensing beam pattern is defined as []
+Given a discretized grid $\{(\vartheta_{a},\varphi_{a})\}_{a=1}^{A},a\in\{1,\dots,A\}$ with $A$ sampled points in total, which samples the azimuth and elevation angular domain over the target sensing area $\mathcal{A}$, along with the desired beampattern $\mathcal{P}_{\text{d}}(\vartheta_{a},\varphi_{a})$, the mean squared error (MSE) between $\mathcal{P}$ and $\mathcal{P}_{\text{d}}$ to evaluate the similarity of the sensing beampattern is defined as [41]
 $\text{MSE}\!=\!\frac{1}{A}\!\sum_{a=1}^{A}\!{\left|\mathcal{P}(\vartheta_{a},\varphi_{a};\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}},\boldsymbol{\theta},\boldsymbol{W})\!-\!\mathcal{P}_{\text{d}}(\vartheta_{a},\varphi_{a})\right|^{2}}\!.\!\!\!$ (20)
-A typical ideal desired beam pattern is given by
+A typical spotlight-type desired beampattern is given by
 […]
-where $\mathcal{T}$ is the set of focused grid points in the region $\mathcal{A}$.
+where $\mathcal{F}$ is the set of focused grid points in the region $\mathcal{A}$.
+While MSE is intuitive, it is scale-sensitive: the same beampattern shape may lead to very different MSE values when the overall beampattern level fluctuates due to transmit-power changes, propagation loss, blockage, or orientation-dependent element gains, which are prevalent in low-altitude geometries.
+To obtain a dimensionless and scale-robust metric that emphasizes shape fidelity rather than absolute amplitude, we adopt the NMSE metric with a scaling parameter $\iota>0$ as
+$\text{NMSE}\triangleq\frac{\sum_{a=1}^{A}{\left|\mathcal{P}_{a}-\iota\mathcal{P}_{d,a}\right|^{2}}}{\sum_{a=1}^{A}{\left|\iota\mathcal{P}_{d,a}\right|^{2}}}=\frac{\left\|\boldsymbol{p}-\iota\boldsymbol{p}_{\text{d}}\right\|_{2}^{2}}{\left\|\iota\boldsymbol{p}_{\text{d}}\right\|_{2}^{2}}.$ (22)
+where $\boldsymbol{p}=[\mathcal{P}_{1},\ldots,\mathcal{P}_{A}]^{T}$ and $\boldsymbol{p}_{\text{d}}=[\mathcal{P}_{\text{d},1},\ldots,\mathcal{P}_{\text{d},A}]^{T}$.
+The scaling parameter $\iota$ decouples the global amplitude mismatch from the spatial shape mismatch: even when $\mathcal{P}_{a}$ is uniformly scaled by channel loss or array gain variations, $\text{NMSE}(\iota)$ can remain small as long as the angular shape matches.
+Moreover, using $\|\iota\boldsymbol{p}_{\text{d}}\|_{2}^{2}$ in the denominator avoids the trivial solution $\iota=0$, which would artificially reduce the numerator, thereby enforcing a meaningful normalization [^2] The adopted beampattern matching NMSE is a design-friendly proxy for sensing quality: a smaller NMSE indicates better mainlobe fidelity in the intended angular sectors and reduced leakage or sidelobes elsewhere, which typically improves the received peak-to-maximum-sidelobe ratio (PSLR) and signal-to-clutter-plus-noise ratio (SCNR), and the reliability of detection or estimation in the region of interest.
+For fixed $(\boldsymbol{W},\boldsymbol{\theta},\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}})$, we update $\iota$ by minimizing $\text{NMSE}(\iota)$ in closed form.
+Let $c=\|\boldsymbol{p}\|_{2}^{2}$, $d=\boldsymbol{p}_{\text{d}}^{T}\boldsymbol{p}$, and $e=\|\boldsymbol{p}_{\text{d}}\|_{2}^{2}$.
+When $d>0$, the minimizer is
+$\iota^{\star}=\arg\min_{\iota>0}\text{NMSE}(\iota)=\frac{\|\boldsymbol{p}\|_{2}^{2}}{\boldsymbol{p}_{\text{d}}^{T}\boldsymbol{p}}.$ (23)
+In practice, to avoid numerical issues when $\boldsymbol{p}_{\text{d}}^{T}\boldsymbol{p}$ is very small (e.g., at initialization), we use a safeguarded update $\iota^{\star}=\|\boldsymbol{p}\|_{2}^{2}/(\boldsymbol{p}_{\text{d}}^{T}\boldsymbol{p}+\varepsilon)$ with a small $\varepsilon>0$.
+With $\iota=\iota^{\star}$, the NMSE in (22) measures how well the synthesized pattern matches the desired pattern up to an overall scaling, which is particularly suitable when (i) the absolute beampattern level is affected by the BS power budget and propagation loss, (ii) the sensing objective primarily concerns mainlobe or sidelobe shape rather than absolute amplitude, and (iii) the system expects the sensing metric to be comparable across different system settings, e.g., $P_{\text{B}}$, $M$, NN, and user or target geometries.
 II-E Problem Formulation
-Accordingly, we jointly optimize the BS and RIS rotations, the RIS phase shifts, and the BS transmit beamforming to maximize the sum-rate performance for communication with the guaranteed sensing requirement on the beam pattern MSE
-$\text{(P1)}:\underset{\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}},\boldsymbol{\theta},\boldsymbol{W}}{\max}$ $\sum_{k=1}^{K}{\log_{2}\left(1+\text{SINR}_{k}\right)}$ (22a)
-s.t. $\text{MSE}\leq\eta_{\text{th}}$ (22b)
-$\|\boldsymbol{W}\|_{F}^{2}=\text{Tr}(\boldsymbol{WW}^{H})\leq P_{\text{B}},$ (22c)
-$|\theta_{n}|=1,\forall n\in\mathcal{N},$ (22d)
-$\zeta_{i}^{\text{B}}\!\in\left[\zeta_{i,\min}^{\text{B}},\zeta_{i,\max}^{\text{B}}\right],i\in\{x,y,z\},$ (22e)
-$\zeta_{i}^{\text{R}}\!\in\left[\zeta_{i,\min}^{\text{R}},\zeta_{i,\max}^{\text{R}}\right],i\in\{x,y,z\},$ (22f)
-where $\eta_{\text{th}}$ in (22b) is the maximum allowed sensing MSE, $P_{\mathrm{B}}$ in (22c) is the BS power budget, the unit-modulus constraint (23d) enforces a passive RIS, and the box constraints (22e) and (22f) capture mechanical limits of the array rotations at BS and RIS, respectively.
+Accordingly, we jointly optimize the BS and RIS rotations, the RIS phase shifts, and the BS transmit beamforming to maximize a weighted-sum utility of downlink communication sum rate and sensing beampattern fidelity.
+Specifically, we consider
+$\text{(P1)}:\underset{\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}},\boldsymbol{\theta},\boldsymbol{W}}{\max}$ $\sum_{k=1}^{K}{\log_{2}\left(1+\text{SINR}_{k}\right)}-\rho\text{NMSE}$ (24a)
+s.t. $\|\boldsymbol{W}\|_{F}^{2}=\mathrm{Tr}(\boldsymbol{WW}^{H})\leq P_{\text{B}},$ (24b)
+$|\theta_{n}|=1,\forall n\in\mathcal{N},$ (24c)
+$r_{\mathrm{i}}^{\text{B}}\!\in\left[r_{\mathrm{i},\min}^{\text{B}},r_{\mathrm{i},\max}^{\text{B}}\right],\mathrm{i}\in\{\mathrm{x},\mathrm{y},\mathrm{z}\},$ (24d)
+$r_{\mathrm{i}}^{\text{R}}\!\in\left[r_{\mathrm{i},\min}^{\text{R}},r_{\mathrm{i},\max}^{\text{R}}\right],\mathrm{i}\in\{\mathrm{x},\mathrm{y},\mathrm{z}\},$ (24e)
+where $\rho\geq 0$ in (24a) is a trade-off weight controlling the relative emphasis on communication throughput versus sensing beampattern matching, $P_{\text{B}}$ in (24b) is the BS power budget, the unit-modulus constraint (24c) enforces a passive RIS, and the box constraints (24d) and (24e) capture mechanical limits of the array rotations at BS and RIS, respectively.
+Note that the proposed formulation in (24a) directly targets the fundamental ISAC trade-off by optimizing a scalarized Pareto objective.
+Compared with a strict threshold form of constraint on NMSE, the weighted-sum design offers three practical advantages.
+First, in low-altitude scenarios, the sensing pattern requirement may become temporarily difficult or even infeasible under fast geometry variations and strong multiuser interference, while the weighted-sum objective avoids infeasibility, being robust to geometry variation, and yields a graceful performance trade-off.
+Second, hard-threshold formulations typically require penalty or dual updates and may introduce non-smooth hinge-type terms, which can complicate algorithm design and slow convergence, while a weighted sum leads to a single-level smooth objective, enabling simpler tuning and more stable optimization.
+Third, by sweeping $\rho$, one can efficiently explore the Pareto frontier between the sum rate and NMSE, which provides a comprehensive characterization of the communication-sensing trade-off beyond enforcing a single scenario-dependent threshold.
 From the formulated problem, we observe that the rotation variables $(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}})$ provide a geometric control that affects the steering vectors and directive gain, thereby complementing the electromagnetic control offered by active and passive beamforming at the RIS and BS arrays; consequently, jointly optimizing these blocks yields a significant DoF for low-altitude ISAC.
-However, (P1) is highly non-convex due to unit-modulus constraints, rotation-induced nonlinearities in the steering vector and directive gain, and the coupling among $\boldsymbol{r}^{\text{B}}$, $\boldsymbol{r}^{\text{R}}$, $\boldsymbol{\theta}$, and $\boldsymbol{W}$ in both SINR and MSE expressions.
+However, (P1) is highly non-convex and challenging to solve directly due to unit-modulus constraints, rotation-induced nonlinearities in the steering vector and directive gain, and the coupling among $\boldsymbol{r}^{\text{B}}$, $\boldsymbol{r}^{\text{R}}$, $\boldsymbol{\theta}$, and $\boldsymbol{W}$ in both SINR and MSE expressions.
 III Proposed Algorithm
-In this section, we convert (P1) to a penalized form and solve it by resorting to AO technique across three blocks: BS transmit beamforming $\boldsymbol{W}$, RIS passive beamforming $\boldsymbol{\theta}$, and array rotations $(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}})$.
-The three subproblems are handled by a quadratic transform (QT) and a majorization-minimization (MM) step for $\boldsymbol{W}$, Riemannian optimization for $\boldsymbol{\theta}$, and the projection gradient-based method for $(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}})$, respectively.
-III-A Penalty Method and AO Framework
-The beam pattern similarity constraints (22b), with a high-order non-convex MSE expression, limit the feasible region and hinder the solution of (P1).
-To handle this difficulty, we incorporate (22b) into the objective function as a hinge penalty term, obtaining a more tractable reformulated problem
-$\text{(P1.1)}:\underset{\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}},\boldsymbol{\theta},\boldsymbol{W}}{\max}$ $\sum_{k=1}^{K}{\log_{2}\left(1+\text{SINR}_{k}\left(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}},\boldsymbol{\theta},\boldsymbol{W}\right)\right)}$
-$\!\!\!\!-\rho\max\!\left\{\text{MSE}\left(\boldsymbol{r}^{\text{B}}\!,\boldsymbol{r}^{\text{R}}\!,\boldsymbol{\theta}\!,\boldsymbol{W}\right)\!-\!\eta_{\text{th}},0\right\}\!\!\!$ (23)
-s.t. $\left(22\text{c}\right),\left(22\text{d}\right),\left(22\text{e}\right),\left(22\text{f}\right)$
-where the adjustable penalty parameter $\rho\geq 0$ penalizes violations of (22b) with the hinge penalty function $\max\!\left\{\text{MSE}\!-\!\eta_{\text{th}},0\right\}$.
-When $\rho\rightarrow\infty$, any solution $\{(\boldsymbol{r}^{\text{B}})^{*},(\boldsymbol{r}^{\text{R}})^{*},\boldsymbol{\theta}^{*},\boldsymbol{W}^{*}\}$ of (P1.1) should meet the condition $\text{MSE}\leq\eta_{\text{th}}$, establishing the equivalence between (P1) and (P1.1).
-Since setting $\rho$ large at the beginning renders this approach ineffective, we initialize $\rho$ by a small value to find a good starting point, solve problem (P1.1) iteratively with gradually increased $\rho$, until $\text{MSE}\leq\eta_{\text{th}}$ holds and the objective function reaches the convergence criteria.
-$\rho$ increases only when $\mathrm{MSE}>\eta_{\text{th}}$ persists; once the constraint is satisfied, $\rho$ is kept fixed to avoid unnecessary rate loss.
-Even with fixed $\rho$, due to the non-concave objective function and the non-convex constraints with closely coupled variables, (P1.1) remains highly non-convex and challenging to solve directly.
-Nevertheless, observing that the problem becomes more tractable when optimizing one variable block at a time, we resort to the AO method: at each outer iteration, we cyclically optimize the transmit beamforming $\boldsymbol{W}$, the RIS phase shifts $\boldsymbol{\theta}$, and the array rotations $\{\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}}\}$, while keeping the others fixed, as detailed in the following subsections.
+In this section, we develop an efficient algorithm to tackle the non-concave objective and the non-convex constraints in (P1), where the optimization variables are tightly coupled through both the multiuser SINRs and the sensing NMSE.
+III-A AO framework
+A key observation is that (P1) becomes considerably more tractable when optimizing one variable block at a time.
+Hence, we resort to an AO framework that alternately updates three blocks in an iterative manner: the BS transmit beamforming matrix $\boldsymbol{W}$, the RIS passive beamforming vector $\boldsymbol{\theta}$, and the array rotations $(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}})$.
+Importantly, the sensing metric is $\text{NMSE}(\iota)$ in (22), where the scaling parameter $\iota$ is introduced to decouple global amplitude mismatch from angular shape mismatch.
+We update $\iota$ once per AO outer iteration using the closed-form minimizer $\iota^{\star}$ in (23), and keep it fixed when solving the three inner subproblems for $\boldsymbol{W}$, $\boldsymbol{\theta}$, and $(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}})$.
+This outer update is computationally negligible and does not introduce additional constraints.
+Moreover, since $\iota^{\star}$ minimizes $\text{NMSE}(\iota)$ for fixed $(\boldsymbol{W},\boldsymbol{\theta},\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}})$, and the sum-rate term is independent of $\iota$, updating $\iota$ with $\iota^{\star}$ makes the objective value in (24a) non-decreasing.
+Therefore, the $\iota$-update does not undermine the monotonic improvement of the overall AO procedure; it simply refines the sensing metric by removing the best-fit scaling mismatch at the current iterate.
+Then, in each iteration, the resulting three subproblems are handled by: i) a QT combined with a MM step for updating $\boldsymbol{W}$, ii) a Riemannian optimization method over the complex circle manifold for updating $\boldsymbol{\theta}$, and iii) a projected gradient-based method for updating $(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}})$ under box constraints, respectively.
+Due to the intrinsic non-convexity of (P1), global optimality is generally intractable.
+The proposed AO-based algorithm aims at computing a locally optimal or suboptimal solution that achieves a favorable communication-sensing trade-off.
+In practice, such a suboptimal solution is meaningful because it is obtained with low complexity, respects all hardware constraints such as unit-modulus phases and mechanical rotation limits, and can be efficiently warm-started and re-optimized as the low-altitude geometry evolves.
 III-B QT- and MM-based Transmit Beamforming
-With fixed $(\boldsymbol{\theta},\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}})$, the subproblem optimizing $\boldsymbol{W}$ reads
-$\text{(P2)}:\underset{\boldsymbol{W}}{\max}\,\,$ $\sum_{k=1}^{K}{\log_{2}\left(1+\text{SINR}_{k}(\boldsymbol{W})\right)}$
-$-\rho\max\left\{\text{MSE}(\boldsymbol{W})-\eta_{\text{th}},0\right\}$ (24)
-s.t. $\left(22\text{c}\right),$
-where the SINR and MSE expressions reduce to
-$\text{SINR}_{k}\left(\boldsymbol{W}\right)=\frac{|\boldsymbol{f}_{k}^{H}\boldsymbol{w}_{k}|^{2}}{\sum_{i=1,i\neq k}^{K+M}{|\boldsymbol{f}_{k}^{H}\boldsymbol{w}_{i}|^{2}}+\sigma_{k}^{2}},$ (25a)
-$\text{MSE}\left(\boldsymbol{W}\right)=\frac{1}{A}\sum_{a=1}^{A}{\left|\|\boldsymbol{W}^{H}\boldsymbol{f}_{\text{S},a}\|_{2}^{2}-\mathcal{P}_{\text{d},a}\right|^{2}}.$ (25b)
-The power constraint (22c) is convex, and the non-concavity of (24) originates from the fractional SINR and fourth-order MSE expressions, which are sequentially tackled as follows.
+With $(\boldsymbol{\theta},\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}})$ and $\iota$ fixed, the subproblem for optimizing $\boldsymbol{W}$ becomes
+$\text{(P2)}:\underset{\boldsymbol{W}}{\max}\,$ $\sum_{k=1}^{K}\log_{2}\left(1+\mathrm{SINR}_{k}(\boldsymbol{W})\right)-\rho\text{NMSE}(\boldsymbol{W})$ (25)
+[…]
+and the SINR and NMSE expressions reduce to
+$\text{SINR}_{k}(\boldsymbol{W})$ $=\frac{|\boldsymbol{f}_{k}^{H}\boldsymbol{w}_{k}|^{2}}{\sum_{i=1,i\neq k}^{K+M}{|\boldsymbol{f}_{k}^{H}\boldsymbol{w}_{i}|^{2}}+\sigma_{k}^{2}},$ (26a)
+$\text{NMSE}(\boldsymbol{W})$ $=\frac{\sum_{a=1}^{A}\big|\|\boldsymbol{W}^{H}\boldsymbol{f}_{\text{S},a}\|_{2}^{2}-\iota\mathcal{P}_{d,a}\big|^{2}}{\sum_{a=1}^{A}\left|\iota\mathcal{P}_{d,a}\right|^{2}}$ (26b)
+$=\frac{1}{D}\sum_{a=1}^{A}\big|\|\boldsymbol{W}^{H}\boldsymbol{f}_{\text{S},a}\|_{2}^{2}-\bar{\mathcal{P}}_{d,a}\big|^{2},$ (26c)
+where we denote $D\triangleq\sum_{a=1}^{A}\left|\iota\mathcal{P}_{d,a}\right|^{2}$ and $\bar{\mathcal{P}}_{d,a}\triangleq\iota\mathcal{P}_{d,a}$ determined by the fixed $\iota$ for brevity.
+Here, $\boldsymbol{f}_{k}$ denotes the effective communication channel of user $k$, and $\boldsymbol{f}_{\text{S},a}$ denotes the effective sensing steering vector at grid point $a$, both determined by the fixed $(\boldsymbol{\theta},\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}})$.
+As the power constraint (24b) is convex, the non-concavity of (25) mainly stems from the fractional SINR terms in (26a) and the fourth-order dependence of the sensing loss $\|\boldsymbol{W}^{H}\boldsymbol{f}_{\text{S},a}\|_{2}^{2}$ in (26c).
+In the sequel, we first apply the QT to reformulate the sum-rate into a tractable form, and then employ an MM step to construct a tight surrogate for the sensing-related quartic term, leading to a closed-form update of $\boldsymbol{W}$ with bisection search.
 III-B1 Sensing MSE term with MM-based second-order upper bound
 For brevity, let
-$g_{a}\left(\boldsymbol{W}\right)$ $\triangleq\big(\lVert\boldsymbol{W}^{H}\boldsymbol{f}_{\text{S},a}\rVert_{2}^{2}-\mathcal{P}_{\text{d},a}\big)^{2},$ (26a)
-$=\big(\text{Tr}(\boldsymbol{W}^{H}\boldsymbol{S}_{a}\boldsymbol{W})-\mathcal{P}_{\text{d},a}\big)^{2},$ (26b)
+$g_{a}(\boldsymbol{W})$ $\triangleq\big(\lVert\boldsymbol{W}^{H}\boldsymbol{f}_{\text{S},a}\rVert_{2}^{2}-\bar{\mathcal{P}}_{\text{d},a}\big)^{2},$ (27a)
+$=\big(\text{Tr}(\boldsymbol{W}^{H}\boldsymbol{S}_{a}\boldsymbol{W})-\bar{\mathcal{P}}_{\text{d},a}\big)^{2},$ (27b)
 with the rank-one positive semi-definite matrix defined by
@@
 $\boldsymbol{G}_{a}\left(\boldsymbol{W}\right)$ $\triangleq\nabla g_{a}\left(\boldsymbol{W}\right)$ (29a)
-$=4\big(\text{Tr}(\boldsymbol{W}^{H}\boldsymbol{S}_{a}\boldsymbol{W})-\mathcal{P}_{\text{d},a}\big)\boldsymbol{S}_{a}\boldsymbol{W}.$ (28b)
+$=4\big(\mathrm{Tr}(\boldsymbol{W}^{H}\boldsymbol{S}_{a}\boldsymbol{W})-\bar{\mathcal{P}}_{\text{d},a}\big)\boldsymbol{S}_{a}\boldsymbol{W}.$ (29b)
 At the current iterate $\boldsymbol{W}^{(t)}$, we apply the second-order Taylor expansion (also known as the Descent Lemma [36]) to construct a convex upper-bound surrogate function for $g_{a}(\boldsymbol{W})$ as
@@
 $\triangleq\mathcal{U}_{a}\big(\boldsymbol{W}\mid\boldsymbol{W}^{(t)}\big),$ (30b)
-where $L_{a}^{\text{lip}}$ is any Lipschitz constant on the Frobenius ball $\{\boldsymbol{W}:\|\boldsymbol{W}\|_{F}^{2}\leq P_{\mathrm{B}}\}$ for $\nabla g_{a}(\boldsymbol{W})$.
+where $L_{a}^{\text{lip}}$ is any Lipschitz constant on the Frobenius ball $\{\boldsymbol{W}:\|\boldsymbol{W}\|_{F}^{2}\leq P_{\text{B}}\}$ for $\nabla g_{a}(\boldsymbol{W})$.
 Using matrix-norm inequalities, one obtains a choice for a global Lipschitz constant
-$L_{a}^{\text{lip}}=12P_{\text{B}}\lVert\boldsymbol{S}_{a}\rVert_{2}^{2}\,+4\mathcal{P}_{\text{d},a}\lVert\boldsymbol{S}_{a}\rVert_{2},$ (30)
-with $\|\boldsymbol{S}_{a}\|_{2}=\|\boldsymbol{f}_{\mathrm{S},a}\|_{2}^{2}$.
-Proof: For any two points $\boldsymbol{W}_{1}$ and $\boldsymbol{W}_{2}$ on $g_{a}(\boldsymbol{W})$, we have
+$L_{a}^{\text{lip}}=12P_{\text{B}}\lVert\boldsymbol{S}_{a}\rVert_{2}^{2}+4\bar{\mathcal{P}}_{\text{d},a}\lVert\boldsymbol{S}_{a}\rVert_{2},$ (31)
+with $\|\boldsymbol{S}_{a}\|_{2}=\|\boldsymbol{f}_{\text{S},a}\|_{2}^{2}$.
+Proof:
+For any two points $\boldsymbol{W}_{1}$ and $\boldsymbol{W}_{2}$ on the function $g_{a}(\boldsymbol{W})$, according to matrix norm inequalities, we have
 $\lVert\nabla g_{a}\left(\boldsymbol{W}_{1}\right)-\nabla g_{a}\left(\boldsymbol{W}_{2}\right)\rVert$
-$=4\big\|(\text{Tr}(\boldsymbol{W}_{1}^{H}\boldsymbol{S}_{a}\boldsymbol{W}_{1})-\mathcal{P}_{\text{d},a})\boldsymbol{S}_{a}(\boldsymbol{W}_{1}-\boldsymbol{W}_{2})$
-$\quad\quad+(\text{Tr}(\boldsymbol{W}_{1}^{H}\boldsymbol{S}_{a}\boldsymbol{W}_{1})-\text{Tr}(\boldsymbol{W}_{2}^{H}\boldsymbol{S}_{a}\boldsymbol{W}_{2}))\boldsymbol{S}_{a}\boldsymbol{W}_{2}\big\|$ (31a)
-$\leq 4\lVert\boldsymbol{S}_{a}\rVert_{2}\big(|\text{Tr}(\boldsymbol{W}_{1}^{H}\boldsymbol{S}_{a}\boldsymbol{W}_{1})-\mathcal{P}_{\text{d},a}|\lVert\boldsymbol{W}_{1}-\boldsymbol{W}_{2}\rVert_{F}$
-$\quad\quad+|\text{Tr}(\boldsymbol{W}_{1}^{H}\boldsymbol{S}_{a}\boldsymbol{W}_{1})\!-\!\text{Tr}(\boldsymbol{W}_{2}^{H}\boldsymbol{S}_{a}\boldsymbol{W}_{2})|\lVert\boldsymbol{W}_{2}\rVert_{F}\big).\!\!\!$ (31b)
-Then, according to the following inequalities for terms in (31b)
-$|\text{Tr}(\boldsymbol{W}_{1}^{H}\boldsymbol{S}_{a}\boldsymbol{W}_{1}t)-\text{Tr}(\boldsymbol{W}_{2}^{H}\boldsymbol{S}_{a}\boldsymbol{W}_{2})|$
-$=|\text{Tr}((\boldsymbol{W}_{1}-\boldsymbol{W}_{2})^{H}\boldsymbol{S}_{a}\boldsymbol{W}_{1})-\text{Tr}(\boldsymbol{W}_{2}^{H}\boldsymbol{S}_{a}(\boldsymbol{W}_{1}-\boldsymbol{W}_{2}))|$
+$=4\big\|(\mathrm{Tr}(\boldsymbol{W}_{1}^{H}\boldsymbol{S}_{a}\boldsymbol{W}_{1})-\bar{\mathcal{P}}_{\text{d},a})\boldsymbol{S}_{a}(\boldsymbol{W}_{1}-\boldsymbol{W}_{2})$
+$\quad+(\mathrm{Tr}(\boldsymbol{W}_{1}^{H}\boldsymbol{S}_{a}\boldsymbol{W}_{1})-\mathrm{Tr}(\boldsymbol{W}_{2}^{H}\boldsymbol{S}_{a}\boldsymbol{W}_{2}))\boldsymbol{S}_{a}\boldsymbol{W}_{2}\big\|$
+$\leq 4\lVert\boldsymbol{S}_{a}\rVert_{2}\big(|\mathrm{Tr}(\boldsymbol{W}_{1}^{H}\boldsymbol{S}_{a}\boldsymbol{W}_{1})-\bar{\mathcal{P}}_{\text{d},a}|\lVert\boldsymbol{W}_{1}-\boldsymbol{W}_{2}\rVert_{F}$
+$\quad+|\mathrm{Tr}(\boldsymbol{W}_{1}^{H}\boldsymbol{S}_{a}\boldsymbol{W}_{1})\!-\!\mathrm{Tr}(\boldsymbol{W}_{2}^{H}\boldsymbol{S}_{a}\boldsymbol{W}_{2})|\lVert\boldsymbol{W}_{2}\rVert_{F}\big).\!$ (32)
+Then, we construct the following inequalities for terms related to traces in (III-B1)
+$|\mathrm{Tr}(\boldsymbol{W}_{1}^{H}\boldsymbol{S}_{a}\boldsymbol{W}_{1})-\mathrm{Tr}(\boldsymbol{W}_{2}^{H}\boldsymbol{S}_{a}\boldsymbol{W}_{2})|$
+$=|\mathrm{Tr}((\boldsymbol{W}_{1}-\boldsymbol{W}_{2})^{H}\boldsymbol{S}_{a}\boldsymbol{W}_{1})-\mathrm{Tr}(\boldsymbol{W}_{2}^{H}\boldsymbol{S}_{a}(\boldsymbol{W}_{1}-\boldsymbol{W}_{2}))|$
 $\leq\lVert\boldsymbol{W}_{1}-\boldsymbol{W}_{2}\rVert_{F}\lVert\boldsymbol{S}_{a}\rVert_{2}(\lVert\boldsymbol{W}_{1}\rVert_{F}+\lVert\boldsymbol{W}_{2}\rVert_{F}),$ (33)
-$|\text{Tr}(\boldsymbol{W}_{1}^{H}\boldsymbol{S}_{a}\boldsymbol{W}_{1})-\mathcal{P}_{\text{d},a}|\leq\lVert\boldsymbol{S}_{a}\rVert_{2}\lVert\boldsymbol{W}_{1}\rVert_{F}^{2}+\mathcal{P}_{\text{d},a},$ (33)
-$\lVert\boldsymbol{W}_{1}\rVert_{F},\lVert\boldsymbol{W}_{2}\rVert_{F}\leq\sqrt{P_{B}},$ (34)
-we substitute (32)-(34) into (31) and finally obtain
+$|\mathrm{Tr}(\boldsymbol{W}_{1}^{H}\boldsymbol{S}_{a}\boldsymbol{W}_{1})-\bar{\mathcal{P}}_{\text{d},a}|\leq\lVert\boldsymbol{S}_{a}\rVert_{2}\lVert\boldsymbol{W}_{1}\rVert_{F}^{2}+\bar{\mathcal{P}}_{\text{d},a},$ (34)
+where the power limitation holds
+$\lVert\boldsymbol{W}_{1}\rVert_{F}\leq\sqrt{P_{\text{B}}},\quad\lVert\boldsymbol{W}_{2}\rVert_{F}\leq\sqrt{P_{\text{B}}}.$ (35)
+Substituting (33), (34), and (35) into (III-B1), we obtain
 $\lVert\nabla g_{a}\left(\boldsymbol{W}_{1}\right)-\nabla g_{a}\left(\boldsymbol{W}_{2}\right)\rVert$
-$\leq 4\lVert\boldsymbol{S}_{a}\rVert_{2}\left(P_{B}\lVert\boldsymbol{S}_{a}\rVert_{2}+\mathcal{P}_{\text{d},a}\right)\lVert\boldsymbol{W}_{1}-\boldsymbol{W}_{2}\rVert_{F}$
-$\quad+4\lVert\boldsymbol{S}_{a}\rVert_{2}(2\sqrt{P_{B}}\lVert\boldsymbol{S}_{a}\rVert_{2}\lVert\boldsymbol{W}_{1}-\boldsymbol{W}_{2}\rVert_{F})\sqrt{P_{B}}$
-$=\left(12P_{B}\lVert\boldsymbol{S}_{a}\rVert_{2}^{2}+4\lVert\boldsymbol{S}_{a}\rVert_{2}\mathcal{P}_{\text{d},a}\right)\lVert\boldsymbol{W}_{1}-\boldsymbol{W}_{2}\rVert_{F},$ (35)
-which is exactly the definition of the Lipschitz constant.    $\blacksquare$
-Consequently, replacing each $g_{a}(\boldsymbol{W})$ by its upper bound yields a concave penalization term $\rho\max\{\mathcal{U}_{a}(\boldsymbol{W}\mid\boldsymbol{W}^{(t)})-\!\eta_{\text{th}},0\}$.
-At each iteration, (29) provides a tight lower bound on $\boldsymbol{W}^{(t)}$ to maximize the objective function (24), guaranteeing non-decreasing penalized objective values in updates to $\boldsymbol{W}$ with a standard MM step.
-III-B2 Communication SINR term with Lagrangian dual transform (LDT) and QT for fractions
+$\leq 4\lVert\boldsymbol{S}_{a}\rVert_{2}\left(P_{\text{B}}\lVert\boldsymbol{S}_{a}\rVert_{2}+\bar{\mathcal{P}}_{\text{d},a}\right)\lVert\boldsymbol{W}_{1}-\boldsymbol{W}_{2}\rVert_{F}$
+$\quad+4\lVert\boldsymbol{S}_{a}\rVert_{2}(2\sqrt{P_{\text{B}}}\lVert\boldsymbol{S}_{a}\rVert_{2}\lVert\boldsymbol{W}_{1}-\boldsymbol{W}_{2}\rVert_{F})\sqrt{P_{\text{B}}}$
+$=\left(12P_{\text{B}}\lVert\boldsymbol{S}_{a}\rVert_{2}^{2}+4\lVert\boldsymbol{S}_{a}\rVert_{2}\bar{\mathcal{P}}_{\text{d},a}\right)\lVert\boldsymbol{W}_{1}-\boldsymbol{W}_{2}\rVert_{F},$ (36)
+which is exactly the definition of the Lipschitz constant. ∎
+Consequently, at each iteration, replacing each $g_{a}(\boldsymbol{W})$ by its concave upper bound $\mathcal{U}_{a}(\boldsymbol{W}\mid\boldsymbol{W}^{(t)})$ on $\boldsymbol{W}^{(t)}$ in (30) provides a tight lower bound to maximize the objective function $\rho\text{NMSE}(\boldsymbol{W})$ of (25), guaranteeing non-decreasing objective values in updates to $\boldsymbol{W}$ with a standard $M$ step.
+III-B2 Communication SINR term with LDT and QT
 Introducing auxiliary real-valued variables $\{\mu_{k}>0,\forall k\}$, we remove the fractional term from the logarithmic function using the LDT [37]
@@
 $=\underset{\mu_{k}\geq 0}{\max}\,\log_{2}(1\!+\!\mu_{k})\!-\!\frac{\mu_{k}}{\ln 2}\!+\!\frac{1+\mu_{k}}{\ln 2}\frac{|\boldsymbol{f}_{k}^{H}\boldsymbol{w}_{k}|^{2}}{\sum_{i=1}^{K+M}{|\boldsymbol{f}_{k}^{H}\boldsymbol{w}_{i}|^{2}}\!+\!\sigma_{k}^{2}}.$ (37)
-Next, introducing complex-valued variables $\{y_{k},\forall k\}$, we apply QT to transform the fractional SINR term into
+Next, introducing complex-valued variables $\{\eta_{k},\forall k\}$, we apply QT to transform the fractional SINR term into
 $\frac{|\boldsymbol{f}_{k}^{H}\boldsymbol{w}_{k}|^{2}}{\sum_{i=1}^{K+M}{|\boldsymbol{f}_{k}^{H}\boldsymbol{w}_{i}|^{2}}+\sigma_{k}^{2}}$
-$=\underset{y_{k}\in\mathbb{C}}{\max}\,\,2\text{Re}\{y_{k}^{*}\boldsymbol{f}_{k}^{H}\boldsymbol{w}_{k}\}\!-\!\left|y_{k}\right|^{2}\Big(\!\sum_{i=1}^{K+M}\!{|\boldsymbol{f}_{k}^{H}\boldsymbol{w}_{i}|^{2}}\!+\!\sigma_{k}^{2}\Big).\!\!$ (37)
+$=\underset{\eta_{k}\in\mathbb{C}}{\max}\,\,2\text{Re}\{\eta_{k}^{*}\boldsymbol{f}_{k}^{H}\boldsymbol{w}_{k}\}\!-\!\left|\eta_{k}\right|^{2}\Big(\!\sum_{i=1}^{K+M}\!{|\boldsymbol{f}_{k}^{H}\boldsymbol{w}_{i}|^{2}}\!+\!\sigma_{k}^{2}\Big).\!\!$ (38)
 These auxiliary variables can be updated in closed form
 $\mu_{k}^{*}=\frac{|\boldsymbol{f}_{k}^{H}\boldsymbol{w}_{k}|^{2}}{\sum_{i=1,i\neq k}^{K+M}{|\boldsymbol{f}_{k}^{H}\boldsymbol{w}_{i}|^{2}}+\sigma_{k}^{2}},$ (39)
-$y_{k}^{*}=\frac{\boldsymbol{f}_{k}^{H}\boldsymbol{w}_{k}}{\sum_{i=1}^{K+M}{|\boldsymbol{f}_{k}^{H}\boldsymbol{w}_{i}|^{2}}+\sigma_{k}^{2}}.$ (39)
+$\eta_{k}^{*}=\frac{\boldsymbol{f}_{k}^{H}\boldsymbol{w}_{k}}{\sum_{i=1}^{K+M}{|\boldsymbol{f}_{k}^{H}\boldsymbol{w}_{i}|^{2}}+\sigma_{k}^{2}}.$ (40)
 Consequently, the communication sum rate is equivalently transformed into a non-fractional quadratic function of $\boldsymbol{W}$
-$\mathcal{Y}\left(\boldsymbol{W}\mid\mu_{k},y_{k}\right)$
-$\!=\!\frac{1\!\!+\!\mu_{k}}{\ln 2}\Big(2\text{Re}\{y_{k}^{*}\boldsymbol{f}_{k}^{H}\boldsymbol{w}_{k}\}\!-\!\left|y_{k}\right|^{2}\!\Big(\!\sum_{i=1}^{K+M}{\!\!}|\boldsymbol{f}_{k}^{H}\boldsymbol{w}_{i}|^{2}\!\!+\!\sigma_{k}^{2}\Big)\!\Big).\!\!\!$ (40)
+$\mathcal{Y}\left(\boldsymbol{W}\mid\mu_{k},\eta_{k}\right)$
+$\!=\!\frac{1\!\!+\!\mu_{k}}{\ln 2}\Big(2\text{Re}\{\eta_{k}^{*}\boldsymbol{f}_{k}^{H}\boldsymbol{w}_{k}\}\!-\!\left|\eta_{k}\right|^{2}\!\Big(\!\sum_{i=1}^{K+M}{\!\!}|\boldsymbol{f}_{k}^{H}\boldsymbol{w}_{i}|^{2}\!\!+\!\sigma_{k}^{2}\Big)\!\Big).\!\!\!$ (41)
 III-B3 Closed form solution for the QCQP
-With $\{\mu_{k},y_{k}\}$ fixed and $g_{a}$ replaced by $\mathcal{U}_{a}$, (P2) reduces to the convex quadratic constrained quadratic program (QCQP) stated as below
-$\text{(P2.1):}\underset{\boldsymbol{W}}{\max}\,$ $\sum_{k=1}^{K}{\mathcal{Y}\left(\boldsymbol{W}\mid\mu_{k},y_{k}\right)}-\rho\max\{\mathcal{U}_{a}(\boldsymbol{W}\mid\boldsymbol{W}^{(t)}),0\}$ (41)
-s.t. $\left(22\text{c}\right).$
+With $\{\mu_{k},\eta_{k}\}$ fixed and $g_{a}$ replaced by $\mathcal{U}_{a}$, (P2) reduces to the convex quadratic constrained quadratic program (QCQP) stated as below
+$\text{(P2.1):}\underset{\boldsymbol{W}}{\max}\,$ $\sum_{k=1}^{K}{\mathcal{Y}\left(\boldsymbol{W}\!\mid\!\mu_{k},\eta_{k}\right)}-\frac{\rho}{D}\sum_{a=1}^{A}{\mathcal{U}_{a}(\boldsymbol{W}\!\mid\!\boldsymbol{W}^{(t)})}$ (42)
+[…]
 In general, (P2.1) can be solved using convex optimization tools.
-However, we derive an efficient solution based on Lagrange duality and the Karush-Kuhn-Tucker (KKT) conditions.
-Specifically, if the MSE threshold is violated at the current $\boldsymbol{W}^{(t)}$, the objective function (41) can be compactly rewritten as a summation with a linear and a quadratic term
-$\underset{\boldsymbol{W}}{\max}\,$ $-\text{tr}(\boldsymbol{W}^{H}\boldsymbol{QW})\,\,+2\text{Re}\{\text{Tr}(\boldsymbol{P}^{H}\boldsymbol{W})\}\quad\text{s.t.}\,\,\text{(22c)},$ (42)
-where $\boldsymbol{Q}$ and $\boldsymbol{P}$ are defined by
+However, we derive a solution based on Lagrange duality and the Karush-Kuhn-Tucker (KKT) conditions.
+Specifically, at the current $\boldsymbol{W}^{(t)}$, the objective (42) can be compactly rewritten as a summation with a linear and a quadratic term
+$\underset{\boldsymbol{W}}{\max}\,\,$ $-\mathrm{Tr}(\boldsymbol{W}^{H}\boldsymbol{QW})+2\text{Re}\{\mathrm{Tr}(\boldsymbol{P}^{H}\boldsymbol{W})\}$ (43)
+[…]
+where we omit constant values that do not affect optimization, and the matrices $\boldsymbol{Q}$ and $\boldsymbol{P}$ are defined respectively by
 $\boldsymbol{Q}=\boldsymbol{Q}_{\text{C}}+\rho\boldsymbol{Q}_{\text{S}},$ (44a)
-$\boldsymbol{P}=\boldsymbol{P}_{C}+\rho\boldsymbol{P}_{\text{S}},$ (43b)
+$\boldsymbol{P}=\boldsymbol{P}_{\text{C}}+\rho\boldsymbol{P}_{\text{S}},$ (44b)
 with equivalent expressions from (30) and (41) as
-$\boldsymbol{Q}_{C}=\frac{1}{\ln 2}\sum_{k=1}^{K}{\left(1+\mu_{k}\right)\left|y_{k}\right|^{2}\boldsymbol{f}_{k}\boldsymbol{f}_{k}^{H}},$ (44a)
-$\boldsymbol{Q}_{\text{S}}=\frac{1}{2A}\sum_{a=1}^{A}{L_{a}^{\text{lip}}}\boldsymbol{I}_{M+K},$ (44b)
-$\boldsymbol{P}_{C}=\frac{1}{\ln 2}\left[\left(1+\mu_{1}\right)y_{1}\boldsymbol{f}_{1},\dots,\left(1+\mu_{K}\right)y_{K}\boldsymbol{f}_{K}\right],$ (44c)
-$\boldsymbol{P}_{\text{S}}=\frac{1}{2A}\sum_{a=1}^{A}{\boldsymbol{G}_{a}\left(\boldsymbol{W}^{\left(t\right)}\right)+L_{a}^{\text{lip}}\boldsymbol{W}^{\left(t\right)}}.$ (44d)
+$\boldsymbol{Q}_{\text{C}}=\frac{1}{\ln 2}\sum_{k=1}^{K}{\left(1+\mu_{k}\right)\left|\eta_{k}\right|^{2}\boldsymbol{f}_{k}\boldsymbol{f}_{k}^{H}},$ (45a)
+$\boldsymbol{Q}_{\text{S}}=\frac{1}{2D}\sum_{a=1}^{A}{L_{a}^{\text{lip}}}\boldsymbol{I}_{M},$ (45b)
+$\boldsymbol{P}_{\text{C}}=\frac{1}{\ln 2}\left[\left(1+\mu_{1}\right)\eta_{1}\boldsymbol{f}_{1},\dots,\left(1+\mu_{K}\right)\eta_{K}\boldsymbol{f}_{K}\right],$ (45c)
+$\boldsymbol{P}_{\text{S}}=\frac{1}{2D}\sum_{a=1}^{A}{L_{a}^{\text{lip}}\boldsymbol{W}^{\left(t\right)}-\boldsymbol{G}_{a}\big(\boldsymbol{W}^{(t)}\big)}.$ (45d)
 The corresponding Lagrangian with a multiplier $\nu>0$ is
-$\mathcal{L}(\boldsymbol{W},\nu)=$ $-\text{tr}(\boldsymbol{W}^{H}\boldsymbol{QW})+2\text{Re}\{\text{tr}(\boldsymbol{P}^{H}\boldsymbol{W})t\}$
+$\mathcal{L}(\boldsymbol{W},\nu)=$ $-\mathrm{Tr}(\boldsymbol{W}^{H}\boldsymbol{QW})+2\text{Re}\{\mathrm{Tr}(\boldsymbol{P}^{H}\boldsymbol{W})\}$
 $-\nu(\lVert\boldsymbol{W}\rVert_{F}^{2}-P_{\text{B}}),$ (46)
 with KKT conditions for stationarity, complementary slackness, primal feasibility, and dual feasibility, respectively, being
-$-\boldsymbol{QW}^{*}+\boldsymbol{P}-\nu^{*}\boldsymbol{I}=0,$ (46a)
+$-(\boldsymbol{Q}+\nu^{*}\boldsymbol{I})\boldsymbol{W}^{*}+\boldsymbol{P}=0,$ (47a)
 $\nu^{*}\left(\lVert\boldsymbol{W}^{*}\rVert_{F}^{2}-P_{\text{B}}\right)=0,$ (47b)
@@
 $\boldsymbol{W}\left(\nu\right)=\left(\boldsymbol{Q}+\nu\boldsymbol{I}\right)^{-1}\boldsymbol{P},\nu\geq 0,$ (48)
-where $\nu$ is found by bisection search to satisfy the power constraint (22c) with equality.
-Since $\nu$ enters as a scalar shift on the identity, we precompute the eigenvalue decomposition (EVD) $\boldsymbol{Q}=\boldsymbol{U\varLambda U}^{H}$ with $\boldsymbol{\varLambda}=\text{diag}(\lambda_{1},\dots,\lambda_{M})$ and $\lambda_{m}$ being the eigenvalue; thus, for $\nu>-\min\{\lambda_{m}\}_{m=1}^{M}$, we have
+where $\nu$ is found by bisection search to satisfy the power constraint (24b) with equality.
+Since $\nu$ enters as a scalar shift on the identity, we precompute the eigenvalue decomposition (EVD) $\boldsymbol{Q}=\boldsymbol{U\varLambda U}^{H}$ with $\boldsymbol{\varLambda}=\mathrm{diag}(\lambda_{1},\dots,\lambda_{M})$ and $\lambda_{m}$ being the eigenvalue; thus, for $\nu>-\min\{\lambda_{m}\}_{m=1}^{M}$, we have
 $\boldsymbol{W}\left(\nu\right)$ $=\boldsymbol{U}\left(\boldsymbol{\varLambda}+\nu\boldsymbol{I}\right)^{-1}\boldsymbol{U}^{H}\boldsymbol{P}$ (49a)
@@
 where $\cdot]_{i,:}$ denote the $i$-th row of a matrix.
-Furthermore, if the MSE threshold is satisfied in current $\boldsymbol{W}^{(t)}$, we let $\boldsymbol{Q}=\boldsymbol{Q}_{\text{C}}$ and $\boldsymbol{P}=\boldsymbol{P}_{C}$ and perform the same procedures in (43)-(49).
 III-B4 Overall algorithm and complexity analysis for subproblem 1
@@
 Each inner iteration performs one EVD of an $M{\times}M$ Hermitian matrix at cost $\mathcal{O}(M^{3})$ and a scalar bisection whose per-step cost is $\mathcal{O}(M(K+M))$ to calculate $\|\boldsymbol{W}(\nu)\|_{F}^{2}$; the number of bisection steps grows only logarithmically with precision.
-The MM construction provides a tight lower bound of the penalized objective; therefore, the penalized objective in (P2) is non-decreasing over $\boldsymbol{W}$-updates.
+The MM construction provides a tight lower bound of the objective in (P2), which is non-decreasing over $\boldsymbol{W}$-updates.
 Under the power ball, the sequence $\{\boldsymbol{W}^{(t)}\}$ admits limit points that satisfy the first-order stationarity conditions of the surrogate built from MM and QT.
 Algorithm 1 Low-complexity transmit beamforming based on QT, $M$, and bisection search for (P2)
-1:  Input: $\{\boldsymbol{f}_{k}\}_{k=1}^{K}$, $\{\boldsymbol{f}_{\mathrm{S},a}\}_{a=1}^{A}$, $\{\mathcal{P}_{\mathrm{d},a}\}$, $P_{\mathrm{B}}$, $\rho$, $\eta_{\mathrm{th}}$.
+1:  Input: $\{\boldsymbol{f}_{k}\}_{k=1}^{K}$, $\{\boldsymbol{f}_{\text{S},a}\}_{a=1}^{A}$, $\{\mathcal{P}_{\text{d},a}\}$, $P_{\text{B}}$, $\rho$.
 2:  Initialization: $\boldsymbol{W}^{(0)}$, using zero forcing in the first outer iterate, or using previous outer iterate results; Set $t=0$.
 3:  repeat
-4:   Compute $\|\boldsymbol{S}_{a}\|_{2}^{2}$ via (27), $L_{a}^{\text{lip}}$ via (30), $\boldsymbol{G}_{a}(\boldsymbol{W}^{(t)})$ via (28), and $\mathcal{U}(\boldsymbol{W}\mid\boldsymbol{W}^{(t)})$ via (29).
-5:   Build $\boldsymbol{Q}_{\mathrm{C}},\boldsymbol{P}_{\mathrm{C}}$ from (44); if $\mathcal{U}(\boldsymbol{W}^{(t)}\mid\boldsymbol{W}^{(t)})>\eta_{\mathrm{th}}$, also build $\boldsymbol{Q}_{\mathrm{S}},\boldsymbol{P}_{\mathrm{S}}$, else set them to zero.
-Set $\boldsymbol{Q}$ and $\boldsymbol{P}$ via (43).
-6:   Compute EVD $\boldsymbol{Q}=\boldsymbol{U}\boldsymbol{\Lambda}\boldsymbol{U}^{H}$. Find $\nu^{\star}$ by bisection using (45); set $\boldsymbol{W}^{(t{+}1)}=\boldsymbol{W}(\nu^{\star})$ via (48).
-7:   Update auxiliaries $\mu_{k}^{(t{+}1)}$ by (38) and $y_{k}^{(t{+}1)}$ by(39).
+4:   Compute $\|\boldsymbol{S}_{a}\|_{2}^{2}$ via (28), $L_{a}^{\text{lip}}$ via (31), $\boldsymbol{G}_{a}(\boldsymbol{W}^{(t)})$ via (29a), and $\mathcal{U}(\boldsymbol{W}\mid\boldsymbol{W}^{(t)})$ via (30).
+5:   Build $\boldsymbol{Q}_{\text{C}}$, $\boldsymbol{P}_{\text{C}}$, $\boldsymbol{Q}_{\text{S}}$, and $\boldsymbol{P}_{\text{S}}$ from (45). Set $\boldsymbol{Q}$ and $\boldsymbol{P}$ via (44).
+6:   Perform EVD by $\boldsymbol{Q}=\boldsymbol{U}\boldsymbol{\Lambda}\boldsymbol{U}^{H}$.
+Find $\nu^{\star}$ by bisection using (49) and (50); set $\boldsymbol{W}^{(t{+}1)}=\boldsymbol{W}(\nu^{\star})$ via (48).
+7:   Update auxiliaries $\mu_{k}^{(t{+}1)}$ by (39) and $\eta_{k}^{(t{+}1)}$ by (40).
 8:   $t\leftarrow t{+}1$.
@@
 Given fixed $\boldsymbol{W}$ and $(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}})$, we optimize $\boldsymbol{\theta}$ via
-$\text{(P3)}:\underset{\boldsymbol{\theta}}{\max}\,\,$ $\sum_{k=1}^{K}{\log_{2}\left(1+\text{SINR}_{k}(\boldsymbol{\theta})\right)}$
-$-\rho\max\left\{\text{MSE}(\boldsymbol{\theta})-\eta_{\text{th}},0\right\}$ (50)
-s.t. $\left(22\text{d}\right).$
-The unique challenge in RIS passive beamforming is the unit-modulus constraint (22d). Nevertheless, it inherently defines a Riemannian manifold, known as the complex circle manifold
+$\text{(P3)}:\underset{\boldsymbol{\theta}}{\max}\,\,$ $\sum_{k=1}^{K}{\log_{2}\left(1+\text{SINR}_{k}(\boldsymbol{\theta})\right)}-\rho\text{NMSE}(\boldsymbol{\theta})$ (51)
+[…]
+The unique challenge in RIS passive beamforming is the unit-modulus constraint. Nevertheless, it inherently defines a Riemannian manifold, known as the complex circle manifold
 $\mathcal{R}_{\boldsymbol{\theta}}=\left\{\boldsymbol{\theta}\in\mathbb{C}^{N}:|\theta_{n}|=1,\forall n\in\mathcal{N}\right\},$ (52)
@@
 $\mathcal{P}_{a}(\vartheta,\varphi;\boldsymbol{\theta})$ $=\boldsymbol{\theta}^{H}\boldsymbol{\Xi}_{\text{S},a}\boldsymbol{\theta}+2\Re\{\boldsymbol{\xi}_{\text{S},a}^{H}\boldsymbol{\theta}\}+\xi_{\text{S},a},$ (53)
-where $\boldsymbol{\Xi}_{\text{S},a}$, $\boldsymbol{\xi}_{\text{S},a}$, and $\xi_{\text{S},a}$ are computed based on (14) by
-$\!\boldsymbol{\Xi}_{\mathrm{S},a}=G^{\mathrm{R}}(\boldsymbol{r}^{\mathrm{R}})\mathrm{diag}(\boldsymbol{t}_{a}^{\mathrm{RT}})\boldsymbol{B}^{H}\boldsymbol{W}\boldsymbol{W}^{H}\boldsymbol{B}\mathrm{diag}(\boldsymbol{t}_{a}^{\mathrm{RT}})^{H},\!\!$ (53a)
-$\!\boldsymbol{\xi}_{\mathrm{S},a}^{H}=\!\sqrt{G^{\mathrm{B}}(\boldsymbol{r}^{\mathrm{B}})G^{\mathrm{R}}(\boldsymbol{r}^{\mathrm{R}})}(\boldsymbol{t}_{a}^{\mathrm{BT}})^{H}\boldsymbol{W}\boldsymbol{W}^{H}\boldsymbol{B}\mathrm{diag}(\boldsymbol{t}_{a}^{\mathrm{RT}}),\!\!\!$ (53b)
-$\!\xi_{\mathrm{S},a}=G^{\mathrm{B}}(\boldsymbol{r}^{\mathrm{B}})(\boldsymbol{t}_{a}^{\mathrm{BT}})^{H}\boldsymbol{W}\boldsymbol{W}^{H}\boldsymbol{t}_{a}^{\mathrm{BT}}.\!\!$ (53c)
+where $\boldsymbol{\Xi}_{\text{S},a}$, $\boldsymbol{\xi}_{\text{S},a}$, and $\xi_{\text{S},a}$ are computed based on (II-C) by
+$\!\boldsymbol{\Xi}_{\mathrm{S},a}=G^{\text{R}}(\boldsymbol{r}^{\text{R}})\mathrm{diag}(\boldsymbol{t}_{a}^{\mathrm{RT}})\boldsymbol{B}^{H}\boldsymbol{W}\boldsymbol{W}^{H}\boldsymbol{B}\mathrm{diag}(\boldsymbol{t}_{a}^{\mathrm{RT}})^{H},\!\!$ (54a)
+$\!\boldsymbol{\xi}_{\mathrm{S},a}^{H}=\!\sqrt{G^{\text{B}}(\boldsymbol{r}^{\text{B}})G^{\text{R}}(\boldsymbol{r}^{\text{R}})}(\boldsymbol{t}_{a}^{\mathrm{BT}})^{H}\boldsymbol{W}\boldsymbol{W}^{H}\boldsymbol{B}\mathrm{diag}(\boldsymbol{t}_{a}^{\mathrm{RT}}),\!\!\!$ (54b)
+$\!\xi_{\mathrm{S},a}=G^{\text{B}}(\boldsymbol{r}^{\text{B}})(\boldsymbol{t}_{a}^{\mathrm{BT}})^{H}\boldsymbol{W}\boldsymbol{W}^{H}\boldsymbol{t}_{a}^{\mathrm{BT}}.\!\!$ (54c)
 Similarly, the signal or interference term of any beam $\boldsymbol{w}_{i}$ for the $k$-th communication user in the SINR expression with effective downlink channel $\boldsymbol{f}_{k}(\boldsymbol{\theta})$ can be written as
@@
 Define real-valued scalars $S_{a}(\boldsymbol{\theta})$, $C_{1,k}(\boldsymbol{\theta})$, and $C_{2,k}(\boldsymbol{\theta})$ as
-$\!\!\!S_{a}(\boldsymbol{\theta})$ $=\boldsymbol{\theta}^{H}\boldsymbol{\Xi}_{\mathrm{S},a}\boldsymbol{\theta}+2\Re\{\boldsymbol{\xi}_{\mathrm{S},a}^{H}\boldsymbol{\theta}\}+\xi_{\mathrm{S},a}-\mathcal{P}_{\mathrm{d},a},$ (56a)
+$\!\!\!S_{a}(\boldsymbol{\theta})$ $=\boldsymbol{\theta}^{H}\boldsymbol{\Xi}_{\mathrm{S},a}\boldsymbol{\theta}+2\Re\{\boldsymbol{\xi}_{\mathrm{S},a}^{H}\boldsymbol{\theta}\}+\xi_{\mathrm{S},a}-\bar{\mathcal{P}}_{\mathrm{d},a},$ (57a)
 $\!\!\!C_{1,k}(\boldsymbol{\theta})$ $=\boldsymbol{\theta}^{H}\boldsymbol{\Xi}_{\mathrm{C},k,k}\boldsymbol{\theta}+2\Re\{\boldsymbol{\xi}_{\mathrm{C},k,k}^{H}\boldsymbol{\theta}\}+\xi_{\mathrm{C},k,k},$ (57b)
@@
 we can rewrite the objective function in (51) as
-$F\left(\boldsymbol{\theta}\right)=$ $\sum_{k=1}^{K}{\log_{2}}\Big(1+\frac{C_{1,k}\left(\boldsymbol{\theta}\right)}{C_{2,k}\left(\boldsymbol{\theta}\right)}\Big)$
-$-\rho\max\Big\{\frac{1}{A}\sum_{a=1}^{A}{|S_{a}\left(\boldsymbol{\theta}\right)|^{2}}-\eta_{\text{th}},0\Big\},$ (57)
+$F(\boldsymbol{\theta})=\sum_{k=1}^{K}\log_{2}\!\left(1+\frac{C_{1,k}(\boldsymbol{\theta})}{C_{2,k}(\boldsymbol{\theta})}\right)-\frac{\rho}{D}\sum_{a=1}^{A}\bigl|S_{a}(\boldsymbol{\theta})\bigr|^{2},$ (58)
 which further arrives at a standard Riemannian unconstrained manifold optimization problem stated as
@@
 III-C2 RCG method
-To solve (P3.1), we adopt the Riemannian conjugate gradient (RCG) method, which generalizes the classical conjugate gradient algorithm over curved spaces, as a manifold is a topological space that locally resembles the Euclidean space [28],[29].
+To solve (P3.1), we adopt the RCG method, which generalizes the classical conjugate gradient algorithm over curved spaces, as a manifold is a topological space that locally resembles the Euclidean space [38, 39].
 At any point, tangent vectors give all feasible updating directions, and their set is the tangent space
@@
 $\nabla_{\boldsymbol{\theta}}C_{2,k}(\boldsymbol{\theta})$ $=\sum_{i=1,\,i\neq k}^{K+M}\big(2\boldsymbol{\Xi}_{\mathrm{C},k,i}\boldsymbol{\theta}+2\boldsymbol{\xi}_{\mathrm{C},k,i}\big).$ (62)
-Using the chain rule, the complete Euclidean gradient expression $\nabla_{\boldsymbol{\theta}}F(\boldsymbol{\theta})$ can be computed by (62) at the top of the next page, which is suitable for the case that the MSE threshold does not satisfy in the current $\boldsymbol{\theta}^{(\jmath)}$; otherwise, omit the MSE-related terms in (62).
-$\nabla_{\boldsymbol{\theta}}F\left(\boldsymbol{\theta}\right)=\frac{1}{\ln 2}\sum_{k=1}^{K}{\frac{C_{2,k}\left(\boldsymbol{\theta}\right)\,\nabla_{\boldsymbol{\theta}}C_{1,k}\left(\boldsymbol{\theta}\right)-C_{1,k}\left(\boldsymbol{\theta}\right)\,\nabla_{\boldsymbol{\theta}}C_{2,k}\left(\boldsymbol{\theta}\right)}{C_{2,k}\left(\boldsymbol{\theta}\right)\left(C_{1,k}\left(\boldsymbol{\theta}\right)+C_{2,k}\left(\boldsymbol{\theta}\right)\right)}}-\frac{\rho}{A}\sum_{a=1}^{A}{2}S_{a}\left(\boldsymbol{\theta}\right)\,\nabla_{\boldsymbol{\theta}}S_{a}\left(\boldsymbol{\theta}\right)$ (62a)
-$\qquad\,\,\,\quad=\frac{2}{\ln 2}\sum_{k=1}^{K}{\frac{C_{2,k}\left(\mathbf{\Xi}_{\text{C},k,k}\boldsymbol{\theta}+\boldsymbol{\xi}_{\text{C},k,k}\right)-C_{1,k}\sum_{i=1,i\neq k}^{K}{\mathbf{\Xi}_{\text{C},k,i}\boldsymbol{\theta}+\boldsymbol{\xi}_{\text{C},k,i}}}{C_{2,k}\left(C_{1,k}+C_{2,k}\right)}}-\frac{4\rho}{A}\sum_{a=1}^{A}{S_{a}}\left(\boldsymbol{\theta}\right)\left(\mathbf{\Xi}_{\text{S},a}\boldsymbol{\theta}+\boldsymbol{\xi}_{\text{S},a}\right)$ (62b)
+Using the chain rule, the complete Euclidean gradient expression $\nabla_{\boldsymbol{\theta}}F(\boldsymbol{\theta})$ can be computed by (63) at the top of the next page.
+$\nabla_{\boldsymbol{\theta}}F\left(\boldsymbol{\theta}\right)=\frac{1}{\ln 2}\sum_{k=1}^{K}{\frac{C_{2,k}\left(\boldsymbol{\theta}\right)\,\nabla_{\boldsymbol{\theta}}C_{1,k}\left(\boldsymbol{\theta}\right)-C_{1,k}\left(\boldsymbol{\theta}\right)\,\nabla_{\boldsymbol{\theta}}C_{2,k}\left(\boldsymbol{\theta}\right)}{C_{2,k}\left(\boldsymbol{\theta}\right)\left(C_{1,k}\left(\boldsymbol{\theta}\right)+C_{2,k}\left(\boldsymbol{\theta}\right)\right)}}-\frac{\rho}{D}\sum_{a=1}^{A}{2}S_{a}\left(\boldsymbol{\theta}\right)\,\nabla_{\boldsymbol{\theta}}S_{a}\left(\boldsymbol{\theta}\right)$ (63a)
+$\qquad\,\,\,\quad=\frac{2}{\ln 2}\sum_{k=1}^{K}{\frac{C_{2,k}\left(\mathbf{\Xi}_{\text{C},k,k}\boldsymbol{\theta}+\boldsymbol{\xi}_{\text{C},k,k}\right)-C_{1,k}\sum_{i=1,i\neq k}^{K}{\mathbf{\Xi}_{\text{C},k,i}\boldsymbol{\theta}+\boldsymbol{\xi}_{\text{C},k,i}}}{C_{2,k}\left(C_{1,k}+C_{2,k}\right)}}-\frac{4\rho}{D}\sum_{a=1}^{A}{S_{a}}\left(\boldsymbol{\theta}\right)\left(\mathbf{\Xi}_{\text{S},a}\boldsymbol{\theta}+\boldsymbol{\xi}_{\text{S},a}\right)$ (63b)
 $\overline${\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ }
@@
 where $\left\langle\cdot,\cdot\right\rangle$ denotes the Riemannian metric, being Euclidean inner product here. The descent direction is then updated as
-$\boldsymbol{\tau}^{(\jmath)}=-\nabla_{\mathcal{R}_{\boldsymbol{\theta}}}F^{\left(\jmath\right)}+\beta_{\boldsymbol{\theta}}^{(\jmath)}\mathsf{Proj}_{\boldsymbol{\theta}^{(\jmath)}}(\boldsymbol{\tau}^{(\jmath-1)}),$ (65)
+$\boldsymbol{\tau}^{(\jmath)}=-\nabla_{\mathcal{R}_{\boldsymbol{\theta}}}F^{\left(\jmath\right)}+\chi^{(\jmath)}\mathsf{Proj}_{\boldsymbol{\theta}^{(\jmath)}}(\boldsymbol{\tau}^{(\jmath-1)}),$ (66)
 here the projection represents a simple vector transport operation.
@@
 3:repeat
-4:  Compute $\nabla_{\boldsymbol{\theta}}F(\boldsymbol{\theta}^{(\jmath)})$ via (62) and $\mathsf{Rgrad}F(\boldsymbol{\theta})$ via (63).
+4:  Compute $\nabla_{\boldsymbol{\theta}}F(\boldsymbol{\theta}^{(\jmath)})$ via (63)) and $\mathsf{Rgrad}F(\boldsymbol{\theta})$ via (64).
 5:  Update $\chi^{(\jmath)}$ via (65) and search direction $\boldsymbol{\tau}^{(\jmath)}$ via (66).
@@
 These costs can be further reduced by avoiding explicit $N\times N$ formations and reusing the rank–1 decomposition for $\boldsymbol{B}^{H}\boldsymbol{W}\boldsymbol{W}^{H}\boldsymbol{B}$ in (54a).
-III-D Subproblem 3: Array Rotation Angles
-Given $\boldsymbol{W}$ and $\boldsymbol{\theta}$ from the previous subproblems, we jointly optimize the array rotations $(\boldsymbol{r}^{\mathrm{B}},\boldsymbol{r}^{\mathrm{R}})$ of the BS and RIS via
-$\text{(P4)}:\underset{\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}}}{\max}$ $\sum_{k=1}^{K}{\log_{2}\left(1+\text{SINR}_{k}\left(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}}\right)\right)}$
-$-\rho\max\{\text{MSE}\left(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}}\right)-\eta_{\text{th}},0\}\!\!\!$ (67)
-s.t. $\left(22\text{e}\right),\left(22\text{f}\right)$
+III-D Subproblem 3: Array Rotations
+Given $\boldsymbol{W}$ and $\boldsymbol{\theta}$ from the previous subproblems, we jointly optimize the array rotations $(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}})$ of the BS and RIS via
+$\text{(P4)}:\underset{\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}}}{\max}$ $\sum_{k=1}^{K}{\log_{2}\left(1+\text{SINR}_{k}\left(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}}\right)\right)}-\rho\text{NMSE}(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}})\!\!\!$ (68)
+[…]
 Due to the highly non-concave relationships involving the cosine and exponential functions in the directivity gain and steering vectors, respectively, in the objective function, and the relatively simple box constraints on the array rotation angles, we employ the projection gradient method (PGD) [40], an effective strategy for this nonlinear maximization problem with boundary value constraints.
@@
 The key step of PGD is to derive the gradient expression for the complex objective function (68).
-In the following, we consider the case that the rotations $(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}})^{(\ell)}$ in the current iteration violate the MSE threshold constraint.
-Thus, the MSE term is involved in the cost function; otherwise, the cost and gradient functions relate only to the communication sum rate.
 We adopt $\boldsymbol{r}^{\text{A}}$ for any array rotation variable with $\text{A}\in\{\text{B,R}\}$ for brevity as defined in (2).
-Further, define a six-dimensional vector $\boldsymbol{r}$ for all rotations as below
-$\boldsymbol{r}\triangleq[(\boldsymbol{r}^{\text{B}})^{T},(\boldsymbol{r}^{\text{R}})^{T}]^{T}=\left[r_{x}^{\text{B}},r_{y}^{\text{B}},r_{z}^{\text{B}},r_{x}^{\text{R}},r_{y}^{\text{R}},r_{z}^{\text{R}}\right]^{T}\in\mathbb{R}^{6}.$ (68)
-The gradient of (67) w.r.t. the generic rotation vector $\boldsymbol{r}$ is
-$\frac{\partial F\left(\boldsymbol{r}\right)}{\partial\boldsymbol{r}}=$ $-\frac{2\rho}{A}\sum_{a=1}^{A}{\left(\mathcal{P}_{a}\left(\boldsymbol{r}\right)-\mathcal{P}_{\text{d},a}\right)\frac{\partial\mathcal{P}_{a}\left(\boldsymbol{r}\right)}{\partial\boldsymbol{r}}}$
+Further, define a stacked six-dimensional vector $\boldsymbol{r}$ for all rotations as below
+$\boldsymbol{r}\triangleq[(\boldsymbol{r}^{\text{B}})^{T},(\boldsymbol{r}^{\text{R}})^{T}]^{T}=\left[r_{\mathrm{x}}^{\text{B}},r_{\mathrm{y}}^{\text{B}},r_{\mathrm{z}}^{\text{B}},r_{\mathrm{x}}^{\text{R}},r_{\mathrm{y}}^{\text{R}},r_{\mathrm{z}}^{\text{R}}\right]^{T}\in\mathbb{R}^{6}.$ (69)
+The gradient of (68) w.r.t. the stacked rotation vector $\boldsymbol{r}$ is
+$\frac{\partial F\left(\boldsymbol{r}\right)}{\partial\boldsymbol{r}}=$ $-\frac{2\rho}{D}\sum_{a=1}^{A}{\left(\mathcal{P}_{a}\left(\boldsymbol{r}\right)-\bar{\mathcal{P}}_{\text{d},a}\right)\frac{\partial\mathcal{P}_{a}\left(\boldsymbol{r}\right)}{\partial\boldsymbol{r}}}$
 $+\frac{1}{\ln 2}\sum_{k=1}^{K}{\frac{1}{1+\text{SINR}_{k}\left(\boldsymbol{r}\right)}\frac{\partial\text{SINR}_{k}\left(\boldsymbol{r}\right)}{\partial\boldsymbol{r}}},$ (70)
@@
 $\omega_{v}^{\text{X}}(\boldsymbol{r}^{\text{A}})=j\frac{2\pi}{\lambda}\left(\boldsymbol{u}^{\text{X}}\right)^{T}\frac{\partial\boldsymbol{d}_{v}^{\text{A}}\left(\boldsymbol{r}^{\text{A}}\right)}{\partial\boldsymbol{r}^{\text{A}}},$ (75b)
-$\frac{\partial\boldsymbol{d}_{v}^{\text{A}}(\boldsymbol{r}^{\text{A}})}{\partial\boldsymbol{r}^{\text{A}}}\!=\!\Big[\frac{\partial\boldsymbol{R}(\boldsymbol{r}^{\text{A}})}{\partial r_{x}^{\text{A}}}\overline{\boldsymbol{d}}_{v}^{\text{A}},\frac{\partial\boldsymbol{R}(\boldsymbol{r}^{\text{A}})}{\partial r_{y}^{\text{A}}}\overline{\boldsymbol{d}}_{v}^{\text{A}},\frac{\partial\boldsymbol{R}(\boldsymbol{r}^{\text{A}})}{\partial r_{z}^{\text{A}}}\overline{\boldsymbol{d}}_{v}^{\text{A}}\Big].\!\!\!$ (74c)
-The derivatives of the rotation matrix $\boldsymbol{R}\left(\boldsymbol{r}^{\text{A}}\right)=\boldsymbol{R}_{x}\left(r_{x}^{\text{A}}\right)\boldsymbol{R}_{y}\left(r_{y}^{\text{A}}\right)\boldsymbol{R}_{z}\left(r_{z}^{\text{A}}\right)$ w.r.t each rotation variable are
-$\frac{\partial\boldsymbol{R}(\boldsymbol{r}^{\text{A}})}{\partial r_{x}^{\text{A}}}=\frac{\partial\boldsymbol{R}_{x}(r_{x}^{\text{A}})}{\partial r_{x}^{\text{A}}}\boldsymbol{R}_{y}(r_{y}^{\text{A}})\boldsymbol{R}_{z}(r_{z}^{\text{A}}),$ (75a)
-$\frac{\partial\boldsymbol{R}(\boldsymbol{r}^{\text{A}})}{\partial r_{y}^{\text{A}}}=\boldsymbol{R}_{x}(r_{x}^{\text{A}})\frac{\partial\boldsymbol{R}_{y}(r_{y}^{\text{A}})}{\partial r_{y}^{\text{A}}}\boldsymbol{R}_{z}(r_{z}^{\text{A}}),$ (75b)
-$\frac{\partial\boldsymbol{R}(\boldsymbol{r}^{\text{A}})}{\partial r_{z}^{\text{A}}}=\boldsymbol{R}_{x}(r_{x}^{\text{A}})\boldsymbol{R}_{y}(r_{y}^{\text{A}})\frac{\partial\boldsymbol{R}_{z}(r_{z}^{\text{A}})}{\partial r_{z}^{\text{A}}},$ (75c)
+$\frac{\partial\boldsymbol{d}_{v}^{\text{A}}(\boldsymbol{r}^{\text{A}})}{\partial\boldsymbol{r}^{\text{A}}}\!=\!\Big[\frac{\partial\boldsymbol{R}(\boldsymbol{r}^{\text{A}})}{\partial r_{\mathrm{x}}^{\text{A}}}\overline{\boldsymbol{d}}_{v}^{\text{A}},\frac{\partial\boldsymbol{R}(\boldsymbol{r}^{\text{A}})}{\partial r_{\mathrm{y}}^{\text{A}}}\overline{\boldsymbol{d}}_{v}^{\text{A}},\frac{\partial\boldsymbol{R}(\boldsymbol{r}^{\text{A}})}{\partial r_{\mathrm{z}}^{\text{A}}}\overline{\boldsymbol{d}}_{v}^{\text{A}}\Big].\!\!\!$ (75c)
+The derivatives of the rotation matrix $\boldsymbol{R}\left(\boldsymbol{r}^{\text{A}}\right)=\boldsymbol{R}_{\mathrm{x}}\left(r_{\mathrm{x}}^{\text{A}}\right)\boldsymbol{R}_{\mathrm{y}}\left(r_{\mathrm{y}}^{\text{A}}\right)\boldsymbol{R}_{\mathrm{z}}\left(r_{\mathrm{z}}^{\text{A}}\right)$ w.r.t each rotation variable are
+$\frac{\partial\boldsymbol{R}(\boldsymbol{r}^{\text{A}})}{\partial r_{\mathrm{x}}^{\text{A}}}=\frac{\partial\boldsymbol{R}_{\mathrm{x}}(r_{\mathrm{x}}^{\text{A}})}{\partial r_{\mathrm{x}}^{\text{A}}}\boldsymbol{R}_{\mathrm{y}}(r_{\mathrm{y}}^{\text{A}})\boldsymbol{R}_{\mathrm{z}}(r_{\mathrm{z}}^{\text{A}}),$ (76a)
+$\frac{\partial\boldsymbol{R}(\boldsymbol{r}^{\text{A}})}{\partial r_{\mathrm{y}}^{\text{A}}}=\boldsymbol{R}_{\mathrm{x}}(r_{\mathrm{x}}^{\text{A}})\frac{\partial\boldsymbol{R}_{\mathrm{y}}(r_{\mathrm{y}}^{\text{A}})}{\partial r_{\mathrm{y}}^{\text{A}}}\boldsymbol{R}_{\mathrm{z}}(r_{\mathrm{z}}^{\text{A}}),$ (76b)
+$\frac{\partial\boldsymbol{R}(\boldsymbol{r}^{\text{A}})}{\partial r_{\mathrm{z}}^{\text{A}}}=\boldsymbol{R}_{\mathrm{x}}(r_{\mathrm{x}}^{\text{A}})\boldsymbol{R}_{\mathrm{y}}(r_{\mathrm{y}}^{\text{A}})\frac{\partial\boldsymbol{R}_{\mathrm{z}}(r_{\mathrm{z}}^{\text{A}})}{\partial r_{\mathrm{z}}^{\text{A}}},$ (76c)
 and according to the definition in (3), we have
-$\frac{\partial\boldsymbol{R}_{x}\left(r_{x}^{\text{A}}\right)}{\partial r_{x}^{\text{A}}}=\left[\begin{matrix}0&0&0\\ 0&-\sin r_{x}^{\text{A}}&-\cos r_{x}^{\text{A}}\\ 0&\cos r_{x}^{\text{A}}&-\sin r_{x}^{\text{A}}\\ \end{matrix}\right],$ (76a)
-$\frac{\partial\boldsymbol{R}_{y}\left(r_{y}^{\text{A}}\right)}{\partial r_{y}^{\text{A}}}=\left[\begin{matrix}-\sin r_{y}^{\text{A}}&0&\cos r_{y}^{\text{A}}\\ 0&0&0\\ -\cos r_{y}^{\text{A}}&0&-\sin r_{y}^{\text{A}}\\ \end{matrix}\right],$ (76b)
-$\frac{\partial\boldsymbol{R}_{z}\left(r_{z}^{\text{A}}\right)}{\partial r_{z}^{\text{A}}}=\left[\begin{matrix}-\sin r_{z}^{\text{A}}&-\cos r_{z}^{\text{A}}&0\\ \cos r_{z}^{\text{A}}&-\sin r_{z}^{\text{A}}&0\\ 0&0&0\\ \end{matrix}\right].$ (76c)
+$\frac{\partial\boldsymbol{R}_{\mathrm{x}}\left(r_{\mathrm{x}}^{\text{A}}\right)}{\partial r_{\mathrm{x}}^{\text{A}}}=\left[\begin{matrix}0&0&0\\ 0&-\sin r_{\mathrm{x}}^{\text{A}}&-\cos r_{\mathrm{x}}^{\text{A}}\\ 0&\cos r_{\mathrm{x}}^{\text{A}}&-\sin r_{\mathrm{x}}^{\text{A}}\\ \end{matrix}\right],$ (77a)
+$\frac{\partial\boldsymbol{R}_{\mathrm{y}}\left(r_{\mathrm{y}}^{\text{A}}\right)}{\partial r_{\mathrm{y}}^{\text{A}}}=\left[\begin{matrix}-\sin r_{\mathrm{y}}^{\text{A}}&0&\cos r_{\mathrm{y}}^{\text{A}}\\ 0&0&0\\ -\cos r_{\mathrm{y}}^{\text{A}}&0&-\sin r_{\mathrm{y}}^{\text{A}}\\ \end{matrix}\right],$ (77b)
+$\frac{\partial\boldsymbol{R}_{\mathrm{z}}\left(r_{\mathrm{z}}^{\text{A}}\right)}{\partial r_{\mathrm{z}}^{\text{A}}}=\left[\begin{matrix}-\sin r_{\mathrm{z}}^{\text{A}}&-\cos r_{\mathrm{z}}^{\text{A}}&0\\ \cos r_{\mathrm{z}}^{\text{A}}&-\sin r_{\mathrm{z}}^{\text{A}}&0\\ 0&0&0\\ \end{matrix}\right].$ (77c)
 For the directional gains and their derivatives, let $\bar{\boldsymbol{n}}=[0,0,1]^{T}$ and $\boldsymbol{n}(\boldsymbol{r}^{\text{A}})=\boldsymbol{R}(\boldsymbol{r}^{\text{A}})\bar{\boldsymbol{n}}$ be the array boresight.
 According to the practical directivity model (11), we have
-$\tfrac{\partial\sqrt{G}}{\partial\boldsymbol{r}}=\frac{1}{2\sqrt{G}}\frac{\partial G}{\partial\boldsymbol{r}}.$ (77a)
-$\frac{\partial G^{\text{A}}}{\partial\boldsymbol{r}^{\text{A}}}=G_{0}^{\text{A}}\,p^{\text{A}}\big(-\boldsymbol{n}^{T}\boldsymbol{u}\big)^{p^{\text{A}}-1}\Big(-\frac{\partial\boldsymbol{n}^{T}}{\partial\boldsymbol{r}^{\text{A}}}\boldsymbol{u}\Big),$ (77b)
+$\frac{\partial\sqrt{G^{\text{A}}}}{\partial\boldsymbol{r}^{\text{A}}}=\frac{1}{2\sqrt{G^{\text{A}}}}\frac{\partial G^{\text{A}}}{\partial\boldsymbol{r}^{\text{A}}},$ (78a)
+[…]
 $\frac{\partial\boldsymbol{n}^{T}}{\partial\boldsymbol{r}^{\text{A}}}=\bar{\boldsymbol{n}}^{T}\frac{\partial\boldsymbol{R}(\boldsymbol{r}^{\text{A}})}{\partial\boldsymbol{r}^{\text{A}}}.$ (78c)
 At the non-differentiable boundary, subgradients are used.
-$\,\,\frac{\partial\boldsymbol{h}_{k}(\boldsymbol{r}^{\text{B}})}{\partial\boldsymbol{r}^{\text{B}}}=\sum_{l=1}^{L_{k}}{a_{k,l}\bigg[\frac{\partial\sqrt{G_{k,l}^{\text{B}}(\boldsymbol{r}^{\text{B}})}}{\partial\boldsymbol{r}^{\text{B}}}\boldsymbol{t}_{k,l}^{\text{BU}}(\boldsymbol{r}^{\text{B}})+\sqrt{G_{k,l}^{\text{B}}(\boldsymbol{r}^{\text{B}})}\frac{\partial\boldsymbol{t}_{k,l}^{\text{BU}}(\boldsymbol{r}^{\text{B}})}{\partial\boldsymbol{r}^{\text{B}}}\bigg]}$ (78a)
-$\,\,\frac{\partial\boldsymbol{g}_{k}(\boldsymbol{r}^{\text{R}})}{\partial\boldsymbol{r}^{\text{R}}}=\sum_{q=1}^{Q_{k}}{c_{k,q}\bigg[\frac{\partial\sqrt{G_{k,q}^{\text{R}}(\boldsymbol{r}^{\text{R}})}}{\boldsymbol{r}^{\text{R}}}\boldsymbol{t}_{k,l}^{\text{RU}}(\boldsymbol{r}^{\text{R}})+\sqrt{G_{k,q}^{\text{R}}(\boldsymbol{r}^{\text{R}})}\frac{\boldsymbol{t}_{k,l}^{\text{RU}}(\boldsymbol{r}^{\text{R}})}{\partial\boldsymbol{r}^{\text{R}}}\bigg]}.$ (78b)
-$\,\,\frac{\partial\boldsymbol{B}(\boldsymbol{r})}{\partial\boldsymbol{r}^{\text{B}}}=\sum_{p=1}^{P}{b_{p}}\bigg[\frac{\partial\sqrt{G_{k,p}^{\text{B}}(\boldsymbol{r}^{\text{B}})}}{\partial\boldsymbol{r}^{\text{B}}}\sqrt{G_{k,p}^{\text{R}}(\boldsymbol{r}^{\text{R}})}\boldsymbol{t}_{p}^{\text{BR}}(\boldsymbol{r}^{\text{B}})(\boldsymbol{t}_{p}^{\text{RB}}(\boldsymbol{r}^{\text{R}}))^{H}+\sqrt{G_{k,p}^{\text{B}}(\boldsymbol{r}^{\text{B}})G_{k,p}^{\text{R}}(\boldsymbol{r}^{\text{R}})}\frac{\partial\boldsymbol{t}_{p}^{\text{BR}}(\boldsymbol{r}^{\text{B}})}{\partial\boldsymbol{r}^{\text{B}}}(\boldsymbol{t}_{p}^{\text{RB}}(\boldsymbol{r}^{\text{R}}))^{H}\bigg],$ (78c)
-$\,\,\frac{\partial\boldsymbol{B}(\boldsymbol{r})}{\partial\boldsymbol{r}^{\text{R}}}=\sum_{p=1}^{P}{b_{p}}\bigg[\sqrt{G_{k,p}^{\text{B}}(\boldsymbol{r}^{\text{B}})}\frac{\partial\sqrt{G_{k,p}^{\text{R}}(\boldsymbol{r}^{\text{R}})}}{\partial\boldsymbol{r}^{\text{R}}}\boldsymbol{t}_{p}^{\text{BR}}(\boldsymbol{r}^{\text{B}}(\boldsymbol{r}^{\text{B}}))(\boldsymbol{t}_{p}^{\text{RB}}(\boldsymbol{r}^{\text{R}}))^{H}+\sqrt{G_{k,p}^{\text{B}}(\boldsymbol{r}^{\text{B}})G_{k,p}^{\text{R}}(\boldsymbol{r}^{\text{R}})}\boldsymbol{t}_{p}^{\text{BR}}(\boldsymbol{r}^{\text{B}})\Big(\frac{\partial\boldsymbol{t}_{p}^{\text{RB}}(\boldsymbol{r}^{\text{R}})}{\partial\boldsymbol{r}^{\text{R}}}\Big)^{H}\bigg],$ (78d)
+$\,\,\frac{\partial\boldsymbol{h}_{k}(\boldsymbol{r}^{\text{B}})}{\partial\boldsymbol{r}^{\text{B}}}=\sum_{l=1}^{L_{k}}{\delta_{k,l}\bigg[\frac{\partial\sqrt{G_{k,l}^{\text{B}}(\boldsymbol{r}^{\text{B}})}}{\partial\boldsymbol{r}^{\text{B}}}\boldsymbol{t}_{k,l}^{\text{BU}}(\boldsymbol{r}^{\text{B}})+\sqrt{G_{k,l}^{\text{B}}(\boldsymbol{r}^{\text{B}})}\frac{\partial\boldsymbol{t}_{k,l}^{\text{BU}}(\boldsymbol{r}^{\text{B}})}{\partial\boldsymbol{r}^{\text{B}}}\bigg]},$ (79a)
+$\,\,\frac{\partial\boldsymbol{g}_{k}(\boldsymbol{r}^{\text{R}})}{\partial\boldsymbol{r}^{\text{R}}}=\sum_{q=1}^{Q_{k}}{\psi_{k,q}\bigg[\frac{\partial\sqrt{G_{k,q}^{\text{R}}(\boldsymbol{r}^{\text{R}})}}{\partial\boldsymbol{r}^{\text{R}}}\boldsymbol{t}_{k,q}^{\text{RU}}(\boldsymbol{r}^{\text{R}})+\sqrt{G_{k,q}^{\text{R}}(\boldsymbol{r}^{\text{R}})}\frac{\boldsymbol{t}_{k,q}^{\text{RU}}(\boldsymbol{r}^{\text{R}})}{\partial\boldsymbol{r}^{\text{R}}}\bigg]},$ (79b)
+$\,\,\frac{\partial\boldsymbol{B}(\boldsymbol{r})}{\partial\boldsymbol{r}^{\text{B}}}=\sum_{p=1}^{P}{\gamma_{p}}\bigg[\frac{\partial\sqrt{G_{p}^{\text{B}}(\boldsymbol{r}^{\text{B}})}}{\partial\boldsymbol{r}^{\text{B}}}\sqrt{G_{p}^{\text{R}}(\boldsymbol{r}^{\text{R}})}\boldsymbol{t}_{p}^{\text{BR}}(\boldsymbol{r}^{\text{B}})(\boldsymbol{t}_{p}^{\text{RB}}(\boldsymbol{r}^{\text{R}}))^{H}+\sqrt{G_{p}^{\text{B}}(\boldsymbol{r}^{\text{B}})G_{p}^{\text{R}}(\boldsymbol{r}^{\text{R}})}\frac{\partial\boldsymbol{t}_{p}^{\text{BR}}(\boldsymbol{r}^{\text{B}})}{\partial\boldsymbol{r}^{\text{B}}}(\boldsymbol{t}_{p}^{\text{RB}}(\boldsymbol{r}^{\text{R}}))^{H}\bigg],$ (79c)
+$\,\,\frac{\partial\boldsymbol{B}(\boldsymbol{r})}{\partial\boldsymbol{r}^{\text{R}}}=\sum_{p=1}^{P}{\gamma_{p}}\bigg[\sqrt{G_{p}^{\text{B}}(\boldsymbol{r}^{\text{B}})}\frac{\partial\sqrt{G_{p}^{\text{R}}(\boldsymbol{r}^{\text{R}})}}{\partial\boldsymbol{r}^{\text{R}}}\boldsymbol{t}_{p}^{\text{BR}}(\boldsymbol{r}^{\text{B}}(\boldsymbol{r}^{\text{B}}))(\boldsymbol{t}_{p}^{\text{RB}}(\boldsymbol{r}^{\text{R}}))^{H}+\sqrt{G_{p}^{\text{B}}(\boldsymbol{r}^{\text{B}})G_{p}^{\text{R}}(\boldsymbol{r}^{\text{R}})}\boldsymbol{t}_{p}^{\text{BR}}(\boldsymbol{r}^{\text{B}})\Big(\frac{\partial\boldsymbol{t}_{p}^{\text{RB}}(\boldsymbol{r}^{\text{R}})}{\partial\boldsymbol{r}^{\text{R}}}\Big)^{H}\bigg],$ (79d)
 $\overline${\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ }
-According to the sensing beam pattern (18) and the sensing channel model (14), for the beam-pattern derivative,
-$\frac{\partial\mathcal{P}_{a}}{\partial\boldsymbol{r}^{\text{A}}}=2\Re\Big\{\Big(\frac{\partial\boldsymbol{f}_{\text{S},a}(\boldsymbol{r})}{\partial\boldsymbol{r}^{\mathrm{A}}}\Big)^{H}\boldsymbol{W}\boldsymbol{W}^{H}\boldsymbol{f}_{\text{S},a}(\boldsymbol{r})\Big\},$ (79)
-where $\frac{\partial\boldsymbol{f}_{\text{S},a}(\boldsymbol{r})}{\partial\boldsymbol{r}^{\mathrm{A}}}$ is computed in (80).
-$\frac{\partial\boldsymbol{f}_{\text{S},a}\left(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}}\right)}{\partial\boldsymbol{r}^{\text{B}}}=a_{a}\frac{\partial\left(\sqrt{G_{a}^{\text{B}}\left(\boldsymbol{r}^{\text{B}}\right)}\boldsymbol{t}^{\text{BT}}\left(\boldsymbol{r}^{\text{B}}\right)\right)}{\partial\boldsymbol{r}^{\text{B}}}+c_{a}\frac{\partial\boldsymbol{B}\left(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}}\right)}{\partial\boldsymbol{r}^{\text{B}}}\mathrm{diag}\left(\boldsymbol{\theta}\right)\sqrt{G_{a}^{\text{R}}\left(\boldsymbol{r}^{\text{R}}\right)}\,\boldsymbol{t}^{\text{RT}}\left(\boldsymbol{r}^{\text{R}}\right),$ (80a)
-$\frac{\partial\boldsymbol{f}_{\text{S},a}\left(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}}\right)}{\partial\boldsymbol{r}^{\text{R}}}=a_{a}\frac{\partial\boldsymbol{B}\left(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}}\right)}{\partial\boldsymbol{r}^{\text{R}}}\mathrm{diag}\left(\boldsymbol{\theta}\right)\sqrt{G_{a}^{\text{R}}}\boldsymbol{t}^{\text{RT}}+c_{a}\boldsymbol{B}\left(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}}\right)\mathrm{diag}\left(\boldsymbol{\theta}\right)\frac{\partial\left(\sqrt{G_{a}^{\text{R}}\left(\boldsymbol{r}^{\text{R}}\right)}\,\boldsymbol{t}^{\text{RT}}\left(\boldsymbol{r}^{\text{R}}\right)\right)}{\partial\boldsymbol{r}^{\text{R}}}$ (80b)
+According to the sensing beampattern (18) and the sensing channel model (II-C), for the beampattern derivative,
+$\frac{\partial\mathcal{P}_{a}}{\partial\boldsymbol{r}^{\text{A}}}=2\Re\Big\{\Big(\frac{\partial\boldsymbol{f}_{\text{S},a}(\boldsymbol{r})}{\partial\boldsymbol{r}^{\text{A}}}\Big)^{H}\boldsymbol{W}\boldsymbol{W}^{H}\boldsymbol{f}_{\text{S},a}(\boldsymbol{r})\Big\},$ (80)
+where $\frac{\partial\boldsymbol{f}_{\text{S},a}(\boldsymbol{r})}{\partial\boldsymbol{r}^{\text{A}}}$ is computed in (81).
+$\frac{\partial\boldsymbol{f}_{\text{S},a}\left(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}}\right)}{\partial\boldsymbol{r}^{\text{B}}}=\frac{\partial\left(\sqrt{G_{a}^{\text{B}}\left(\boldsymbol{r}^{\text{B}}\right)}\boldsymbol{t}^{\text{BT}}\left(\boldsymbol{r}^{\text{B}}\right)\right)}{\partial\boldsymbol{r}^{\text{B}}}+\frac{\partial\boldsymbol{B}\left(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}}\right)}{\partial\boldsymbol{r}^{\text{B}}}\mathrm{diag}\left(\boldsymbol{\theta}\right)\sqrt{G_{a}^{\text{R}}\left(\boldsymbol{r}^{\text{R}}\right)}\,\boldsymbol{t}^{\text{RT}}\left(\boldsymbol{r}^{\text{R}}\right),$ (81a)
+$\frac{\partial\boldsymbol{f}_{\text{S},a}\left(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}}\right)}{\partial\boldsymbol{r}^{\text{R}}}=\frac{\partial\boldsymbol{B}\left(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}}\right)}{\partial\boldsymbol{r}^{\text{R}}}\mathrm{diag}\left(\boldsymbol{\theta}\right)\sqrt{G_{a}^{\text{R}}}\boldsymbol{t}^{\text{RT}}+\boldsymbol{B}\left(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}}\right)\mathrm{diag}\left(\boldsymbol{\theta}\right)\frac{\partial\left(\sqrt{G_{a}^{\text{R}}\left(\boldsymbol{r}^{\text{R}}\right)}\,\boldsymbol{t}^{\text{RT}}\left(\boldsymbol{r}^{\text{R}}\right)\right)}{\partial\boldsymbol{r}^{\text{R}}}$ (81b)
 $\overline${\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ }
-It is also convenient to gather partials into Jacobians $\boldsymbol{J}_{\text{S},a}^{(\text{A})}\triangleq\left[\frac{\partial\boldsymbol{f}_{\text{S},a}}{\partial r_{x}^{\text{A}}},\frac{\partial\boldsymbol{f}_{\text{S},a}}{\partial r_{y}^{\text{A}}},\frac{\partial\boldsymbol{f}_{\text{S},a}}{\partial r_{z}^{\text{A}}}\right]\in\mathbb{C}^{N_{t}\times 3}$ similar to (74), so that (79) compacts to
-$\frac{\partial\mathcal{P}_{a}}{\partial\boldsymbol{r}^{\mathrm{A}}}=2\,\Re\!\Big\{\boldsymbol{J}_{\text{S},a}^{(\text{A})\,H}\boldsymbol{W}\boldsymbol{W}^{H}\boldsymbol{f}_{\text{S},a}\Big\}\in\mathbb{R}^{3}.$ (81)
+It is also convenient to gather partials into Jacobians $\boldsymbol{J}_{\text{S},a}^{(\text{A})}\triangleq\left[\frac{\partial\boldsymbol{f}_{\text{S},a}}{\partial r_{\mathrm{x}}^{\text{A}}},\frac{\partial\boldsymbol{f}_{\text{S},a}}{\partial r_{\mathrm{y}}^{\text{A}}},\frac{\partial\boldsymbol{f}_{\text{S},a}}{\partial r_{\mathrm{z}}^{\text{A}}}\right]\in\mathbb{C}^{M\times 3}$ similar to (75), so that (80) compacts to
+$\frac{\partial\mathcal{P}_{a}}{\partial\boldsymbol{r}^{\text{A}}}=2\,\Re\!\Big\{\boldsymbol{J}_{\text{S},a}^{(\text{A})\,H}\boldsymbol{W}\boldsymbol{W}^{H}\boldsymbol{f}_{\text{S},a}\Big\}\in\mathbb{R}^{3}.$ (82)
 III-D2 Projected-gradient framework for the rotation block.
-Let $\mathcal{D}\triangleq\mathcal{D}_{\text{B}}\times\mathcal{D}_{\text{R}}$ be the box of angle limits in (P4) with
-$\mathcal{D}_{\text{A}}=\{\boldsymbol{r}^{\text{A}}:\boldsymbol{r}_{\text{lb}}^{\text{A}}\leq\boldsymbol{r}^{\text{A}}\leq\boldsymbol{r}_{\text{ub}}^{\text{A}}\},$ (82)
-and denote $\boldsymbol{r}_{\text{lb}}\!\triangleq\![(\boldsymbol{r}_{\text{lb}}^{\text{B}})^{T}\!,(\boldsymbol{r}_{\text{lb}}^{\text{R}})^{T}]^{T}$, and $\boldsymbol{r}_{\text{ub}}\!\triangleq\![(\boldsymbol{r}_{\text{ub}}^{\text{B}})^{T}\!,(\boldsymbol{r}_{\text{ub}}^{\text{R}})^{T}]^{T}$.
-Let $\Pi_{\mathcal{D}}(\cdot)$ be the Euclidean projection onto $\mathcal{D}$.
-We perform projected gradient ascent on $F$ with the projected step
-$\boldsymbol{\beta}\left(\alpha\right)=\Pi_{\mathcal{D}}\left(\boldsymbol{r}+\alpha\nabla_{\boldsymbol{r}}F\left(\boldsymbol{r}\right)\right)-\boldsymbol{r},$ (83)
+Let $\mathcal{B}$ be the box of array rotation angle limits in (P4)
+$\mathcal{B}=\{\boldsymbol{r}:\boldsymbol{r}_{\text{lb}}\preceq\boldsymbol{r}\preceq\boldsymbol{r}_{\text{ub}}\},$ (83)
+where we denote $\boldsymbol{r}_{\text{lb}}\triangleq[(\boldsymbol{r}_{\text{lb}}^{\text{B}})^{T},(\boldsymbol{r}_{\text{lb}}^{\text{R}})^{T}]^{T}$ and $\boldsymbol{r}_{\text{ub}}\triangleq[(\boldsymbol{r}_{\text{ub}}^{\text{B}})^{T},(\boldsymbol{r}_{\text{ub}}^{\text{R}})^{T}]^{T}$, and here we let $\boldsymbol{r}_{\text{lb}}^{\text{A}}=[r_{\mathrm{x},\min}^{\text{A}},r_{\mathrm{y},\min}^{\text{A}},r_{\mathrm{z},\min}^{\text{A}}]$ and $\boldsymbol{r}_{\text{ub}}^{\text{A}}=[r_{\mathrm{x},\max}^{\text{A}},r_{\mathrm{y},\max}^{\text{A}},r_{\mathrm{z},\max}^{\text{A}}]$ corresponding to (24d) and (24e), respectively, for brevity.
+Let $\Pi_{\mathcal{B}}(\cdot)$ be the Euclidean projection onto $\mathcal{B}$.
+We perform projected gradient ascent of $F\left(\boldsymbol{r}\right)$ on the feasible region with the projected step
+$\boldsymbol{\beta}\left(\alpha\right)=\Pi_{\mathcal{B}}\left(\boldsymbol{r}+\alpha\nabla_{\boldsymbol{r}}F\left(\boldsymbol{r}\right)\right)-\boldsymbol{r},$ (84)
 for any step size $\alpha>0$.
-Define the projected-gradient mapping $\boldsymbol{H}_{\alpha}\left(\boldsymbol{r}\right)\triangleq\frac{1}{\alpha}\boldsymbol{\beta}\left(\alpha\right)$, for any $\alpha>0$, $\boldsymbol{r}^{\star}$ is a first-order stationary point of (P4) if and only if $\boldsymbol{H}_{\alpha}\left(\boldsymbol{r}^{\star}\right)=\boldsymbol{0}$, according to the KKT condition.
-Hence $\|\boldsymbol{H}_{\alpha}(\boldsymbol{r})\|_{2}$ is a valid criticality measure for stopping.
-Because $\mathcal{D}$ is a box, the projection (83) reduces to the componentwise clipping
-$\boldsymbol{\beta}^{(\ell)}(\alpha)=\text{clip}(\boldsymbol{r}^{\left(\ell\right)}+\alpha\nabla_{\boldsymbol{r}}F^{(\ell)}(\boldsymbol{r}))_{[r_{\text{lb}},r_{\text{ub}}]}-\boldsymbol{r}^{(\ell)},$ (84)
-where the clip $\text{clip}\left(x\right)_{\left[y,z\right]}=\min\left\{\max\left\{x,y\right\},z\right\}$ acts elementwise.
-Given the previous $\left(\boldsymbol{r},F\left(\boldsymbol{r}\right),\nabla F\left(\boldsymbol{r}\right)\right)^{(\ell-1)}$, the stepsize uses Barzilai–Borwein initialization
+Define the projected-gradient mapping $\boldsymbol{\upsilon}_{\alpha}\left(\boldsymbol{r}\right)\triangleq\frac{1}{\alpha}\boldsymbol{\beta}\left(\alpha\right)$, for any $\alpha>0$, $\boldsymbol{r}^{\star}$ is a first-order stationary point of (P4) if and only if $\boldsymbol{\upsilon}_{\alpha}\left(\boldsymbol{r}^{\star}\right)=\boldsymbol{0}$, according to the KKT condition.
+Hence $\|\boldsymbol{\upsilon}_{\alpha}(\boldsymbol{r})\|_{2}$ is a valid criticality measure for stopping.
+Because $\mathcal{B}$ is a box, the projection (84) reduces to the component-wise clipping
+$\boldsymbol{\beta}^{(\ell)}(\alpha)=\mathrm{clip}(\boldsymbol{r}^{\left(\ell\right)}+\alpha\nabla_{\boldsymbol{r}}F^{(\ell)}(\boldsymbol{r}))_{[r_{\text{lb}},r_{\text{ub}}]}-\boldsymbol{r}^{(\ell)},$ (85)
+where the clip operation is defined by $\mathrm{clip}\left(\mathfrak{a}\right)_{\left[\mathfrak{b},\mathfrak{c}\right]}=\min\left\{\max\left\{\mathfrak{a},\mathfrak{b}\right\},\mathfrak{c}\right\}$ that acts element-wise.
+Given the previous $\left(\boldsymbol{r},F\left(\boldsymbol{r}\right),\nabla F\left(\boldsymbol{r}\right)\right)^{(\ell-1)}$, the step size uses the Barzilai–Borwein initialization
 $\alpha_{\text{BB}}=\frac{\|\boldsymbol{r}^{\left(\ell\right)}-\boldsymbol{r}^{\left(\ell-1\right)}\|^{2}}{\left(\boldsymbol{r}^{\left(\ell\right)}-\boldsymbol{r}^{\left(\ell-1\right)}\right)^{T}\left(\nabla F\left(\boldsymbol{r}^{\left(\ell\right)}\right)-\nabla F\left(\boldsymbol{r}^{\left(\ell-1\right)}\right)\right)},$ (86)
 and clip to $\alpha_{\min},\alpha_{\max}]$ to ensure robustness. An Armijo backtracking is then performed until the ascent rule holds
-$F\left(\boldsymbol{r}+\boldsymbol{\beta}\left(\alpha\right)\right)\geq F\left(\boldsymbol{r}\right)+c_{1}\nabla_{\boldsymbol{r}}F\left(\boldsymbol{r}\right)^{T}\boldsymbol{\beta}\left(\alpha\right),$ (86)
-with $c_{1}\!\in\!(0,1)$ and the backtracking using $\alpha\!\leftarrow\!c_{2}\alpha$, $c_{2}\in(0,1)$.
-The iteration terminates when it meets the stopping criteria that the projected gradient norm is small, i.e., $\|\boldsymbol{H}_{\alpha}(\boldsymbol{r}^{(\ell)})\|_{2}\ \leq\ \varepsilon_{\text{pg}}$, or the relative step is small, i.e., $\frac{\|\boldsymbol{r}^{(\ell)}-\boldsymbol{r}^{(\ell-1)}\|_{2}}{\max\{1,\|\boldsymbol{r}^{(\ell-1)}\|_{2}\}}\ \leq\ \varepsilon_{\text{rel}}$.
-For box constraints, $\boldsymbol{H}_{\alpha}(\boldsymbol{r}^{(\ell)})$ admits the following componentwise form for $\imath$-th entry independent of $\alpha$
-$\!\!\!\big[\boldsymbol{H}_{\alpha}(\boldsymbol{r}^{(\ell)})\big]_{\imath}\!=\!\!\begin{cases}[\nabla_{\boldsymbol{r}}F^{\left(\ell\right)}(\boldsymbol{r})]_{\imath},\!&\!\!\!r_{\text{lb},i}\!\!<r_{i}\!<r_{\text{ub},i}\!,\\ \max\{0,[\nabla_{\boldsymbol{r}}F^{(\ell)}(\boldsymbol{r})]_{\imath}\},\!&\!\!\!r_{\imath}=r_{\text{lb},\imath},\\ \min\{0,[\nabla_{\boldsymbol{r}}F^{(\ell)}(\boldsymbol{r})]_{\imath}\},\!&\!\!\!r_{\imath}=r_{\text{ub},\imath}.\end{cases}\!\!\!\!$ (87)
-The gradient cost is dominated by Jacobians (80) of rotation-related steering vectors and derivatives (78) with matrix-chain products.
-Because the directional-gain model is piecewise smooth, we use subgradients at the visibility boundary and a smooth gating for stable line search.
+$F\left(\boldsymbol{r}+\boldsymbol{\beta}\left(\alpha\right)\right)\geq F\left(\boldsymbol{r}\right)+\kappa_{1}\nabla_{\boldsymbol{r}}F\left(\boldsymbol{r}\right)^{T}\boldsymbol{\beta}\left(\alpha\right),$ (87)
+with $\kappa_{1}\!\in\!(0,1)$ and the backtracking using $\alpha\!\leftarrow\!\kappa_{2}\alpha$, $\kappa_{2}\in(0,1)$.
+The iteration terminates when it meets the stopping criteria that the projected gradient norm is smaller than a threshold $\varepsilon_{\text{pg}}$, i.e., $\|\boldsymbol{\upsilon}_{\alpha}(\boldsymbol{r}^{(\ell)})\|_{2}\ \leq\ \varepsilon_{\text{pg}}$, or the relative step is smaller than a threshold $\varepsilon_{\text{rel}}$, i.e., $\frac{\|\boldsymbol{r}^{(\ell)}-\boldsymbol{r}^{(\ell-1)}\|_{2}}{\max\{1,\|\boldsymbol{r}^{(\ell-1)}\|_{2}\}}\ \leq\ \varepsilon_{\text{rel}}$.
+For box constraints, $\boldsymbol{\upsilon}_{\alpha}(\boldsymbol{r}^{(\ell)})$ admits the following component-wise form for $\imath$-th entry, according to (69), independent of $\alpha$
+$\!\!\!\big[\boldsymbol{\upsilon}_{\alpha}(\boldsymbol{r}^{(\ell)})\big]_{\imath}\!=\!\!\begin{cases}[\nabla_{\boldsymbol{r}}F^{\left(\ell\right)}(\boldsymbol{r})]_{\imath},\!&\!\!\!r_{\text{lb},i}\!\!<\!r_{i}\!<\!r_{\text{ub},i},\\ \max\{0,[\nabla_{\boldsymbol{r}}F^{(\ell)}(\boldsymbol{r})]_{\imath}\},\!&\!\!\!r_{\imath}=r_{\text{lb},\imath},\\ \min\{0,[\nabla_{\boldsymbol{r}}F^{(\ell)}(\boldsymbol{r})]_{\imath}\},\!&\!\!\!r_{\imath}=r_{\text{ub},\imath}.\end{cases}\!\!\!\!$ (88)
+The gradient cost is dominated by Jacobians of rotation-related steering vectors and derivatives with matrix-chain products.
+Because the directional-gain model is piecewise-smooth, we use subgradients in (88) along the visibility boundary to ensure stable line search.
 Algorithm 3 PGD for Array Rotation Subproblem (P4)
-1:  Input: fixed $(\boldsymbol{W},\boldsymbol{\theta})$, box $\mathcal{D}$, tolerances $\varepsilon_{\mathrm{pg}},\varepsilon_{\mathrm{rel}}>0$, and maximum iterations $I_{\text{PGD}}^{\text{max}}$.
-2:  Initialization: initial point $\boldsymbol{r}^{(0)}\in\mathcal{D}$
+1:  Input: fixed $(\boldsymbol{W},\boldsymbol{\theta})$, box $\mathcal{B}$, tolerances $\varepsilon_{\mathrm{pg}},\varepsilon_{\mathrm{rel}}>0$, and maximum iterations $I_{\text{PGD}}^{\text{max}}$.
+2:  Initialization: initial point $\boldsymbol{r}^{(0)}\in\mathcal{B}$
 3:  for $\ell=0,1,\ldots,I_{\text{PGD}}^{\text{max}}-1$ do
@@
 12:   Update optimization variables by $\boldsymbol{r}^{(\ell+1)}=\boldsymbol{r}^{(\ell)}+\boldsymbol{\beta}^{(\ell)}(\alpha)$.
-13:   Compute projected gradient mapping $\boldsymbol{H}_{\alpha}(\boldsymbol{r}^{(\ell)})$ via (87).
-14:   if $\|\boldsymbol{H}_{\alpha}(\boldsymbol{r}^{(\ell+1)})\|_{2}\leq\varepsilon_{\mathrm{pg}}$ or $\dfrac{\|\boldsymbol{r}^{(\ell+1)}-\boldsymbol{r}^{(\ell)}\|_{2}}{\max\{1,\|\boldsymbol{r}^{(\ell)}\|_{2}\}}\leq\varepsilon_{\mathrm{rel}}$ then
+13:   Compute projected gradient mapping $\boldsymbol{\upsilon}_{\alpha}(\boldsymbol{r}^{(\ell)})$ via (88).
+14:   if $\|\boldsymbol{\upsilon}_{\alpha}(\boldsymbol{r}^{(\ell+1)})\|_{2}\leq\varepsilon_{\mathrm{pg}}$ or $\dfrac{\|\boldsymbol{r}^{(\ell+1)}-\boldsymbol{r}^{(\ell)}\|_{2}}{\max\{1,\|\boldsymbol{r}^{(\ell)}\|_{2}\}}\leq\varepsilon_{\mathrm{rel}}$ then
 15:    break;
@@
 17:  end for
-18:  Output: $\boldsymbol{r}^{\mathrm{B}}\leftarrow\boldsymbol{r}^{(\ell+1)}_{1:3}$, $\boldsymbol{r}^{\mathrm{R}}\leftarrow\boldsymbol{r}^{(\ell+1)}_{4:6}$.
-III-D3 Overall algorithm under penalty-based AO framework for original problem 1
+18:  Output: $\boldsymbol{r}^{\text{B}}\leftarrow\boldsymbol{r}^{(\ell+1)}_{1:3}$, $\boldsymbol{r}^{\text{R}}\leftarrow\boldsymbol{r}^{(\ell+1)}_{4:6}$.
+III-D3 Overall algorithm and complexity analysis for subproblem 3
 Algorithm 3 summarizes the overall projected gradient method.
 For the per-iteration computational complexity, directly forming $\boldsymbol{B}$ and the Jacobians as dense $M\times N$ matrices yields $\mathcal{O}(PMN)$ computations per rotation axis, and evaluating the sensing gradients over $A$ grids would add $\mathcal{O}(AMN)$.
-However, using the rank-$P$ factorization for $\boldsymbol{B}$ and its derivatives, all costly matrix–vector products, involving $\boldsymbol{B}$ and $\partial\boldsymbol{B}/\partial r_{\imath}^{\mathrm{A}}$, are computed without materializing $M\times N$ blocks and costing $\mathcal{O}(NP)$ plus $\mathcal{O}(MP)$.
+However, using the rank-$P$ factorization for $\boldsymbol{B}$ and its derivatives, all costly matrix–vector products, involving $\boldsymbol{B}$ and $\partial\boldsymbol{B}/\partial r_{\imath}^{\text{A}}$, are computed without materializing $M\times N$ blocks and costing $\mathcal{O}(NP)$ plus $\mathcal{O}(MP)$.
 With these reuses, one PGD iteration with full gradient of $F$ has the leading cost: i) precomputation on steering vectors and their Jacobians with $\mathcal{O}\big(P(M+N)\big)$, ii) sensing and communication channel–Jacobian products $\mathcal{O}\big((A+K)P(N+M)\big)$, iii) use $\boldsymbol{W}(\boldsymbol{W}^{H}\boldsymbol{f})$ to apply $\boldsymbol{W}\boldsymbol{W}^{H}$ in (72) and (82) with $\mathcal{O}\!\big((M+K)M(A+K)\big)$.
 Therefore, the per-iteration complexity scales as $\mathcal{O}\big(P(M+N)(A+K+1)+(M+K)M(A+K)\big)$ with a constant coefficient affected by backtracking searches.
-III-E Overall algorithm
-Finally, we adopt a penalty-assisted AO-based algorithm framework over $(\boldsymbol{W},\boldsymbol{\theta},\boldsymbol{r}^{\mathrm{B}},\boldsymbol{r}^{\mathrm{R}})$, resulting in an overall algorithm with nested loops summarized in Algorithm 4, where we obtain a suboptimal solution for the original problem (P1).
-At the $\delta$-th outer iteration, with the other blocks fixed, we first solve (P2) by Algorithm 1 to update $\boldsymbol{W}^{(\delta+1)}$; next solve (P3) by Algorithm 2 to update $\boldsymbol{\theta}^{(\delta+1)}$; then solve (P4) by Algorithm 3 (projected-gradient over the box $\mathcal{D}$) to obtain $(\boldsymbol{r}^{\mathrm{B}},\boldsymbol{r}^{\mathrm{R}})^{(\delta+1)}$.
-After the three inner loops, the penalty weight is refreshed as $\rho^{(\delta+1)}\leftarrow\gamma_{\rho}\rho^{(\delta)}$ if the MSE constraint is violated (i.e., $\mathrm{MSE}>\eta_{\text{th}}$); otherwise $\rho^{(\delta+1)}=\rho^{(\delta)}$.
-Each inner solver uses its own line search/tolerances to guarantee ascent of the penalized objective and is warm-started from the previous iterate.
-The outer loop terminates when the fractional increase of the penalized objective in (P1.1) falls below $\varepsilon_{\mathrm{AO}}$ and the MSE constraint is satisfied ($\mathrm{MSE}\leq\eta_{\text{th}}$).
-Algorithm 4 Penalty-assisted AO-based Algorithm for (P1)
-1:  Input: $\boldsymbol{u}_{k,l}^{\text{BU}},\boldsymbol{u}_{p}^{\text{RB}},\boldsymbol{u}_{p}^{\text{BR}},\boldsymbol{u}_{k,q}^{\text{RU}},\boldsymbol{u}_{a}^{\text{BT}},\boldsymbol{u}_{a}^{\text{RT}}$, $\gamma_{\rho}\!>\!0$, and $\mathcal{P}_{\text{d}}\left(\vartheta_{a},\varphi_{a}\right)$ .
-2:  Initialization: $\boldsymbol{W}^{\left(0\right)},\boldsymbol{\theta}^{\left(0\right)},\left(\boldsymbol{r}^{\text{A}},\boldsymbol{r}^{\text{B}}\right)^{\left(0\right)}$, $\rho^{(0)}$ and $\delta=0$
+III-E Overall algorithm for the original problem
+Finally, Algorithm 4 summarizes the overall AO procedure for (P1).
+At the $\zeta$-th outer iteration, we first update the NMSE scaling parameter by $\iota^{(\zeta)}\leftarrow\iota^{\star}$ in (23).
+Then, with the other blocks fixed, we update $\boldsymbol{W}$ by solving (P2), update $\boldsymbol{\theta}$ by solving (P3) on the complex circle manifold, and update $(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}})$ by solving (P4) via projected gradient ascent under box constraints.
+Since each block update is designed to be non-decreasing in the objective of (P1) with $\iota$ fixed, and the $\iota$-update minimizes $\text{NMSE}(\iota)$ for fixed $(\boldsymbol{W},\boldsymbol{\theta},\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}})$, the overall objective in (24a) is monotonically non-decreasing over outer iterations.
+The algorithm thus converges to a stationary point of (P1) under standard regularity assumptions.
+However, due to the intrinsic non-convexity of (P1), global optimality is generally intractable, and the proposed AO-based algorithm obtains a suboptimal solution for feasible system design.
+Algorithm 4 AO-based Algorithm for (P1)
+1:Input: channel and angular parameters, $P_{\text{B}}$, $\rho$, desired pattern $\{\mathcal{P}_{\text{d},a}\}_{a=1}^{A}$, tolerances $\epsilon_{\mathrm{AO}}$.
+2:Initialize: feasible $\boldsymbol{W}^{(0)}$, $\boldsymbol{\theta}^{(0)}$, and $(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}})^{(0)}$ and set $\zeta=0$.
 3:repeat
-4:   Solve subproblem (P2) via Algorithm 1 to obtain $\boldsymbol{W}^{\left(\delta+1\right)}$.
-5:   Solve subproblem (P3) via Algorithm 2 to obtain $\boldsymbol{\theta}^{\left(\delta+1\right)}$.
-6:   Solve subproblem (P4) via Algorithm 3 to obtain $(\boldsymbol{r}^{\text{A}}\!,\!\boldsymbol{r}^{\text{B}})^{(\delta\!+\!1)}$.
-7:   if Current MSE value violates the threshold $\eta_{\text{th}}$ then
-8:    $\rho^{(\delta+1)}=\gamma\rho^{(\delta)}$.
-9:   else
-10:    $\rho^{(\delta+1)}=\rho^{(\delta)}$.
-11:   end if
-12:   $\delta\leftarrow\delta+1$.
-13:  until the fractional increase of the objective value of problem (P1.1) is smaller than a threshold $\varepsilon_{\text{AO}}$ and the MSE is below the threshold $\eta_{\text{th}}$.
-14:  Output: $\boldsymbol{W}\leftarrow\boldsymbol{W}^{(\delta)},\boldsymbol{\theta}\leftarrow\boldsymbol{\theta}^{(\delta)},\boldsymbol{r}^{\text{A}}\leftarrow\boldsymbol{r}^{\text{A},(\delta)}$, $\boldsymbol{r}^{\text{B}}\leftarrow\boldsymbol{r}^{\text{B},(\delta)}$.
+4:  Compute current beampattern samples $\boldsymbol{p}^{(\zeta)}=[\mathcal{P}_{1},\ldots,\mathcal{P}_{A}]^{T}$ and update $\iota^{(\zeta)}\leftarrow\iota^{\star}$ via (23).
+5:  Update $\boldsymbol{W}^{(\zeta+1)}$ by solving (P2) via QT, $M$, and bisection search.
+6:  Update $\boldsymbol{\theta}^{(\zeta+1)}$ by solving (P3) via RCG on the complex circle manifold.
+7:  Update $(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}})^{(\zeta+1)}$ by solving (P4) via projected gradient ascent with Armijo backtracking.
+8:  $\zeta\leftarrow\zeta+1$.
+9:until the relative increase of the objective in (P1) is below $\epsilon_{\mathrm{AO}}$ (or a maximum iteration is reached).
+10:Output: $(\boldsymbol{W},\boldsymbol{\theta},\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}})$.
 IV Simulation Results
-This section presents numerical results for the proposed rotatable BS–RIS ISAC design.
-Unless otherwise stated, the BS is equipped with a $M=M_{x}M_{y}=2\times 2$ antennas by default with inter-element spacing $d_{x}=d_{y}=\lambda/2$ for a carrier wavelength $\lambda=0.1$ m.
-The RIS is modeled as a $6\times 6$ UPA with spacing $e_{x}=e_{y}=\lambda/2$, located at $(10,0,0)$ m relative to the BS.
-The unrotated BS and RIS boresights are aligned with the positive $z$-axis, and their orientations are parameterized by rotation vectors $\boldsymbol{r}^{\mathrm{B}}$ and $\boldsymbol{r}^{\mathrm{R}}$.
-The small-scale fading of all links follows a geometry-based multipath model: each BS–user and RIS–user channel comprises $L_{k}=Q_{k}=2$ paths with independently generated azimuth and elevation angles, uniformly distributed over $\pi,\pi]$ and $\pi/3,\pi/3]$, respectively, and complex Gaussian path gains with normalized variance.
-The cascaded BS–RIS channel contains $P=2$ paths with angles and gains drawn independently.
-Thermal noise at each user is modeled as circularly symmetric complex Gaussian with variance $\sigma_{k}^{2}=10^{-6}$.
-For sensing, the BS forms a monostatic beampattern over a two-dimensional angular grid with $A_{\mathrm{az}}=11$ azimuth and $A_{\mathrm{el}}=6$ elevation samples, uniformly spanning $\phi\in[-\pi,\pi]$ and $\vartheta\in[-\pi/4,\pi/4]$.
-The desired sensing pattern corresponds to two spotlight-like angular sectors, and the target threshold of sensing MSE is $\eta_{\mathrm{th}}=0.2$.
-The total BS transmit power is $P_{\mathrm{B}}=30$ dBm.
-We adopt both half-space isotropic and directional element patterns.
-In the directional case, the BS and RIS element gains follow the piecewise model (11), while $G_{0,\mathrm{B}}=G_{0,\mathrm{R}}=2$ are constant in the isotopic case.
-For each random channel realization, the proposed alternating optimization algorithm jointly updates the BS precoder $\boldsymbol{W}$, the RIS phase vector $\boldsymbol{\theta}$, and the rotations $(\boldsymbol{r}^{\mathrm{B}},\boldsymbol{r}^{\mathrm{R}})$ so as to maximize the downlink sum rate $\sum_{k=1}^{K}\log_{2}(1+\mathrm{SINR}_{k})$ under the sensing MSE constraint.
-To assess the gain of spatial geometry control and RIS assistance, we compare the fully rotatable BS–RIS architecture (“Rot-BS+Rot-RIS”) with four benchmarks: (i) only the BS is rotatable while the RIS orientation is fixed (“Rot-BS+Fix-RIS”); (ii) only the RIS is rotatable (“Fix-BS+Rot-RIS”); (iii) both BS and RIS are fixed (“Fix-BS+Fix-RIS”); and two non-RIS baseline (iv) (“Rot-BS+No-RIS”) and (v) (“Fix-BS+No-RIS”), where the RIS–user link is disabled by setting the RIS–user path gains to zero.
-All reported curves are obtained by averaging the communication and sensing metrics over a large number of independent channel realizations.
-Figure 2: Sum rate versus maximum transmit power under $p^{\text{R}}=p^{\text{B}}=0$ with $N=6\times 6$, $M=2\times 2$ and $K=2$.
-Figure 3: Sum rate versus maximum transmit power under $p^{\text{R}}=p^{\text{B}}=2$ with $N=6\times 6$, $M=2\times 2$ and $K=2$.
+This section evaluates the proposed low-altitude ISAC design with cooperatively rotatable BS and RIS arrays.
+For performance characterization, we report three metrics: the downlink sum rate $R\triangleq\sum_{k=1}^{K}\log_{2}(1+\mathrm{SINR}_{k})$, the beampattern mismatch quantified by the NMSE in (22) with $\iota$ updated by (23) at each AO outer iteration, and the weighted ISAC utility $\mathcal{U}\triangleq\sum_{k=1}^{K}\log_{2}(1+\mathrm{SINR}_{k})-\rho\text{NMSE}$ of (P1).
+IV-A Simulation Setup and Benchmarks
+Unless otherwise stated, we set $\rho=10$.
+The BS located at (0,0,0)(0,0,0)employs a $M=M_{\text{col}}M_{\text{row}}=2\times 2$ UPA with inter-antenna spacing $\lambda/2$ at carrier wavelength $\lambda=0.1$ m.
+The RIS is a $6\times 6$ UPA with element spacing $\lambda/2$, located at $(10,0,0)$ m relative to the BS.
+The small-scale fading of all links follows a rotation-aware geometry-based multipath model: each BS-user and RIS-user channel contains $L_{k}=Q_{k}=2$ paths, and the BS-RIS channel contains $P=2$ paths.
+The azimuth and elevation angles are independently generated with $\varphi\in[-\pi,\pi]$ and $\vartheta\in[-\pi/3,\pi/3]$, and the complex path gains are drawn from independent circularly symmetric Gaussian distributions with normalized variance.
+For sensing, we discretize the angular region $\phi\in[-\pi,\pi]$ and $\vartheta\in[-\pi/4,\pi/4]$ with $A_{\mathrm{az}}=11$ azimuth sampling points and $A_{\mathrm{el}}=6$ elevation points, resulting in $A=A_{\mathrm{az}}A_{\mathrm{el}}=66$ grid points.
+The desired pattern is a spotlight-type template consisting of two angular sectors, and the NMSE is computed with the optimal scaling $\iota^{\star}$ in (23).
+The maximum rotation angle of the active and passive array is set by $\pi/2$ for each Euler angle in (24d) and (24e) by $r_{\text{i},\min}^{\text{R}}=r_{\text{i},\min}^{\text{B}}=-\frac{\pi}{2},r_{\text{i},\max}^{\text{R}}=r_{\text{i},\max}^{\text{B}}=\frac{\pi}{2},\text{i}\in\{\mathrm{x},\mathrm{y},\mathrm{z}\}$.
+We consider two settings on the directional radiation pattern gain of the BS antenna and the RIS element: half-space isotropic ($b^{\text{B}}=b^{\text{R}}=0$) and directional ($b^{\text{B}}=b^{\text{R}}=2$), where the latter yields a narrower mainlobe and stronger gain variations over directions.
+We compare the proposed jointly rotatable BS and RIS scheme (“Rot-BS & Rot-RIS”) with the following baselines: (i) “Rot-BS & Fix-RIS”, where only the BS array is rotatable while the RIS array orientation is fixed, (ii) “Fix-BS & Rot-RIS”, where only the RIS array is rotatable while the BS array orientation is fixed, (iii) “Fix-BS & Fix-RIS”, where both BS and RIS arrays are fixed, (iv) “Rot-BS & No-RIS”, where the RIS–user link is disabled by setting the RIS–user path gains to zero, and (v) “Fix-BS & No-RIS”.
+All reported curves are obtained by averaging the communication and sensing metrics over 100 independent channel realizations.
+Figure 2: Weighted ISAC utility versus maximum transmit power under $b^{\text{R}}=b^{\text{B}}=0$ with $\rho=10$, $N=6\times 6$, $M=2\times 2$ and $K=2$.
+Figure 3: Weighted ISAC utility versus maximum transmit power under $b^{\text{R}}=b^{\text{B}}=2$ with $\rho=10$, $N=6\times 6$, $M=2\times 2$ and $K=2$.
+Figure 4: Sum rate versus maximum transmit power under $b^{\text{R}}=b^{\text{B}}=0$ with $\rho=10$, $N=6\times 6$, $M=2\times 2$ and $K=2$.
+Figure 5: Sum rate versus maximum transmit power under $b^{\text{R}}=b^{\text{B}}=2$ with $\rho=10$, $N=6\times 6$, $M=2\times 2$ and $K=2$.
+Figure 6: Beampattern NMSE versus maximum transmit power under $b^{\text{R}}=b^{\text{B}}=0$ with $\rho=10$, $N=6\times 6$, $M=2\times 2$ and $K=2$.
+Figure 7: Beampattern NMSE versus maximum transmit power under $b^{\text{R}}=b^{\text{B}}=2$ with $\rho=10$, $N=6\times 6$, $M=2\times 2$ and $K=2$.
+IV-B Impact of Transmit Power
 Figs.
-2 and 3 illustrate the sum rate versus the BS transmit power under the sensing beampattern MSE constraint for a $2\times 2$ transmit array, two communication users, and two sensing directions.
-In the half-space isotropic case (Fig.
-2), all schemes benefit from increasing power, showing roughly linear growth at low–to–moderate SNR and gradual saturation when multiuser interference and the sensing constraint become dominant.
-Across the entire power range, the proposed joint BS–RIS rotation (“Rot-BS+Rot-RIS”) consistently achieves the highest sum rate, followed by the partially rotatable schemes.
-At the same time, the non-RIS baselines perform worse due to their reduced effective aperture.
-However, the RIS array rotations seem not to be effective in the half-space isotropic case, because rotations in this case cannot reap the directivity gains; one can say that isotropic elements avoid severe signal degradation in side-lobe directions.
-With a directional half-space pattern (Fig.
-3, element directivity factor $2$), all curves shift upward, confirming the SNR and throughput gains brought by element directivity.
-At the same time, the rate gap between Rot-BS+Rot-RIS and the benchmark schemes becomes more pronounced, as the proposed design better aligns the main lobes with user directions and steers reduced-gain regions toward the sensing directions, thereby alleviating the performance penalty of the sensing constraint.
-This indicates that geometric adaptation, especially when both active and passive arrays are rotatable, becomes increasingly valuable in the interference-limited regime and practical BS antenna or RIS element radiation patterns.
-Figure 4: Sum rate versus the number of BS antennas under $p^{\text{R}}=p^{\text{B}}=0$ with $N=6\times 6$, $M_{x}=1$ $P=30$ dBm, and $K=3$.
-Figure 5: Sum rate versus the number of BS antennas under $p^{\text{R}}=p^{\text{B}}=2$ with $N=6\times 6$, $M_{x}=1$ $P=30$ dBm, and $K=3$.
+2-7 evaluate the impact of the BS transmit-power budget $P_{\text{B}}$ under $M=2\times 2$, $N=6\times 6$, and $K=2$.
 Figs.
-4 and 5 examine the impact of the number of transmit antennas for a $1\times M_{y}$ BS array with three users.
-In the isotropic case (Fig.
-4), the sum rate of all schemes increases with $M_{y}$, since a larger array aperture provides stronger array gain and finer spatial resolution to separate the three users while shaping the sensing beampattern.
-Schemes with only BS rotation or only RIS rotation still benefit from adding antennas, but their curves remain noticeably below the fully rotatable configuration, indicating that partial geometric adaptation cannot fully unlock the potential of a large array in the presence of both multiuser interference and sensing requirements.
-The non-RIS baselines lie at the bottom in all $M_{y}$, highlighting the importance of the extra reflecting aperture with more spatial resources.
-In the directional case (Fig.
-5), all curves are further elevated and the advantage of Rot-BS+Rot-RIS becomes more marked.
-The directivity amplifies the relative advantage of the joint BS–RIS rotation: by aligning the main lobe of the transmit array and the dominant reflecting directions of the RIS with the user angles while steering reduced-gain regions toward the sensing directions, the proposed design can more efficiently decouple communication beams from sensing beams, leading to a larger rate gain over the partially rotatable and fully fixed benchmarks.
-In summary, combining array scaling with geometric adaptation is effective for one-dimensional arrays in ISAC systems, especially with directional arrays with practical radiation patterns.
-Figure 6: Sum rate versus the number of communication users under $p^{\text{R}}=p^{\text{B}}=2$ with $N=6\times 6$, $M=2\times 4$, and $P=22$ dBm.
-Figure 7: Sum rate versus the number of communication users under $p^{\text{R}}=p^{\text{B}}=2$ with $N=6\times 6$, $M=2\times 2$, and $P=30$ dBm.
+2-3 plot the weighted utility $\mathcal{U}$, while Figs.
+4-5 and Figs.
+6-7 plot the associated sum rate and NMSE, respectively.
+Half-space isotropic elements ($b=0$): From Figs.
+2, 4, and 6, increasing $P_{\text{B}}$ monotonically improves the achievable utility and sum rate for all schemes, as higher power boosts the effective SNR.
+Meanwhile, the NMSE changes mildly with $P_{\text{B}}$, indicating that with the optimal scaling $\iota^{\star}$, the NMSE is dominated by spatial shape mismatch rather than global amplitude.
+Across the entire power range, “Rot-BS & Rot-RIS” achieves the best performance in both $\mathcal{U}$ and $R$, together with the smallest NMSE.
+This gain comes from the additional geometric DoF provided by jointly rotating the active and passive arrays, which reshapes the array manifolds and effectively enlarges the feasible set of beampatterns under the same electronic beamforming resources.
+Moreover, the RIS-assisted schemes significantly outperform the no-RIS baselines, confirming that the additional passive aperture not only enhances the downlink channels but also provides additional spatial controllability for sensing beampattern shaping.
+Directional elements ($b=2$): Figs.
+3, 5, and 7 show that the benefit of joint rotation becomes substantially more pronounced when practical directional element patterns are employed.
+In this case, misalignment between the array boresight and the desired propagation directions can severely reduce the effective channel gain (and thus increase NMSE), potentially driving the weighted utility to negative values for fixed-orientation baselines.
+By contrast, “Rot-BS & Rot-RIS” maintains the highest utility and sum rate and achieves the lowest NMSE, since it can simultaneously align the BS and RIS boresights with the dominant user/target directions and avoid low-gain regions.
+These results highlight a key LAWN-specific insight: mechanical geometry control is most valuable when the system operates with directional arrays and/or becomes interference-limited, where purely electronic beamforming alone is insufficient to reliably meet both communication and sensing objectives.
+Figure 8: Weighted ISAC utility the number of BS antennas under $b^{\text{R}}=b^{\text{B}}=0$ with $\rho=10$, $N=6\times 6$, $M_{\text{col}}=1$, $P=30$ dBm, and $K=3$.
+Figure 9: Sum rate versus the number of BS antennas under $b^{\text{R}}=b^{\text{B}}=0$ with $\rho=10$, $N=6\times 6$, $M_{\text{col}}=1$, $P=30$ dBm, and $K=3$.
+Figure 10: Beampattern NMSE versus the number of BS antennas under $b^{\text{R}}=b^{\text{B}}=0$ with $\rho=10$, $N=6\times 6$, $M_{\text{col}}=1$, $P=30$ dBm, and $K=3$.
+IV-C Impact of BS Array Size
 Figs.
-6 investigate user scaling for a $2\times 4$ BS array with fixed transmit power $P_{\mathrm{B}}=22$ dBm.
-The sum rate increases with the number of users due to multiuser multiplexing gains, but the growth gradually slows as the system becomes more interference-limited and the available spatial degrees of freedom must be shared between user separation and sensing beampattern shaping.
-The Rot-BS+Rot-RIS scheme consistently delivers the highest sum rate, demonstrating that joint geometry control is especially beneficial in the multiuser regime, whereas partial rotation or fixed architectures suffer from stronger interference and tighter sensing constraints.
+8-10 study the impact of the BS array size by increasing $M_{\text{row}}$ for a $1\times M_{\text{row}}$ BS array under $K=3$ and $P_{\text{B}}=30$ dBm (with $b=0$).
+As $M_{\text{row}}$ increases, the weighted utility and sum rate (Figs.
+8-9) increase steadily because a larger aperture provides higher array gain and more spatial DoF to separate users while shaping the sensing beampattern.
+In addition, the NMSE decreases with $M_{\text{row}}$ (Fig.
+10), showing that more antennas enable a finer beampattern approximation for the same desired template.
+Notably, the proposed “Rot-BS & Rot-RIS” scheme consistently achieves the best utility-rate-NMSE trade-off across all $M_{\text{row}}$, while the partially rotatable and fully fixed schemes exhibit higher NMSE floors.
+This indicates that array scaling and geometry adaptation are complementary: increasing $M$ expands the electronic beamforming space, while rotation expands the geometric manifold of feasible array responses, and the combination yields the most robust ISAC performance.
+Figure 11: Weighted ISAC utility versus the number of communication users under $b^{\text{R}}=b^{\text{B}}=2$ with $\rho=10$, $N=6\times 6$, $M=2\times 4$, and $P=30$ dBm.
+Figure 12: Sum rate versus the number of communication users under $b^{\text{R}}=b^{\text{B}}=2$ with $\rho=10$, $N=6\times 6$, $M=2\times 4$, and $P=30$ dBm.
+Figure 13: Beampattern NMSE versus the number of communication users under $b^{\text{R}}=b^{\text{B}}=2$ with $\rho=10$, $N=6\times 6$, $M=2\times 4$, and $P=30$ dBm.
+IV-D Impact of User Loading
+Figs.
+11-13 investigate user scaling under a $2\times 4$ BS array, $N=6\times 6$, $P_{\text{B}}=30$ dBm, and directional elements ($b=2$).
+As the number of users $K$ increases, the sum rate increases due to multiuser spatial multiplexing gain (Fig.
+12), and consequently the weighted utility also increases (Fig.
+11).
+However, Fig.
+13 shows that the NMSE generally increases with $K$, which reflects the intrinsic ISAC tension: serving more users requires allocating more spatial resources to interference management and data transmission, leaving less flexibility for beampattern shaping.
+Importantly, “Rot-BS & Rot-RIS” consistently yields the highest utility and sum rate while achieving the lowest NMSE across all $K$.
+This demonstrates that joint BS and RIS rotation is particularly beneficial in heavily loaded regimes, where geometric adaptation provides additional leverage to mitigate interference and preserve sensing beampattern fidelity under directional element constraints.
+Figure 14: Weighted ISAC utility versus the maximum rotation range $r$ (degrees) under $b^{\text{R}}=b^{\text{B}}=2$ with $N=6\times 6$, $M=2\times 2$, and $P_{\text{B}}=30$ dBm.
+Figure 15: Sum rate versus the maximum rotation range $r$ under $b^{\text{R}}=b^{\text{B}}=2$ with $N=6\times 6$, $M=2\times 2$, and $P_{\text{B}}=30$ dBm.
+Figure 16: Beampattern NMSE versus the maximum rotation range $r$ (degrees) under $b^{\text{R}}=b^{\text{B}}=2$ with $N=6\times 6$, $M=2\times 2$, and $P_{\text{B}}=30$ dBm.
+IV-E Impact of Rotation Range Constraints
+Figs.
+14-16 evaluate the effect of mechanical rotation limits under $b=2$ with $M=2\times 2$, $N=6\times 6$, and $P_{\text{B}}=30$ dBm.
+For ease of evaluation, we let each Euler angle, in (24d) and (24e), of both arrays sharing the same rotation range $r$ defined by $r_{\mathrm{i},\min}^{\text{A}}=-r,r_{\mathrm{i},\max}^{\text{A}}=r,\forall\text{A}\in\left\{\text{B},\text{R}\right\},\forall\mathrm{i}\in\{\mathrm{x},\mathrm{y},\mathrm{z}\}$.
+As the maximum allowable rotation range $r$ increases, the utility and sum rate improve (Figs.
+14-15), and the NMSE decreases (Fig.
+16), since a larger rotation range provides more geometric DoF to align array boresights and reshape the array response.
+We further observe diminishing returns: a moderate rotation capability already captures most of the gain, and further increasing $r$ yields only marginal improvements.
+Moreover, allowing the BS to rotate is generally more influential than rotating the RIS alone, because BS rotation directly impacts both the direct BS-user channels and the BS-RIS illumination, while RIS rotation mainly affects the reflected link.
+These results suggest that in practical designs, equipping the BS with moderate rotation capability and combining it with even limited RIS rotation can offer a near-optimal rate-sensing trade-off with manageable mechanical complexity.
+[…]
+IV-F Pareto Frontier of the Rate-Sensing trade-off
 Fig.
-7 illustrates the impact of the maximum BS rotation range $\zeta$ on the achievable sum rate for different RIS array rotation ranges.
-As $\zeta$ increases, the sum rate of all schemes improves and gradually saturates, since a larger rotation range provides more geometric degrees of freedom to steer the BS beams toward the users and away from interference directions.
-Across the whole range of $\zeta$, both rotatable-RIS configurations significantly outperform the fixed-IRS and no-IRS baselines, confirming that jointly adapting the BS orientation and the IRS reflection direction is highly beneficial.
-Moreover, the curves for “rot-IRS $90^{\circ}$” and “rot-IRS $45^{\circ}$” almost coincide once $\zeta$ exceeds $30^{\circ}$, indicating that a moderate IRS rotation capability already captures most of the gain brought by passive array rotation, while the BS rotation range remains the dominant factor in leveraging geometric flexibility.
+17 plots the achievable rate-sensing trade-off by sweeping the weight $\rho$ in the weighted-sum objective of (24a).
+Each marker corresponds to the converged solution of (P1) for a given $\rho$, where we sweep $\rho$ on a logarithmic-like grid $\rho\in\{0.1,0.2,\dots,0.9,1,2,$ $\dots,9,10,20,\dots,90,100,200,\dots,1000\}$.
+The horizontal axis reports the beampattern NMSE in (22), where a smaller NMSE indicates a more accurate match to the desired sensing beampattern.
+Hence, moving rightward along each curve corresponds to increasing $\rho$ and placing a stronger emphasis on sensing fidelity, at the cost of reduced downlink sum rate.
+Several key observations can be drawn.
+First, all schemes exhibit a clear and monotonic rate-NMSE trade-off: as $\rho$ increases, the obtained NMSE decreases, but the sum rate gradually degrades.
+Notably, the curves display a “knee” behavior, where moderate sensing improvement can be achieved with a relatively small rate loss, while pushing NMSE further down incurs a disproportionately large throughput sacrifice.
+This knee region offers a practical operating point for low-altitude ISAC deployment that balances sensing fidelity and multiuser throughput.
+Second, the fully rotatable architecture “Rot-BS & Rot-RIS” yields the outermost boundary.
+Intuitively, BS rotation provides a global geometric DoF that simultaneously impacts the direct BS-user links and the BS-side steering of the cascaded BS-RIS-user paths, thereby improving user separation and interference management.
+RIS rotation further enriches the controllable spatial response of the reflected link and enables additional beampattern shaping capability, which becomes particularly valuable when stringent sensing fidelity is required.
+Third, comparing the partially rotatable cases reveals the roles of active versus passive rotations.
+“Rot-BS & Fix-RIS” achieves a markedly better frontier than “Fix-BS & Rot-RIS”, indicating that BS rotation is more influential in this setup as it directly affects all communication and sensing steering vectors.
+Nevertheless, enabling RIS rotation still provides a significant frontier expansion beyond “Rot-BS & Fix-RIS”, especially in the sensing-centric regime, where the passive array can help match the desired beampattern with less distortion to the multiuser downlink beams.
 V Conclusion
-This paper studied a low-altitude ISAC system with cooperative rotatable active and passive arrays.
-We considered a downlink transmission where a BS equipped with an active RA serves multiple ground users and senses low-altitude targets in monostatic mode, assisted by a rotatable RIS, i.e., passive RA.
-To capture the impact of 3D array rotations, we developed a rotation-aware geometry-based multipath channel model in which BS–user, BS–RIS, and RIS–user links are parameterized by path-dependent 3D angles and complex gains, and array rotations jointly affect steering vectors and direction-dependent element gains.
-On this basis, we formulated a unified design that maximizes the multiuser sum rate under a transmit-power budget, RIS unit-modulus constraints, mechanical rotation limits, and a sensing beampattern MSE constraint.
-To tackle the resulting highly coupled and non-convex problem, we proposed a penalty-based AO framework that alternately updates the BS ISAC precoder, RIS phase-shift vector, and BS/RIS rotation angles.
-The BS precoding block is solved via a QT– and MM-aided method, leading to a tractable QCQP; the RIS block is handled on the complex-circle manifold using an RGD algorithm; and the rotation block is optimized via PGD with Barzilai–Borwein step sizes and Armijo backtracking.
-Extensive simulations under LAWN-inspired 3D geometries showed that the proposed jointly rotatable BS–RIS architecture achieves significant sum-rate gains over fixed or partially rotatable baselines while satisfying stringent sensing MSE requirements, with the advantages further amplified with directional antennas.
+This paper investigated a low-altitude ISAC downlink architecture that exploits mechanical rotations at both the active BS array and the passive RIS to introduce additional spatial DoF beyond conventional electronic beamforming.
+By incorporating rotation-dependent array geometry and element directivity into a unified directional channel model, we formulated a joint ISAC design that maximizes a weighted utility comprising the downlink sum rate and the sensing beampattern similarity, measured by NMSE.
+To solve the highly nonconvex problem with coupled variables, we developed an AO framework that iteratively updates the BS transmit beamforming, RIS phase shifts, and the BS and RIS array rotation angles.
+Specifically, the transmit beamforming subproblem is addressed by combining a QT for the sum-rate term with an MM surrogate for the sensing NMSE, yielding a structured QCQP with an efficient closed-form update under a single power constraint.
+The RIS phase update is performed on the complex unit-modulus manifold using RCG, while the rotation variables are updated using a projected gradient method with analytically derived Jacobians.
+Each block update is designed to monotonically improve the objective, and the overall procedure converges to a stationary point of the alternating iterations.
+Numerical results demonstrated that enabling rotations at both the BS and the RIS consistently enlarges the achievable ISAC performance region compared with fixed-array or partially rotatable baselines, with the gains becoming more pronounced under practical directional element patterns.
+Moreover, by sweeping the utility weight, we obtained achievable rate-sensing Pareto frontiers that explicitly quantify how mechanical rotations expand the trade-off boundary.
+Several important extensions remain for future work.
+First, incorporating imperfect CSI and sensing-target uncertainty would enable robust designs aligned with practical estimation errors.
+Second, the time-scale mismatch and actuation overhead of mechanical rotations (e.g., latency, energy consumption, and motion constraints) should be modeled to capture dynamic operation.
+Finally, extending the framework to wideband channels, multi-target sensing, and discrete or quantized RIS hardware constraints would further strengthen the applicability of rotatable active-passive arrays in realistic low-altitude ISAC systems.